### v3

# Mirror Descent Under Generalized Smoothness

Abstract
Smoothness is crucial for attaining fast rates in first-order optimization. However, many optimization problems in modern machine learning involve non-smooth objectives. Recent studies relax the smoothness assumption by allowing the Lipschitz constant of the gradient to grow with respect to the gradient norm, which accommodates a broad range of objectives in practice. Despite this progress, existing generalizations of smoothness are restricted to Euclidean geometry with $\ell_{2}$-norm and only have theoretical guarantees for optimization in the Euclidean space. In this paper, we address this limitation by introducing a new $\ell*$-smoothness concept that measures the norm of Hessians in terms of a general norm and its dual, and establish convergence for mirror-descent-type algorithms, matching the rates under the classic smoothness. Notably, we propose a generalized self-bounding property that facilitates bounding the gradients via controlling suboptimality gaps, serving as a principal component for convergence analysis. Beyond deterministic optimization, we establish sharp convergence for stochastic mirror descent, matching state-of-the-art under classic smoothness. Our theory also extends to non-convex and composite optimization, which may shed light on practical usages of mirror descent, including pre-training and post-training of LLMs.
Generalized smoothness, mirror descent, non-Euclidean geometry
Table 1: Summary of our main results.
Algorithm Convergence Rate Type* Convexity
Mirror Descent (Beck and Teboulle, 2003) $O(1/T)$ (Theorem 3.5) a&l
Accelerated Mirror Descent (Lan, 2020) $O(1/T^{2})$ (Theorem 3.7) l
Optimistic Mirror Descent (Chiang et al., 2012) $O(1/T)$ (Equation 11) a
Mirror Prox (Nemirovski, 2004) $O\left(1/T\right)$ (Equation 13) a
Stochastic Mirror Descent (Nemirovski et al., 2009) $\widetilde{O}(1/\sqrt{T})$ (Theorem 4.2) l
Composite Mirror Descent (Duchi et al., 2010) $O(1/T)$ (Theorem 5.1) g
    Types of convergence, either average-iterate (a), last-iterate (l), or gradient-mapping (g).
1 Introduction
This paper considers the optimization problem $\min_{\mathbf{x}\in\mathcal{X}}\ f(\mathbf{x})$, where $f$ is a convex differentiable function defined on the domain $\mathcal{X}$. It is well-known that gradient descent converges at the rate of $O(1/\sqrt{T})$ for any Lipschitz continuous $f$ (Nesterov, 2013). If $f$ is smooth, i.e., has a Lipschitz continuous gradient, then a faster rate of $O(1/T)$ can be obtained. Furthermore, an optimal $O(1/T^{2})$ rate is attained by incorporating acceleration schemes (Nesterov, 1983). However, optimization problems arising from modern machine learning (ML) are typically non-smooth (Defazio and Bottou, 2019). Even for standard objectives such as $\ell_{2}$-regression, the global smoothness parameter could still be unbounded (Gorbunov et al., 2025).
Recently, efforts have been made to bridge this gap between theory and practice. Zhang et al. (2020b) propose a generalized $(L_{0},L_{1})$-smooth condition, which assumes $\left\|\nabla^{2}f(\mathbf{x})\right\|_{2}\leq L_{0}+L_{1}\left\|\nabla f(\mathbf{x})\right\|_{2}$ for nonnegative constants $L_{0},L_{1}$, on the basis of extensive experimental findings on language and vision models. Under this condition, they analyze the convergence of gradient clipping (Mikolov and others, 2012) and theoretically justify its advantages over gradient descent during neural network training. Later, Li et al. (2023a) propose $\ell$-smoothness which replaces the affine function specified by $(L_{0},L_{1})$ with an arbitrary non-decreasing sub-quadratic function $\ell(\cdot)$, i.e., $\left\|\nabla^{2}f(\mathbf{x})\right\|_{2}\leq\ell(\left\|\nabla f(\mathbf{x})\right\|_{2})$. The mild condition on $\ell(\cdot)$ ensures compatibility with objectives in modern ML and sheds light on new algorithmic designs.
However, existing studies on generalized smoothness suffer from a critical limitation: the definition is limited to $\ell_{2}$-norm and thus can only relate to optimization algorithms based on gradient descent in the Euclidean space. In particular, none of the existing work has considered mirror descent (MD) (Beck and Teboulle, 2003), a powerful optimization algorithm tailored for non-Euclidean problems. Over the years, algorithmic adaptivity to diverse geometries has received tremendous success in traditional ML (Ben-Tal et al., 2001; Zhang et al., 2020c), reinforcement learning (Montgomery and Levine, 2016; Wang et al., 2019; Tomar et al., 2022; Lan, 2023; Alfano et al., 2025), network quantization (Ajanthan et al., 2021), over-parameterization regimes (Sun et al., 2022), watermarking diffusion models (Liu et al., 2023a), pretraining and post-training of LLMs (Xie et al., 2023; Munos et al., 2024; Zhang et al., 2025e, d; Wu et al., 2025). Given the indispensable role of MD in these advancements, a natural question arises:
Can we extend mirror descent under the generalized smoothness condition?
We answer this question affirmatively by devising a new notion of generalized smoothness called $\ell*$-smoothness, which is defined under an arbitrary norm rather than $\ell_{2}$-norm, with full coverage of the $\ell$-smoothness proposed by Li et al. (2023a). To be more specific, we provide a fine-grained characterization of the matrix norm by utilizing a general norm and its dual. Under the $\ell*$-smoothness condition, we establish the convergence of mirror descent, accelerated mirror descent, optimistic mirror descent, and mirror prox algorithms, all matching classic results under $L$-smoothness in deterministic convex optimization. The key ingredient in our analysis is to bridge the gradients’ dual norm with the suboptimality gap. To this end, we establish a generalized self-bounding property under $\ell*$-smoothness (Lemma 3.4), which bounds the gradient’s dual norm by the product of the suboptimality gap and the local Lipschitz constant. For different types of mirror descent algorithms, we employ various strategies to demonstrate that the suboptimality gaps along the optimization trajectory can be bounded by the initial one, which may be of independent interest.
Furthermore, we extend our methodology to stochastic convex optimization and non-convex composite optimization. For the former, we effectively manage the gradient norms by delving into the last-iterate behavior of stochastic mirror descent (SMD) and provide a simple proof based on the “chain-of-events” (Section 4.2) analysis. We derive a state-of-the-art high probability convergence of $O(\sqrt{\log(T)}/\sqrt{T})$, matching the rate for standard $L$-smooth functions. For the latter, we show that the standard $O(1/T)$ convergence for $L$-smooth functions also holds for $\ell*$-smooth function class. Our main results are summarized in Table 1.
Notations   The optima of the objective is denoted by $\mathbf{x}_{*}:\in\operatorname*{argmin}_{\mathbf{x}\in\mathcal{X}}$ and that $f^{*}:=f(\mathbf{x}_{*})$. Let $\left\|\cdot\right\|$ represent a general norm on a finite-dimensional Banach space $\mathcal{E}$, with its corresponding dual norm defined as $\left\|\mathbf{x}\right\|_{*}:=\sup_{\mathbf{y}\in\mathcal{E}}\left\{\langle\mathbf{x},\mathbf{y}\rangle\mid\left\|\mathbf{y}\right\|\leq 1\right\}$ for any $\mathbf{x}\in\mathcal{E}^{*}$, where $\mathcal{E}^{*}$ is the dual space of $\mathcal{E}$. We denote the ball centered at $\mathbf{x}_{0}$ with radius $r_{0}$ by $\mathcal{B}(\mathbf{x}_{0},r_{0}):=\left\{\mathbf{x}\in\mathcal{E}\mid\left\|\mathbf{x}-\mathbf{x}_{0}\right\|\leq r_{0}\right\}$. The rest notaions can be found in Appendix C.
Related work   We briefly introduce two seminal works on generalized smoothness, with the comprehensive literature review deferred to Appendix AZhang et al. (2020b) first observe in LSTM and ResNet training that the loss functions satisfy an affine bound on the Hessian norm, which they formalize as $(L_{0},L_{1})$-smoothness. Under this assumption, they show that simple gradient clipping finds an $\epsilon$-stationary point in $O(\epsilon^{-2})$ iterations for deterministic non-convex problems and $O(\epsilon^{-4})$ for the stochastic counterpart provided bounded noise. More recently, Li et al. (2023a) propose a broader notion called $\ell$-smoothness, where the Hessian is bounded by a non-decreasing function of the gradient norm. This condition captures a wider range of practical models, and under it, they recover the classic convergence rates: logarithmic in $\epsilon^{-1}$ for strongly convex problems, $O(\epsilon^{-1})$ for convex ones, and $O(\epsilon^{-2})$ for non-convex objectives. They further prove that (modified) Nesterov’s accelerated method achieves $O(\epsilon^{-1/2})$ on convex losses, and that SGD still matches the optimal $O(\epsilon^{-4})$ rate in stochastic non-convex settings (Arjevani et al., 2023).
2 Generalized Smooth Function Class
In this section, we formally introduce a refined notion of generalized smoothness on the basis of Li et al. (2023a), so as to accommodate non-Euclidean geometries.
2.1 Definitions and Properties
[…]
Definition 2.1.
($\ell*$-smoothness) A differentiable function $f:\mathcal{X}\mapsto\mathbb{R}$ belongs to the class $\mathcal{F}_{\ell}(\left\|\cdot\right\|)$ w.r.t. a non-decreasing continuous link function $\ell:\mathbb{R}_{+}\mapsto\mathbb{R}_{++}$ if
$\forall\mathbf{h}\in\mathcal{X},\ \left\|\nabla^{2}f(\mathbf{x})\mathbf{h}\right\|_{*}\leq\ell(\left\|\nabla f(\mathbf{x})\right\|_{*})\left\|\mathbf{h}\right\|$ (1)
holds almost everywhere for $\mathbf{x}\in\mathcal{X}$.
Remark 2.2.
Through careful analysis, we discover that it suffices to take the supremum over the domain $\mathcal{X}$ rather than the whole space $\mathcal{E}$, indicating that Definition 2.1 enjoys a slightly weaker condition than Li et al. (2023a).
Remark 2.3.
When $\left\|\cdot\right\|=\left\|\cdot\right\|_{2}$Definition 2.1 simplifies to the $\ell$-smoothness definition in Li et al. (2023a), which recovers the $L$-smooth condition (Nesterov and others, 2018) and recently proposed $(L_{0},L_{1})$-smoothness (Zhang et al., 2020b) with $\ell(\alpha)\equiv L$ and $\ell(\alpha)=L_{0}+L_{1}\alpha$, respectively. In fact, due to the equivalence of norms (Boyd and Vandenberghe, 2004), $\ell$-smoothness can be transformed into $\ell*$-smoothness with a rescaled $\ell(\cdot)$. Whereas, the rescaling brings an additional factor dependent on the problem dimension, ultimately giving rise to dimension-dependent convergence rates.
In general, our formulation encompasses a broader class of functions and also plays a pivotal role in applying mirror-descent-type optimization algorithms. We emphasize that strict twice-differentiability is not required here, as Definition 2.1 permits (1) to fail on a set of measure zero [^1] One can also follow the similar arguments as in Zhang et al. (2020b) to relax the twice-differentiability condition by considering $\limsup_{\mathbf{h}\to\mathbf{0}}\left\|\nabla f(\mathbf{x}+\mathbf{h})-\nabla f(\mathbf{x})\right\|_{*}/\left\|\mathbf{h}\right\|\leq\ell(\left\|\nabla f(\mathbf{x})\right\|_{*})$.Definition 2.1 can be seen as a global condition on the Hessian, but it lacks local characterization of the gradients, which proves to be essential for convergence analysis (Zhang et al., 2020b). Therefore, we introduce the following definition to capture the local Lipschitzness of the gradient, which is equivalent to Definition 2.1 under mild conditions, as illustrated in the sequel.
Definition 2.4.
($(\ell,r)*$-smoothness) A differentiable function $f:\mathcal{X}\mapsto\mathbb{R}$ belongs to the class $\mathcal{F}_{\ell,r}(\left\|\cdot\right\|)$ w.r.t. a non-decreasing continuous link function $\ell:\mathbb{R}_{+}\mapsto\mathbb{R}_{++}$ and a non-increasing continuous radius function $r:\mathbb{R}_{+}\mapsto\mathbb{R}_{++}$ if (i) $\forall\mathbf{x}\in\mathcal{X},\mathcal{B}(\mathbf{x},r(\left\|\nabla f(\mathbf{x})\right\|_{*}))\subseteq\mathcal{X}$; (ii) $\forall\mathbf{x}_{1},\mathbf{x}_{2}\in\mathcal{B}(\mathbf{x},r(\left\|\nabla f(\mathbf{x})\right\|_{*})),$
$\left\|\nabla f(\mathbf{x}_{1})-\nabla f(\mathbf{x}_{2})\right\|_{*}\leq\ell(\left\|\nabla f(\mathbf{x})\right\|_{*})\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|.$ (2)
Similar to Li et al. (2023a), we introduce the following standard assumption to bridge the gap between two function classes.
Assumption 2.5.
The objective function $f$ is differentiable and closed within its open domain $\mathcal{X}$.
Note that when the optimization problem degenerates to the unconstrained setting, this assumption holds naturally for all continuous functions (Boyd and Vandenberghe, 2004). Now we present the following proposition, indicating that $\mathcal{F}_{\ell}(\left\|\cdot\right\|)$ and $\mathcal{F}_{\ell,r}(\left\|\cdot\right\|)$ are nearly equivalent.
Proposition 2.6.
[…]
For simplicity, we restrict our focus to $\mathcal{F}_{\ell}(\left\|\cdot\right\|)$ for the algorithms discussed hereafter. Using Proposition 2.6, we can transform Definition 2.1 into Definition 2.4, enabling a quantitative depiction of the local smoothness property around a point $\mathbf{x}\in\mathcal{X}$, as demonstrated in the following lemma.
Lemma 2.7.
For any $f\in\mathcal{F}_{\ell}(\left\|\cdot\right\|)$ satisfying $\left\|\nabla f(\mathbf{x})\right\|_{*}\leq G\in\mathbb{R}_{+}$ and given $\mathbf{x}\in\mathcal{X}$, it holds that: (i) $\mathcal{B}(\mathbf{x},G/L)\subseteq\mathcal{X}$ where $L:=\ell(2G)$; (ii) $\forall\mathbf{x}_{1},\mathbf{x}_{2}\in\mathcal{B}(\mathbf{x},G/L),\ \left\|\nabla f(\mathbf{x}_{1})-\nabla f(\mathbf{x}_{2})\right\|_{*}\leq L\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|$; (iii) $\forall\mathbf{x}_{1},\mathbf{x}_{2}\in\mathcal{B}(\mathbf{x},G/L),\ f(\mathbf{x}_{1})\leq f(\mathbf{x}_{2})+\langle\nabla f(\mathbf{x}_{2}),\mathbf{x}_{1}-\mathbf{x}_{2}\rangle+L\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|^{2}/2$.
Lemma 2.7 provides (i) an optimistic estimate of the local Lipschitz parameter $\ell(2\left\|\nabla f(\mathbf{x})\right\|_{*})$, and (ii) functional estimation bounds analogous to $L$-smoothness. This establishes a connection between $\mathcal{F}_{\ell}(\left\|\cdot\right\|)$ and the $L$-smooth function class, demonstrating that the former exhibits similar properties as the latter in a local region. As Lemma 2.7 requires $\left\|\nabla f(\mathbf{x})\right\|_{*}\leq G$, the boundedness of gradients plays a critical role in this transformation which will be discussed further in Section 3.2.
2.2 Examples
Li et al. (2023a) provide a series of concrete examples satisfying the $\ell$-smooth condition. Since our formulation generalizes theirs, these example functions also satisfy $\ell*$-smoothness. Nonetheless, we provide additional examples to illustrate the potential emergence of dimension-dependent factors, highlighting the advantages of the algorithmic adaptivity brought by mirror descent. The comprehensive discussions and omitted details can be found in Appendix C.
Consider the logistic kernel and logistic regression function:
$f_{1}(\mathbf{x}):=\exp\left(\mathbf{w}^{\top}\mathbf{x}+b\right),f_{2}(\mathbf{x})=\log\left(1+\exp(-\mathbf{w}^{\top}\mathbf{x})\right)$
for $\mathbf{w},\mathbf{x}\in\mathbb{R}^{n}$, both ubiquitous in machine learning. Under standard $L$-smoothness, $f_{1}$ has an unbounded Hessian (unless $\mathbf{w}=0$), and $f_{2}$’s largest singular value grows like $\|\mathbf{w}\|_{2}^{2}$, which can be arbitrarily large in $\mathbb{R}^{n}$. In contrast, the $(L_{0},L_{1})$-smoothness of Zhang et al. (2020b) remedies this issue, as formalized below.
Proposition 2.8.
(i) $f_{1},f_{2}\in\mathcal{F}_{\widehat{\ell}}(\left\|\cdot\right\|_{2})$ with $\widehat{\ell}(\alpha)=\|\mathbf{w}\|_{2}\alpha$. (ii) $f_{1},f_{2}\in\mathcal{F}_{\widetilde{\ell}}(\left\|\cdot\right\|_{1})$ with $\widetilde{\ell}(\alpha)=\|\mathbf{w}\|_{\infty}\alpha$. (iii) The ratio of link functions satisfies $1/\sqrt{n}\leq\widetilde{\ell}(\alpha)/\widehat{\ell}(\alpha)\leq 1$.
Proposition 2.8 shows that our $\ell*$-smoothness can be significantly tighter than $\ell$-smoothness of Li et al. (2023a), with a worst-case dimensional factor of $1/\sqrt{n}$. In Sections C.2 and C.3, we further exhibit examples where $\widetilde{\ell}(\alpha)/\widehat{\ell}(\alpha)=O(1/n)$ and $O(n^{-0.4})$, respectively, demonstrating even stronger gains. Since both Li et al. (2023a) and our Theorem 3.5 yield convergence rates proportional to the link function $\ell$, mirror descent under $\ell*$-smoothness achieves improved rates by these dimension-dependent factors (see Section C.4 for detailed discussions).
Remark 2.9.
The examples and the justifications in Proposition 2.8 are not confined to unbounded domains like $\mathbb{R}^{n}$. They remain valid under Assumption J.2, as elaborated in Section C.1.3. Moreover, we underline that the constant-link function example in Section C.2 as well as the empirical example in Section C.3 are defined on bounded domains well-modeled by Assumption J.2.
2.3 Empirical Demonstration
Consider the layer-wise non-Euclidean smoothness model proposed in Riabinin et al. (2025, Assumption 1):
$\frac{\|\nabla_{i}f(X)-\nabla_{i}f(Y)\|_{(i)\star}}{\|X_{i}-Y_{i}\|_{(i)}}\leq L^{0}_{i}+L^{1}_{i}\|\nabla_{i}f(X)\|_{(i)\star},$ (3)
where $i\in[K]$ denotes the $i$-th layer of a deep neural network. Notice that our Definitions 2.1 and 2.4 actually cover the above assumption since setting $\left\|\cdot\right\|$ in (2) to $\left\|\cdot\right\|_{(i)}$ directly generates (3). Motivated by this, we experiment on nanoGPT pretraining to demonstrate the address the $\ell*$-smoothness, with results shown in Figure 1. A strong agreement between real and approximated $L_{i}$ values suggests that our $\ell*$-smooth assumption aligns well with the real optimization trajectory of LLMs. Further experimental details can be found in Appendix B.
Figure 1: Trajectory of NanoGPT training on FineWeb dataset
3 Mirror Descent Meets $\ell*$-smoothness
This section formally presents the convergence results for a series of first-order optimization algorithms based on mirror descent, whose analyses can be found in Appendices E, F, G and H.
3.1 Preliminaries
First, we state the general setup of mirror descent algorithms (Beck and Teboulle, 2003).
Definition 3.1.
We call a continuous function $\psi:\mathcal{X}\mapsto\mathbb{R}_{+}$ a distance-generating function with modulus $\alpha$ w.r.t. $\left\|\cdot\right\|$, if (i) the set $\mathcal{X}^{o}=\{\mathbf{x}\in\mathcal{X}|\partial\psi(\mathbf{x})\neq\emptyset\}$ is convex; (ii) $\psi$ is continuously differentiable and $\alpha$-strongly convex w.r.t. $\left\|\cdot\right\|$, i.e., $\langle\nabla\psi(\mathbf{x}_{1})-\nabla\psi(\mathbf{x}_{2}),\mathbf{x}_{1}-\mathbf{x}_{2}\rangle\geq\alpha\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|^{2},\ \forall\mathbf{x}_{1},\mathbf{x}_{2}\allowbreak\in\mathcal{X}^{o}$.
Definition 3.2.
Define the Bregman function $B:\mathcal{X}\times\mathcal{X}^{o}\mapsto\mathbb{R}_{+}$ associated with the distance-generating function $\psi$ as $B(\mathbf{x},\mathbf{x}^{o})=\psi(\mathbf{x})-\psi(\mathbf{x}^{o})-\langle\nabla\psi(\mathbf{x}^{o}),\mathbf{x}-\mathbf{x}^{o}\rangle$.
We equip domain $\mathcal{X}$ with a distance-generating function $\psi(\cdot)$, which is, WLOG, $1$-strongly convex w.r.t. a certain norm $\left\|\cdot\right\|$ endowed on $\mathcal{E}$. We also define the corresponding Bregman divergence according to Definition 3.2. Then, we introduce the following assumptions on the objective function, which controls the growth rate of $\ell(\cdot)$, as also required in Li et al. (2023a).
Assumption 3.3.
The objective satisfies $f\in\mathcal{F}_{\ell}(\left\|\cdot\right\|)$, where $\ell$ is sub-quadratic, i.e., $\lim_{\alpha\to\infty}\allowbreak\alpha^{2}/\ell(\alpha)=\infty$.
For convenience, we also define the following prox-mapping function (Nemirovski, 2004). For any $\mathbf{g}\in\mathcal{E},\mathbf{y}\in\mathcal{X}$,
$\mathcal{P}_{\mathbf{y}}(\mathbf{g}):=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{X}}\left\{\langle\mathbf{g},\mathbf{x}\rangle+B(\mathbf{x},\mathbf{y})\right\},$ (4)
which serves as the basic component of the optimization algorithms being considered.
3.2 Main Idea and Technical Challenges
As discussed previously, bounding the gradients is a prerequisite for Lemma 2.7, which enables us to reduce the $\ell*$-smoothness analysis to the conventional analysis of $L$-smooth functions. In the Euclidean case, Li et al. (2023a) bound the gradients along the optimization trajectory by tracking the gradients directly. More specifically, they manage to show that for gradient descent, the $\ell_{2}$-norm of the gradients decrease monotonically provided that the learning rates are sufficiently small. However, this vital discovery builds upon the geometry of the Euclidean setting, i.e., $\langle\mathbf{x},\mathbf{x}\rangle=\left\|\mathbf{x}\right\|_{2}^{2}$, which is invalid in the non-Euclidean setting.
Technical challenges   The underlying problem is that for mirror descent, the dual norms of the gradients do not necessarily form a monotonic decreasing sequence. The difficulty stems from the nature of mirror descent, where the descent step is performed in the dual space, making the explicit analysis of the dual norms very difficult and rarely addressed. Even under $L$-smoothness, tracking the gradients directly for mirror descent remains challenging, as discussed in Zhang and He (2018); Lan (2020); Huang et al. (2021) and references therein, unless the problem exhibits special structures (Zhou et al., 2017) or extra assumptions are exerted (Lei and Zhou, 2017).
Main idea   To tackle this problem, we introduce the following generalized version of the reversed Polyak-Łojasiewicz inequality (Polyak, 1963; Lojasiewicz, 1963), aka the self-bounding property of smooth functions (Srebro et al., 2010; Zhang et al., 2017; Zhang and Zhou, 2019).
Lemma 3.4.
For any $f\in\mathcal{F}_{\ell}(\left\|\cdot\right\|)$, we have $\left\|\nabla f(\mathbf{x})\right\|_{*}^{2}\leq 2\ell(2\left\|\nabla f(\mathbf{x})\right\|_{*})(f(\mathbf{x})-f^{*})$ for any $\mathbf{x}\in\mathcal{X}$.
Lemma 3.4 measures the growth of gradient’s dual norm by the suboptimality gap $f(\mathbf{x})-f^{*}$. By Assumption 3.3, if we further assume $\ell(\alpha)=(\alpha/2)^{\beta},\beta\in[0,2)$, then $\left\|\nabla f(\mathbf{x})\right\|_{*}\leq 2[f(\mathbf{x})-f^{*}]^{1/(2-\beta)}$ is immediately implied. Hence, Lemma 3.4 suggests that $\left\|\nabla f(\mathbf{x})\right\|_{*}$ is effectively under control if $\ell(\cdot)$ does not grow too rapidly. In this way, the gradient can be bounded by the suboptimality gap of the objective function, which is generally much easier in the context of mirror descent. We rigorously show that the suboptimality gap at an arbitrary iterate is bounded by that of the starting point (see Lemma E.2), which further suggests the boundness of the gradients. Once the gradients have an upper bound, we can construct an effective smoothness parameter $L$ that serves as an optimistic estimation of the local curvatures along the trajectory. Formally, we define
$G=\sup\left\{\alpha\in\mathbb{R}_{+}|\alpha^{2}\leq 2\ell(2\alpha)(f(\mathbf{x}_{0})-f^{*})\right\},$ (5)
$\text{ and }L=\ell(2G),$
where $\mathbf{x}_{0}=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{X}}\psi(\mathbf{x})$ is selected as the starting point for all optimization algorithms considered in this paper, whose trajectory provably satisfies $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G$ (will be stated rigorously in the theorems to come) and consequently, $\ell(2\left\|\nabla f(\mathbf{x}_{t})\right\|_{*})\leq L$. We also emphasize that both $G$ and $L$ are absolute constants, i.e., $G,L<\infty$ (Lemma D.3), which is essential to the validity of our theoretical derivations.
3.3 Mirror Descent and Accelerated Mirror Descent
In this section, we provide theoretical guarantees for mirror descent and its accelerated variant under $\ell*$-smoothness. We start with the basic mirror descent algorithm (Nemirovski and Yudin, 1983; Beck and Teboulle, 2003) given by
$\mathbf{x}_{t+1}=\mathcal{P}_{\mathbf{x}_{t}}(\eta\nabla f(\mathbf{x}_{t})),$ (6)
where $\eta$ represents the learning rates, and the prox-mapping function $\mathcal{P}_{\cdot}(\cdot)$ is defined in (4). Our convergence analysis relies on the interesting fact that the suboptimality gap along the trajectory is a non-increasing sequence: $f(\mathbf{x}_{t})-f^{*}\leq f(\mathbf{x}_{t-1})-f^{*}\leq\cdots\leq f(\mathbf{x}_{0})-f^{*}$, provided that learning rates are chosen appropriately. Consequently, the gradients along the trajectory are bounded by Lemma 3.4, and so are the local Lipschitz constants. The formal statements are illustrated in the following theorem.
Theorem 3.5.
Under Assumptions 2.5 and 3.3, if $0<\eta\leq 1/L$, then the gradients along the trajectory generated by (6) satisfy $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G,\forall t\in\mathbb{N}$, and the average-iterate as well as the last-iterate convergence rate is given by
$\max\left\{f(\overline{\mathbf{x}}_{T})-f^{*},f(\mathbf{x}_{T})-f^{*}\right\}\leq\frac{B(\mathbf{x}_{*},\mathbf{x}_{0})}{\eta T},$ (7)
where $\overline{\mathbf{x}}_{T}=\frac{1}{T}\sum_{t=1}^{T}\mathbf{x}_{t}$.
Our convergence result recovers the classical $O(1/T)$ rate for deterministic convex optimization (Bubeck and others, 2015; Lan, 2020), for both the average-iterate and the last-iterate.
Remark 3.6.
The fact that $G$ is a finite constant is proven after we analyze the trajectory, rather than being used as an explicit assumption to analyze the trajectory. At first glance, it will incur a “circular” analysis. However, this seemingly circularity is exactly the key to our analysis techniques, where we break rigorously in Lemmas D.3, E.2 and 3.5 by induction.
Next, we investigate the accelerated version of mirror descent. The acceleration scheme for smooth convex optimization originates from Nesterov (1983) and has numerous variants (Allen-Zhu and Orecchia, 2017; d’Aspremont et al., 2021). In this paper, we study one of its simplest versions (Lan, 2020):
$\mathbf{y}_{t}=(1-\alpha_{t})\mathbf{x}_{t-1}+\alpha_{t}\mathbf{z}_{t-1},$ (8)
$\mathbf{z}_{t}=\mathcal{P}_{\mathbf{z}_{t-1}}(\eta_{t}\nabla f(\mathbf{y}_{t})),$
$\mathbf{x}_{t}=(1-\alpha_{t})\mathbf{x}_{t-1}+\alpha_{t}\mathbf{z}_{t}.$
We present the following theoretical guarantee, which achieves the optimal $O(1/T^{2})$ convergence rate of first-order methods (Nesterov and others, 2018).
Theorem 3.7.
Under Assumptions 2.5 and 3.3, define [^2] Due to technical reasons, $L=\ell(4G)$ needs to be slightly larger than the one used in mirror descent. This adjustment affects only accelerated mirror descent; other algorithms still use $L=\ell(2G)$.
$L=\ell(4G),\tau=\left\lceil\frac{4\sqrt{2B(\mathbf{x}_{*},\mathbf{x}_{0})}L}{G}\right\rceil-1,\alpha_{t}=\frac{2}{t+1},$
$\eta=\min\left\{1,\frac{3(\tau-1)}{2(\tau-3)(\tau-2)}\right\},\eta_{t}=\frac{t\eta}{2L}.$
Then the convergence rate of (8) is given by
$f(\mathbf{x}_{T})-f^{*}\leq\frac{LB(\mathbf{x}_{*},\mathbf{x}_{0})}{\eta T(T+1)},$ (9)
and the gradients along the trajectory satisfy $\max\{\left\|\nabla f(\mathbf{y}_{t})\right\|_{*},2\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\}\leq 2G$ for all $t\in\mathbb{N}$.
Remark 3.8.
The general idea is similar to mirror descent, that is, bounding the gradients along the trajectory via controlling the suboptimality gaps. However, as (8) maintains three sequences, the analysis becomes significantly more challenging compared to the single sequence $\left\{\mathbf{x}_{t}\right\}_{t\in\mathbb{N}}$ in (6). While $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\lesssim G$ follows directly from the convergence in (9), it cannot be directly extended to the sequence $\{\mathbf{y}_{t}\}_{t\in\mathbb{N}}$, whose gradients also need to be bounded. To address this issue, we focus on the key quantity $e_{t}:=\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|$, which measures the distance between the two sequences. If $e_{t}\lesssim G/L$, we can readily deduce $\left\|\nabla f(\mathbf{y}_{t})\right\|_{*}\lesssim G$ by leveraging the local smoothness property in Definition 2.4. To achieve this goal, we develop a “time partition” technique illustrated in Lemma F.3. Specifically, when $t$ is below the threshold $\tau$, we bound $e_{t}$ by a contraction mapping, which effectively limits its growth for small $t$. Conversely, when $t$ exceeds $\tau$, $e_{t}$ provably decays hyperbolically. Combining the two cases, we can derive that $e_{t}\lesssim G/L$, which suggests the boundness of the gradients.
Comparison with Li et al. (2023a, Theorem 4.4)   For $f\in\mathcal{F}_{\ell}(\left\|\cdot\right\|_{2})$Li et al. (2023a) study a variant of Nesterov’s accelerated gradient algorithm (NAG) and also recover the optimal bound. They introduce an auxiliary sequence into the original NAG, which aims to aggressively stabilize the optimization trajectory. The stabilization mainly refers to $\left\|\mathbf{y}_{t}-\mathbf{x}_{t}\right\|$ in their language, sharing a similar spirit as $e_{t}$ in our paper. To this end, they have to use a smaller step size ($\eta\simeq 1/L^{2}$) and impose a more complex analysis than the original NAG. Unlike their framework, we concentrate on a more stable acceleration scheme in (8), without manually tuning down the base learning rate $\eta$ or bringing in any additional algorithmic components. Moreover, our proof is much simpler and more intuitive, as reflected in our constructive ways of analyzing the stability term $e_{t}$ by partitioning the timeline. It is also worth mentioning that we do not need the fine-grained characterization on the modulus of $\ell$, while Li et al. (2023a) further assume $\ell(u)=O(u^{\alpha}),\alpha\in(0,2]$.
3.4 Optimistic Mirror Descent and Mirror Prox
In this subsection, we consider another genre of the classic optimization algorithm: optimistic mirror descent and mirror prox. Since these algorithms are often treated as a single framework in the online learning community (Chen et al., 2024; Wang et al., 2024b), we adopt the terminology in Mokhtari et al. (2020); Azizian et al. (2021) to avoid potential confusion.
Unlike mirror descent, optimistic mirror descent and mirror prox perform two prox-mappings per iteration and incorporate current observed information to estimate the functional curvature. Firstly, we consider the following optimistic mirror descent algorithm (Popov, 1980; Chiang et al., 2012; Rakhlin and Sridharan, 2013a, b), which requires $1$ gradient query per iteration.
$\mathbf{y}_{t}=\mathcal{P}_{\mathbf{x}_{t-1}}(\eta\nabla f(\mathbf{y}_{t-1})),\ \mathbf{x}_{t}=\mathcal{P}_{\mathbf{x}_{t-1}}(\eta\nabla f(\mathbf{y}_{t})),$ (10)
where we initialize $\mathbf{y}_{0}=\mathbf{x}_{0}=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{X}}\psi(\mathbf{x})$. The following theorem establishes the classic $O(1/T)$ convergence rate for this algorithm.
Theorem 3.9.
Under Assumptions 2.5 and 3.3, if $0<\eta\leq 1/(3L)$, then the gradients along the trajectory generated by (10) satisfy $\max\{\left\|\nabla f(\mathbf{y}_{t})\right\|_{*},\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\}\leq G,\forall t\in\mathbb{N}$, and the convergence rate is given by
[…]
Next, we investigate the mirror prox algorithm (Korpelevich, 1976; Nemirovski, 2004; Nesterov, 2007), which is defined via the following updates:
$\mathbf{y}_{t}=\mathcal{P}_{\mathbf{x}_{t-1}}(\eta\nabla f(\mathbf{x}_{t-1})),\ \mathbf{x}_{t}=\mathcal{P}_{\mathbf{x}_{t-1}}(\eta\nabla f(\mathbf{y}_{t})).$ (12)
The only difference between (10) and (12) is that the latter does not reuse gradient information, which results in $2$ gradient queries per iteration. For a comprehensive overview and comparison between these two algorithms, one may refer to Azizian et al. (2021); Cai et al. (2022a). The same $O(1/T)$ convergence rate is also obtained, as shown in the theorem below.
Theorem 3.10.
Under Assumptions 2.5 and 3.3, if $0<\eta\leq 1/(2L)$, then the gradients along the trajectory generated by (12) satisfy $\max\{\left\|\nabla f(\mathbf{y}_{t})\right\|_{*},\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\}\leq G,\forall t\in\mathbb{N}$, and the convergence rate is given by
[…]
The high-level idea behind Equations 11 and 13 aligns with that in Section 3.3, i.e., bounding the dual norms of the gradients by analyzing suboptimality gaps. However, the implication from the last-iterate convergence in Theorems 3.5 and 3.7 to the bounded suboptimality gap does not apply for optimistic mirror descent and mirror prox, as elaborated below.
Hardness results on the last-iterate   Different from the algorithms in Section 3.3 which enjoys a non-asymptotic last-iterate convergence under $L$-smoothness, the last-iterate of (10) as well as (12) had been only known to converge asymptotically, as shown by Popov (1980); Hsieh et al. (2019) for (10); Korpelevich (1976); Facchinei and Pang (2007) for (12). Even in the Euclidean setting, the finite-time last-iterate convergence is very challenging (Golowich et al., 2020; Wei et al., 2021; Gorbunov et al., 2022) and requires complicated analysis as well as intricate techniques (Cai et al., 2022b), e.g., computer-aided proofs based on sum-of-squares programming (Nesterov, 2000; Parrilo, 2003).
Circumvent the last-iterate analysis   Recall the discussions in Section 3.2 that it suffices to ensure the boundness of the suboptimality gap, which can be implied by the last-iterate convergence. Since the latter one is a stronger statement, we can bypass the last-iterate convergence analysis, and control the suboptimality gap directly through careful analysis. Therefore, the cumbersome procedures in Cai et al. (2022b) can be avoided. Below, we briefly outline our approach.
Solution   For the mirror prox method, we establish the descent property for both sequences (Lemma H.2), i.e., $\max\left\{f(\mathbf{x}_{t}),f(\mathbf{y}_{t})\right\}\leq f(\mathbf{x}_{t-1})$, which guarantees the bounded suboptimality gap. For the optimistic mirror descent method, we track the stability term $\left\|\mathbf{y}_{t}-\mathbf{y}_{t-1}\right\|$ and show that it grows no faster than the geometric series (Lemma G.2). The upper bound on stability terms suggests that the optimization trajectory is effectively stabilized, ultimately yielding $\max\left\{f(\mathbf{x}_{t}),f(\mathbf{y}_{t})\right\}\leq f(\mathbf{x}_{0})$. Consequently, we bound both the suboptimality gap and the gradient.
4 Stochastic Convex Optimization
In this section, we extend from deterministic optimization to stochastic convex optimization.
4.1 Theoretical Results
Consider the following stochastic mirror descent (SMD) algorithm (Nemirovski et al., 2009):
$\mathbf{x}_{t+1}=\mathcal{P}_{\mathbf{x}_{t}}(\eta_{t+1}\mathbf{g}_{t}),$ (14)
where $\mathbf{g}_{t}$ is an noisy estimate of the true gradient $\nabla f(\mathbf{x}_{t})$. Let $\boldsymbol{\epsilon}_{t}:=\mathbf{g}_{t}-\nabla f(\mathbf{x}_{t})$ denote the noise. We make the following sub-Gaussian assumption on $\boldsymbol{\epsilon}_{t}$, which is common for the high probability convergence (Juditsky et al., 2011; Vershynin, 2018; Zhang et al., 2023; Liu and Zhou, 2024).
Assumption 4.1.
For all $t\in\mathbb{N}$, the stochastic gradients are unbiased: $\mathbb{E}_{t-1}[\boldsymbol{\epsilon}_{t}]=0$, where the expectation $\mathbb{E}_{t-1}$ is conditioned on the past stochasticity $\left\{\boldsymbol{\epsilon}_{s}\right\}_{s=0}^{t-1}$. The noise level further satisfies:
$\mathbb{E}_{t-1}\left[\exp\left(\lambda\left\|\boldsymbol{\epsilon}_{t}\right\|_{*}^{2}\right)\right]\leq\exp\left(\lambda\sigma^{2}\right),\forall\lambda\in[0,\sigma^{-2}].$
Now, we formally introduce the convergence of SMD, where the analysis is postponed to Appendix I.
Theorem 4.2.
Under Assumptions 2.5, 3.3 and 4.1, for any $0<\delta<1$, define
$G:=\sup\left\{\alpha\in\mathbb{R}_{+}|\alpha^{2}\leq 2\ell(2\alpha)\left(f(\mathbf{x}_{0})-f^{*}+\frac{64\sigma}{\sqrt{\delta}}\right)\right\},$
$\eta=\min\left\{\frac{64}{G},\sqrt{\frac{B(\mathbf{x}_{*},\mathbf{x}_{0})}{\sigma^{2}\log\left(\frac{1}{\delta}\right)\log(T)}}\right\},\quad L:=\ell(2G).$
Set $\eta_{t}=\min\left\{\frac{1}{2L},\frac{\eta}{\sqrt{T}}\right\}$, with probability at least $1-\delta$:
$f(\mathbf{x}_{T})-f^{*}$
$\leq$ $O\left(\frac{LB(\mathbf{x}_{*},\mathbf{x}_{0})}{T}+\sigma\sqrt{\frac{B(\mathbf{x}_{*},\mathbf{x}_{0})\log\left(\frac{1}{\delta}\right)\log(T)}{T}}\right).$
Moreover, $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G<+\infty$ holds for all $0\leq t\leq T-1$ with probability at least $1-\delta$.
The above theorem matches the state-of-the-art last-iterate convergence rate in Liu and Zhou (2024). Notably, our result is valid under $\ell*$-smooth function class, which is more general than the $(L,M)$-function class [^3] Liu and Zhou (2024, Assumption 3) state that $f$ is $(L,M)$-smooth such that $f(\mathbf{x})\leq f(\mathbf{y})+\langle\mathbf{g}_{\mathbf{y}},\mathbf{x}-\mathbf{y}\rangle+\frac{L}{2}\left\|\mathbf{x}-\mathbf{y}\right\|^{2}+M\left\|\mathbf{x}-\mathbf{y}\right\|,\forall\mathbf{x},\mathbf{y}\in\mathcal{X},\mathbf{g}_{\mathbf{y}}\in\partial f(\mathbf{y})$. in Liu and Zhou (2024). More importantly, our analysis is adaptive to noise, i.e., when $\sigma=0$ degenerates to deterministic optimization, the convergence in Theorem 4.2 encompasses that in Theorem 3.5 (deterministic mirror descent). Our theory answers affirmatively to the application of SMD to train real-world ML models, which is typically non-smooth and may satisfy our $\ell*$-smoothness condition (see our nanoGPT experiments in Section 2.3 as well as Appendix B).
Apart from the aforementioned convergence under sub-Gaussian noise, we also propose a new generalized bounded noise model (Assumption J.1) and establish a time-uniform convergence rate of $\widetilde{O}(1/\sqrt{t})$ (Theorem J.3). Due to space limitations, we defer this part to Appendix J.
4.2 Proof Sketch
Similar to the deterministic case, there are two major challenges underlying the convergence analysis for generalized smooth functions: (i) ensure in each step $t$, the iterates $\mathbf{x}_{t+1},\mathbf{x}_{t}$ are close enough to exploit the effective smooth property in Lemma 2.7; and (ii) ensure the local curvatures $\ell(2\left\|\nabla f(\mathbf{x}_{t})\right\|_{*})$ to be bounded uniformly by an absolute constant. For the second challenge, we can achieve this goal by bounding gradient norms $\left\{\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\right\}_{t=0}^{T-1}$, sharing a similar spirit as in Section 3. For the first challenge, the deterministic case begins with $\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|\leq\eta\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}$. Since we have already controlled $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}$ in the second challenge, the first challenge has been resolved simultaneously if the learning rate $\eta$ is set appropriately. However, for SCO, we write
$\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|\leq\eta_{t+1}\left\|\mathbf{g}_{t}\right\|_{*}\leq\eta_{t+1}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}}+\eta_{t+1}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\left\|\boldsymbol{\epsilon}_{t}\right\|_{*}}.$
The gradient norm term $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}$} is effectively managed due to challenge (ii). For the noise term $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\left\|\boldsymbol{\epsilon}_{t}\right\|_{*}$}, we seek help from Assumption 4.1, which has good properties. We show that throughout $T$ iterations, the probability of “bad” noise estimates (with large noise norms $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\left\|\boldsymbol{\epsilon}_{t}\right\|_{*}$}) under Assumption 4.1 can be relatively low.
How do we manage $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}$} and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\left\|\boldsymbol{\epsilon}_{t}\right\|_{*}$}?   We define the following events
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}A_{t}}:=\left\{f(\mathbf{x}_{t})-f^{*}\leq F\right\},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}B_{t}}:=\left\{\left\|\boldsymbol{\epsilon}_{t}\right\|_{*}\leq\frac{G}{2\eta L}\right\},$
and show that $\cup_{t=0}^{T-1}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}A_{t}}$ and $\cup_{t=0}^{T-1}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}B_{t}}$ happen simultaneously with high probability along the optimization trajectory. For $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}A_{t}$}, it is achieved by carefully analyzing the last-iterate behavior of (14): $f(\mathbf{x}_{t})-f(\mathbf{x})\leq b_{t}(B(\mathbf{x},\mathbf{x}_{0}))$, where $b_{t}$ denotes the bound depending on $B(\mathbf{x},\mathbf{x}_{0})$ for step $t$. We take the reference point $\mathbf{x}=\mathbf{x}_{0}$ to enable cancellation of terms, and obtain $f(\mathbf{x}_{t})-f(\mathbf{x}_{0})\leq b_{t}(0)$. By showing that $b_{t}$ is non-increasing in $t$, we are able to bound $f(\mathbf{x}_{t})-f^{*}\leq f(\mathbf{x}_{0})-f^{*}+\max_{t}b_{t}=O(1)$. For $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}B_{t}$}, we use Chebyshev’s inequality and take a union bound over the time horizon. The most subtle point underlying the analysis is that we have to perform a “chain-of-event” analysis, rather than taking union bounds brute-force. This is because when we analyze the failure probability of $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}A_{t}$}, previous events must happen, otherwise we are unable to exploit generalized smoothness. Such “chain-like” characteristics requires taking union bounds based on mathematical induction.
5 Non-convex Optimization
In this section, we extend $\ell*$-smoothness to non-convex composite optimization (Ghadimi et al., 2016; Lan, 2020), which can be formulated as
$F^{*}:=\min_{\mathbf{x}\in\mathcal{X}}\left\{F(\mathbf{x}):=f(\mathbf{x})+\phi(\mathbf{x})\right\}$ (15)
where $f\in\mathcal{F}_{\ell}(\left\|\cdot\right\|)$ and is possibly non-convex; $\phi:\mathcal{X}\mapsto\mathbb{R}_{+}$ is a simple convex regularizer, but possibly non-smooth (e.g., $\phi(\mathbf{x})\equiv 0$ or $\phi(\mathbf{x})=\left\|\mathbf{x}\right\|_{1}$). (15) can model a wide range of problems in ML since the objective $F$ is neither convex nor smooth. To optimize (15), the de facto method is composite mirror descent (CMD) (Duchi et al., 2010; Ghadimi et al., 2016; Lei and Tang, 2018; Lan, 2020; Wang et al., 2024b):
$\mathbf{x}_{t+1}=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{X}}\left\{\langle\nabla f(\mathbf{x}_{t}),\mathbf{x}\rangle+\frac{B(\mathbf{x},\mathbf{x}_{t})}{\eta}+\phi(\mathbf{x})\right\},$
where we assume access to exact gradients. In the sequel, we study the convergence behavior of CMD under $\ell*$-smoothness and show that it still achieves the classic rate under our relaxed smoothness model.
To start, we define the following gradient mapping function (Nesterov and others, 2018) as the convergence criterion:
$\mathcal{G}_{t}:=\frac{\mathbf{x}_{t}-\mathbf{x}_{t+1}}{\eta},$ (16)
which is widely used in the context of mirror descent (Ghadimi et al., 2016; Lan, 2020; Huang et al., 2021). As pointed out by Lan (2020, Lemma 6.3), when the size of $\mathcal{G}_{t}$ vanishes, $\mathbf{x}_{t+1}$ approaches to a stationary point of (15). For more comprehensive illustrations, the readers may refer to Lan (2020), and references therein. Below, we formally present the theoretical guarantee, whose proof is located in Appendix K.
Theorem 5.1.
Under Assumptions 2.5 and 3.3, let
$G:=\sup\left\{\alpha\in\mathbb{R}_{+}|\alpha^{2}\leq 2\ell(2\alpha)\left(f(\mathbf{x}_{0})-f^{*}+\phi(\mathbf{x}_{0})\right)\right\},$
$L:=\ell(2G),\text{ and }0<\eta\leq\frac{1}{L}.$
Then, CMD enjoys the following convergence rate:
$\frac{1}{T}\sum_{t=0}^{t-1}\left\|\mathcal{G}_{t}\right\|^{2}\leq\frac{F(\mathbf{x}_{0})-F^{*}}{\eta T}.$ (17)
The convergence in (17) matches the one derived under the standard smoothness conditions (Ghadimi et al., 2016; Lan, 2020). This result provides strong theoretical support for the practical application of (composite) mirror descent, as modern neural networks are non-convex and non-smooth. Our result may also shed light on the application of CMD in LLM pretraining (Xie et al., 2023) and preference alignment (Munos et al., 2024; Zhang et al., 2025e, d; Wu et al., 2025), which has been empirically observed to exhibit generalized smooth properties (Crawshaw et al., 2022; Liu et al., 2025b; Riabinin et al., 2025).
6 Conclusion
In this paper, we propose a refined smoothness notion called $\ell*$-smoothness, which characterizes the norm of the Hessian by leveraging a general norm and its dual. This extension encompasses the current generalized smooth conditions while highlighting non-Euclidean problem geometries. We establish new convergence results of mirror descent, accelerated mirror descent, optimistic mirror descent, mirror prox, and stochastic mirror descent algorithms under this setting. Beyond convex optimization, we further delve into non-convex composite mirror descent and achieve tight convergence rates. Our theory provides insights into the practical usage of mirror descent, especially in pretraining and preference alignment of LLMs (Xie et al., 2023; Zhang et al., 2025e, d).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Related Work
In this section, we review some concepts of generalized smoothness from the literature.
A.1 $(L_{0},L_{1})$-smoothness
$(L_{0},L_{1})$-smooth: non-convex case
The concept of $(L_{0},L_{1})$-smoothness is firstly proposed by Zhang et al. (2020b) based on empirical observations from LSTMs (Merity et al., 2018) and ResNets (He et al., 2016), which allows the function to have an affine-bounded Hessian norm. Under this new condition, they analyze gradient clipping (Mikolov and others, 2012; Pascanu et al., 2013) for deterministic non-convex optimization and derive a complexity of $O(\epsilon^{-2})$ [^4] In non-convex settings, the goal is to find $\epsilon$-stationary point $\mathbf{x}$ satisfying $\left\|\nabla f(\mathbf{x})\right\|_{2}\leq\epsilon$. In convex settings, we aim at $\epsilon$-suboptimal point $\mathbf{x}$ satisfying $f(\mathbf{x})-f^{*}\leq\epsilon$. The meaning of $\epsilon$ depends on the context., which matches the lower bound (Carmon et al., 2020) up to constant factors. They further extend to stochastic settings and provide an $O(\epsilon^{-4})$ complexity bound under the uniformly bounded noise assumption. Gaash et al. (2025) investigate the high-probability convergence of SGD with clipping, and obtain convergence rate that matches SGD up to polylogarithmic factors and additive terms. Yu et al. (2026) analyze both vector and matrix sign-based optimizers, namely SignSGD (Bernstein et al., 2018), Lion (Chen et al., 2023a), Muon (Jordan et al., 2024), and Muonlight (Liu et al., 2025a), under the generalized heavy-tailed noise conditions. They consider coordinate-wise $(L_{0},L_{1})$-smooth function class and obtain a complexity of $O(\epsilon^{-\frac{3p-2}{p-1}})$ for all methods, matching state-of-the-art.
$(L_{0},L_{1})$-smooth: convex case
Koloskova et al. (2023) show that gradient clipping has $O(\epsilon^{-1})$ complexity, matching the classic result in deterministic optimization. Takezawa et al. (2024) establish the same complexity bound for gradient descent with polyak stepsizes (Polyak, 1987). However, besides $(L_{0},L_{1})$-smoothness, these two works further impose an additional $L$-smooth assumption, where $L$ could be significantly larger than $L_{0}$ and $L_{1}$Gorbunov et al. (2025) address the limitation and recover their results without the extra $L$-smoothness assumption. Moreover, they study a variant of adaptive gradient descent (Malitsky and Mishchenko, 2020) and provide an $O(\epsilon^{-1})$ complexity result, albeit with worse constant terms compared to the original one. For acceleration schemes in convex optimization, they modify the method of similar triangles (MST) (Gasnikov and Nesterov, 2018) and prove an optimal complexity of $O(\epsilon^{-0.5})$Lobanov et al. (2024) provide a refined convergence analysis of gradient descent and achieve a linear speedup.
Other explorations
Ever since Zhang et al. (2020b) proposed $(L_{0},L_{1})$-smoothness condition, this generalized notion of smoothness has been flourishing in minimax optimization (Xian et al., 2024), bilevel optimization (Hao et al., 2024; Gong et al., 2024), multi-objective optimization (Zhang et al., 2025a), sign-based optimization (Crawshaw et al., 2022), zero-th order optimization (Lobanov and Gasnikov, 2025), distributionally robust optimization (Jin et al., 2021; Zhang et al., 2025b) and variational inequality (Vankov et al., 2024b, a). Numerous attempts have been made to refine existing algorithms under this weaker assumption, including variance reduction (Reisizadeh et al., 2025), clipping/normalized gradient (Zhang et al., 2020a; Qian et al., 2021; Zhao et al., 2021; Hübler et al., 2024; Yang et al., 2024), error feedback (Khirirat et al., 2024b) and trust region methods (Xie et al., 2024a). Notably, Faw et al. (2022, 2023); Wang et al. (2023b); Li et al. (2023b); Wang et al. (2024a); Zhang et al. (2025c) explore the convergence of AdaGrad (Duchi et al., 2011), RMSprop (Hinton et al., 2012), and Adam (Kingma and Ba, 2015) under generalized smoothness. Beyond the optimization community, it has also garnered significant attention in federated learning (Khirirat et al., 2024a; Demidovich et al., 2025), meta-learning (Chayti and Jaggi, 2024), and fine-tuning LLMs (Ma et al., 2025).
A.2 $\alpha$-symmetric Generalized Smoothness
An important generalization of $(L_{0},L_{1})$-smoothness is $\alpha$-symmetric $(L_{0},L_{1})$-smoothness proposed by Chen et al. (2023b), which introduces symmetry into the original formulation and also allows the dependency on the gradient norm to be polynomial with degree of $\alpha$. The formal definition is given as follows:
$\left\|\nabla f(\mathbf{x})-\nabla f(\mathbf{y})\right\|_{2}\leq\left(L_{0}+L_{1}\sup_{\theta\in[0,1]}\left\|\nabla f\left(\theta\mathbf{x}+(1-\theta)\mathbf{y}\right)\right\|_{2}\right)\left\|\mathbf{x}-\mathbf{y}\right\|_{2},\ \forall\mathbf{x},\mathbf{y}\in\mathcal{E},$ (18)
for some $L_{0},L_{1}\in\mathbb{R}_{+}$. For deterministic non-convex optimization, Chen et al. (2023b) establish the optimal complexity of $O(\epsilon^{-2})$ for a variant of normalized gradient descent (Nesterov, 1984; Cortés, 2006). They also show that the popular SPIDER algorithm (Fang et al., 2018) achieves the optimal $O(\epsilon^{-3})$ complexity in the stochastic setting. As pointed out by Chen et al. (2023b); Vankov et al. (2025), any twice-differentiable function satisfying (18) is also $(L_{0}^{\prime},L_{1}^{\prime})$-smooth for different constant factors. As an extension, Jiang et al. (2024) study sign-based finite-sum non-convex optimization and obtain an improved complexity over SignSVRG algorithm (Chzhen and Schechtman, 2023).
A.3 $\ell$-smoothness
Recently, Li et al. (2023a) significantly generalize $(L_{0},L_{1})$-smoothness to $\ell$-smoothness, which assumes $\left\|\nabla^{2}f(\mathbf{x})\right\|_{2}\leq\ell(\left\|\nabla f(\mathbf{x})\right\|_{2})$ for an arbitrary non-decreasing function $\ell$. The enhanced flexibility of $\ell$ accommodates a broader range of practical ML problems (Devlin et al., 2019; Caron et al., 2021; Radford et al., 2021), and can model certain real-world problems where $(L_{0},L_{1})$-smoothness fails (Cooper, 2024). Under this condition, they study gradient descent for strongly-convex, convex, and non-convex objectives, obtaining classic results of $O(\log(\epsilon^{-1}))$, $O(\epsilon^{-1})$, and $O(\epsilon^{-2})$, respectively. They also prove that Nesterov’s accelerated gradient method (Nesterov, 1983) attains the optimal $O(\epsilon^{-0.5})$ complexity for convex objectives. Furthermore, they delve into stochastic non-convex optimization and show that stochastic gradient descent achieves the optimal complexity of $O(\epsilon^{-4})$ under finite variance condition, matching the lower bound in Arjevani et al. (2023).
Tyurin (2025) improves the convergence rate from Li et al. (2023a) in the convex setting, and also enhances the rate in the non-convex setting when $\ell$ has a sub-quadratic and quadratic growth. However, for super-quadratic or even exponential $\ell$ in the non-convex regimes, Tyurin (2025) imposes an additional assumption of bounded gradients. Xie et al. (2024b) study online convex optimization problems by assuming each online function $f_{t}:\mathcal{X}\mapsto\mathbb{R}$ is $\ell_{t}$-smooth and that $\ell_{t}(\cdot)$ can be queried arbitrarily by the learner. Based on $\ell_{t}$-smoothness, they derive gradient-variation regret bounds for convex and strongly convex functions, albeit with the limitation of assuming a globally constant upper bound on the adversary’s smoothness.
A.4 Other Generalizations of Classic Smoothness
It is known that a function $f$ is $L$-smooth if LhfLh-f is convex (Nesterov and others, 2018) for $h(\cdot)=\left\|\cdot\right\|^{2}/2$. Advancements have been made to relax $h$ into an arbitrary Legendre function, leading to the development of a notion called relative smoothness (Birnbaum et al., 2011; Bauschke et al., 2017; Lu et al., 2018; Hanzely et al., 2021; Hanzely and Richtárik, 2021), which substitutes the canonical quadratic upper bound with $f(\mathbf{x})\leq f(\mathbf{y})+\langle\nabla f(\mathbf{y}),\mathbf{x}-\mathbf{y}\rangle+LB_{h}(\mathbf{x},\mathbf{y})$, where $B_{h}$ is the Bregman divergence associated with $h$. Since relative smoothness is now standard in the analysis of mirror descent, we provide a detailed comparison with our $\ell*$-smoothness in Section A.5.
More recently, Mishkin et al. (2024) propose directional smoothness, a measure of local gradient variation that preserves the global smoothness along specific directions. To address the imbalanced Hessian spectrum distribution in practice (Sagun et al., 2016; Pan et al., 2022)Liu et al. (2025b); Jiang et al. (2025) propose anisotropic smoothness and obtain improved bounds for adaptive gradient methods.
A.5 Comparison with Relative Smoothness
According to Lu et al. (2018), a function $f$ is $L$-smooth relative to $h$ iff LhfLh-f is convex, which is equivalent to $\nabla^{2}f(x)\preceq L\nabla^{2}h(x)$ when both functions are twice-differentiable on the interior of the domain. Notably, $h$ is not presumed to possess any special properties like strict or strong convexity. In the sequel, we compare this notion with $\ell*$-smoothness and highlight why ours is more practical.
Practical infeasibility
Relative smoothness constrains the Hessian spectrum via a reference function $h$. However, for any $f$, there exists infinitely many valid $(L,h)$ pairs satisfying the definition. Therefore, finding the most appropriate pair is essential under this framework. Unfortunately, in most cases, $(L,h)$ can only be determined through delicate design (requiring prior knowledge of $f$). In contrast, $\ell*$-smoothness bounds the matrix primal-dual norm of the Hessian via a link function of the gradient dual norm, i.e., $\left\|\nabla^{2}f(\mathbf{x})\right\|_{*,\cdot}:=\sup_{\mathbf{h}}\left\{\left\|\nabla^{2}f(\mathbf{x})\mathbf{h}\right\|_{*}/\left\|\mathbf{h}\right\|\right\}\leq\ell(\left\|\nabla f(\mathbf{x})\right\|_{*})$. This formulation does not involve Löwner partial order $\preceq$ and is generally much easier to verify, as it compresses the information of eigenvalues into a specific norm and connects it with gradients. Moreover, recent advances in ML suggest that $\ell*$-smoothness is satisfied by many practical objectives both theoretically and empirically (see references discussed earlier), whereas relative smoothness remains less prevalent in the ML community. To summarize, both notions extend the traditional smoothness from different perspectives: relative smoothness is more direct and strict, while generalized smoothness is more flexible and practical. Below, we provide concrete examples and further reasoning to support our claims.
Examples
Consider the example in Lu et al. (2018, Proposition 2.1), where $\left\|\nabla^{2}f(\mathbf{x})\right\|_{2}\leq\allowbreak p_{r}(\left\|\mathbf{x}\right\|_{2})$ for some $r$-degree polynomial $p_{r}$. Then $f$ is $L$-smooth relative to $h$ with $L=\sup_{a>0}p_{r}(a)/(1+a^{r}),h(\mathbf{x})=\left\|\mathbf{x}\right\|_{2}^{r+2}/(r+2)+\left\|\mathbf{x}\right\|_{2}^{2}/2$. In this case, the subproblem has closed-form solutions only for $r=1,2,3$, otherwise an extra root-finding oracle is required, which is unsatisfactory. More importantly, this framework only accommodates functions whose Hessian is bounded by polynomials of $\left\|\mathbf{x}\right\|_{2}$, which is restrictive. For instance, consider $g(\mathbf{x})=\exp(\left\|\mathbf{x}\right\|_{2}^{2}/2)$. It holds that $\left\|\nabla^{2}g(\mathbf{x})\right\|_{2}\leq(1+\left\|\mathbf{x}\right\|_{2}^{2})\exp(\left\|\mathbf{x}\right\|_{2}^{2}/2)$. Thus the Hessian grows exponentially w.r.t. $\left\|\mathbf{x}\right\|_{2}$, rendering polynomial-based characterization inadequate. However, since $\|\nabla g(\mathbf{x})\|_{2}=\left\|\mathbf{x}\right\|_{2}\exp(\left\|\mathbf{x}\right\|_{2}^{2}/2)$, it follows that $g$ is $(0,\left\|\mathbf{x}\right\|_{2}+1)$-smooth. Evidently, generalized smoothness can model this function class better than relative smoothness. Furthermore, the rest examples in Lu et al. (2018) all conform to our $\ell*$-smoothness framework.
Key limitations
In summary, relative smoothness has the following crucial limitations.
  1. 1.
    In practice, selecting a meaningful $(L,h)$ is challenging even with complete problem information. Finding solvable subproblems remains difficult.
  2. 2.
    For real-world applications where only a first-order oracle of $f$ is available (e.g., neural networks), relative smoothness lacks an effective solution. In contrast, generalized smoothness remains applicable by empirically estimating the link function $\ell$ (Zhang et al., 2020b).
Appendix B Experimental Details for LLMs
(a) Validation of layer-wise $(L^{0},L^{1})$-smoothness for the group of parameters from the transformer blocks of NanoGPT-124M along unScion training trajectories. The group norms are non-euclidean norms $\|\cdot\|_{(i)}=\sqrt{n_{i}/m_{i}}\|\cdot\|_{2\to 2}$.
All experiments for the NanoGPT model (Karpathy, 2022) are conducted using PyTorch [^5] PyTorch Documentation. Available at: https://pytorch.org/docs/stable/index.html with Distributed Data Parallel (DDP) [^6] Distributed Data Parallel (DDP) in PyTorch. Available at: https://pytorch.org/docs/stable/notes/ddp.html across 8 NVIDIA 4090 GPUs (24GB each). The experiments are based on an open-source codebase which can be found at https://github.com/artem-riabinin/Experiments-estimating-smoothness-for-NanoGPT-and-CNN/. Theoretical formulations and details for the codebase can be found at Riabinin et al. (2025, Appendix E).
We mainly aimed at empirically validating our generalized smoothness model under non-Euclidean norms. In the procedure of training nanoGPT on the Fineweb dataset (Penedo et al., 2024) with unScion (Riabinin et al., 2025), we plot the estimated trajectory smoothness
$\hat{L}_{i}[k]\coloneqq\frac{\|\nabla_{i}f_{\xi^{k+1}}(X^{k+1})-\nabla_{i}f_{\xi^{k}}(X^{k})\|_{(i)\star}}{\|X_{i}^{k+1}-X_{i}^{k}\|_{(i)}},$
and its approximation
$\hat{L}_{i}^{\text{approx}}[k]\coloneqq L_{i}^{0}+L_{i}^{1}\|\nabla_{i}f_{\xi^{k+1}}(X^{k+1})\|_{(i)\star}$
as functions of the iteration index $k$. We observed similar findings to Riabinin et al. (2025, Appendix E.3.1), where a there exihibits clear alignment between $\hat{L}_{i}[k]$ and $\hat{L}_{i}^{\text{approx}}[k]$, suggesting that our smoothness model holds approximately under non-Euclidean norms. Figure 5(a) shows more results for parameter groups at different transformer blocks of NanoGPT-124M.
Appendix C The Gain of Algorithmic Adaptivity
In this section, we present the omitted details for Section 2.2. The proof of Proposition 2.8 is essentially the discussions underlying Section C.1, which are inspired by Gorbunov et al. (2025). First, we present the omitted notations in the main text.
Omitted notations
[…]
C.1 Theoretical Justifications: Non-constant Link Functions
We provide a detailed elaboration on the examples in Section 2.2, which are widely encountered in practice. The unnormalized softmax logits, aka the logistic kernel function in (19), frequently appear in machine learning (Bishop and Nasrabadi, 2006), deep learning (Goodfellow et al., 2016), statistics (Keener, 2010; Wainwright, 2019), and statistical mechanics (Landau and Lifshitz, 2013). Similarly, the logistic regression function in (20) is a fundamental building block in the ML community.
C.1.1 Unnormalized Softmax Logits
Consider the following logistic kernel function:
[…]
We focus on the unbounded domain like $\mathbb{R}^{n}$, which is often the case in reality. For bounded domains (satisfying Assumption J.2) considered in this paper, the discussions are deferred to Section C.1.3. We will first show that the Hessian $\nabla^{2}f(\mathbf{x})$ is unbounded unless $\mathbf{w}=0$, indicating $f$ is not $L$-smooth for any constant $L$. Then, we demonstrate that generalized smoothness can effectively handle this problem. Finally, we give characterizations of $\ell*$-smoothness and $\ell$-smoothness, highlighting the edge of our formulation.
The gradient and Hessian of $f$ are given by
$\nabla f(\mathbf{x})=C\exp(\mathbf{w}^{\top}\mathbf{x}+b)\mathbf{w},\quad\nabla^{2}f(\mathbf{x})=C\exp(\mathbf{w}^{\top}\mathbf{x}+b)\mathbf{w}\mathbf{w}^{\top}.$
Then we have
$\left\|\nabla^{2}f(\mathbf{x})\right\|_{2}=\left|C\right|\exp(\mathbf{w}^{\top}\mathbf{x}+b)\left\|\mathbf{w}\right\|_{2}^{2}.$
Since the domain of $\mathbf{x}$ is unbounded, we conclude that $\left\|\nabla^{2}f(\mathbf{x})\right\|_{2}$ is unbounded unless $\mathbf{w}=0$. Hence, standard smoothness can not be applied here.
Next, we focus on $(L_{0},L_{1})$-smoothness. First, we argue that $(L_{0},L_{1})$-smoothness serves as a good remedy to the above failure case. Second, we reveal the gaps between $\ell$-smoothness and $\ell*$-smoothness.
$\ell$-smoothness
We have
$\left\|\nabla f(\mathbf{x})\right\|_{2}=\left|C\right|\exp(\mathbf{w}^{\top}\mathbf{x}+b)\left\|\mathbf{w}\right\|_{2},\quad\left\|\nabla^{2}f(\mathbf{x})\right\|_{2}=\left|C\right|\exp(\mathbf{w}^{\top}\mathbf{x}+b)\left\|\mathbf{w}\right\|_{2}^{2}.$
Thus,
$\left\|\nabla^{2}f(\mathbf{x})\right\|_{2}=\left|C\right|\left\|\mathbf{w}\right\|_{2}\cdot\left\|\nabla f(\mathbf{x})\right\|_{2},$
which implies $f\in\mathcal{F}_{\widehat{\ell}}(\left\|\cdot\right\|_{2})$ with $\widehat{\ell}(\alpha)=\left\|\mathbf{w}\right\|_{2}\alpha$. Obviously, $(L_{0},L_{1})$-smoothness can model the curvature of this function more appropriately.
$\ell*$-smoothness
Consider the norm pairs as $\left\|\cdot\right\|=\left\|\cdot\right\|_{1}$ and $\left\|\cdot\right\|_{*}=\left\|\cdot\right\|_{\infty}$. By Definition 2.1 in our manuscript, we have
$\sup_{\mathbf{h}\in\mathbb{R}^{n}}\frac{\left\|\nabla^{2}f(\mathbf{x})\mathbf{h}\right\|_{\infty}}{\left\|\mathbf{h}\right\|_{1}}=$ $\left|C\right|\exp(\mathbf{w}^{\top}\mathbf{x}+b)\sup_{\mathbf{h}\in\mathbb{R}^{n}}\frac{\left\|\mathbf{w}\mathbf{w}^{\top}\mathbf{h}\right\|_{\infty}}{\left\|\mathbf{h}\right\|_{1}}$
$=$ $\left|C\right|\exp(\mathbf{w}^{\top}\mathbf{x}+b)\left\|\mathbf{w}\right\|_{\infty}\sup_{\mathbf{h}\in\mathbb{R}^{n}}\frac{\left|\mathbf{w}^{\top}\mathbf{h}\right|}{\left\|\mathbf{h}\right\|_{1}}$
$=$ $\left\|\nabla f(\mathbf{x})\right\|_{\infty}\sup_{\mathbf{h}\in\Delta_{d}}\left|\mathbf{w}^{\top}\mathbf{h}\right|$
$=$ $\left\|\nabla f(\mathbf{x})\right\|_{\infty}\left\|\mathbf{w}\right\|_{\infty}.$
Thus, $f\in\mathcal{F}_{\widetilde{\ell}}(\left\|\cdot\right\|_{1})$ with $\widetilde{\ell}(\alpha)=\left\|\mathbf{w}\right\|_{\infty}\alpha$.
Comparison
Define the ratio between $\ell*$-smoothness and $\ell$-smoothness as
$\phi(\mathbf{w}):=\frac{\left\|\mathbf{w}\right\|_{\infty}}{\left\|\mathbf{w}\right\|_{2}}=\frac{\widetilde{\ell}(\alpha)}{\widehat{\ell}(\alpha)}\in\left[\frac{1}{\sqrt{n}},1\right].$
From above, we can assert that our formulation is strictly better than that in Li et al. (2023a). Taking $\mathbf{w}=\mathbf{1}_{n}$ satisfies the left side of the inequality. In this case, $f\in\mathcal{F}_{\widehat{\ell}}(\left\|\cdot\right\|_{2})$ with $\widehat{\ell}(\alpha)=\sqrt{n}\alpha$ and $f\in\mathcal{F}_{\widetilde{\ell}}(\left\|\cdot\right\|_{1})$ with $\widetilde{\ell}(\alpha)=\alpha$. Clearly, our definition surpasses $\ell$-smoothness by a dimension-dependent factor.
Even more practical case
This time, we consider $\mathbf{w}\in\mathbb{R}^{n}$ as a random variable uniformly distributed on a unit sphere. Since the slopes of the generalized smooth functions $\widehat{\ell}$$\widetilde{\ell}$ are now randomized, we measure the smoothness parameter in expectation. We clearly have $\left\|\mathbf{w}\right\|_{2}=1$ by definition. As for $\widetilde{\ell}$, it is well-known (Schechtman and Zinn, 2000; Boucheron et al., 2013) that
$\mathbb{E}[\left\|\mathbf{w}\right\|_{\infty}]\preceq\sqrt{\frac{\log n}{n}}.$
Hence, the (expected) ratio becomes
$\phi_{\mathbb{E}}(\mathbf{w}):=\frac{\mathbb{E}[\left\|\mathbf{w}\right\|_{\infty}]}{\mathbb{E}[\left\|\mathbf{w}\right\|_{2}]}=\mathbb{E}[\left\|\mathbf{w}\right\|_{\infty}]\preceq\sqrt{\frac{\log n}{n}}.$
Mathematically, we establish the bound $\widetilde{\ell}/\widehat{\ell}\preceq\sqrt{\log(n)/n}$ in expectation, demonstrating the quantitative advantage of $\ell*$-smoothness. This benefit becomes particularly pronounced when handling high-dimensional optimization problems, mirroring practical scenarios where our adaptive approach offers meaningful improvements over classical smoothness analysis.
C.1.2 Logistic Regression
Consider the loss function of logistic regression:
[…]
whose gradient and Hessian are given by
$\nabla f(\mathbf{x})=\frac{-C\mathbf{w}}{1+\exp(\mathbf{w}^{\top}\mathbf{x})},\quad\nabla^{2}f(\mathbf{x})=\frac{C\mathbf{w}\mathbf{w}^{\top}}{\exp(-\mathbf{w}^{\top}\mathbf{x})\left(1+\exp(\mathbf{w}^{\top}\mathbf{x})\right)^{2}}.$
Thus we have
$\left\|\nabla^{2}f(\mathbf{x})\right\|_{2}=$ $\frac{\left|C\right|\left\|\mathbf{w}\right\|_{2}^{2}}{\left(\exp(\frac{1}{2}\mathbf{w}^{\top}\mathbf{x})+\exp(-\frac{1}{2}\mathbf{w}^{\top}\mathbf{x})\right)^{2}}$
$\leq$ $\frac{\left|C\right|\left\|\mathbf{w}\right\|_{2}^{2}}{\left(2\sqrt{\exp\left(\frac{1}{2}\mathbf{w}^{\top}\mathbf{x}-\frac{1}{2}\mathbf{w}^{\top}\mathbf{x}\right)}\right)^{2}}=\frac{\left|C\right|\left\|\mathbf{w}\right\|_{2}^{2}}{4},$
implying that $f$ is $\left|C\right|\left\|\mathbf{w}\right\|_{2}^{2}/4$-smooth. This smoothness constant can become very large if $\left|C\right|\cdot\left\|\mathbf{w}\right\|_{2}\gg 1$. Therefore, it’s meaningful to seek an alternative. Let’s consider the generalized smoothness notion as in Section C.1.1.
$\ell$-smoothness
We have
$\left\|\nabla f(\mathbf{x})\right\|_{2}=\frac{\left|C\right|\left\|\mathbf{w}\right\|_{2}}{1+\exp(\mathbf{w}^{\top}\mathbf{x})},\quad\left\|\nabla^{2}f(\mathbf{x})\right\|_{2}=\frac{\left|C\right|\left\|\mathbf{w}\right\|_{2}^{2}}{\exp(-\mathbf{w}^{\top}\mathbf{x})\left(1+\exp(\mathbf{w}^{\top}\mathbf{x})\right)^{2}}.$
Thus,
$\frac{\left\|\nabla^{2}f(\mathbf{x})\right\|_{2}}{\left\|\nabla f(\mathbf{x})\right\|_{2}}=$ $\frac{(1+\exp(\mathbf{w}^{\top}\mathbf{x}))\left\|\mathbf{w}\right\|_{2}}{\exp(-\mathbf{w}^{\top}\mathbf{x})\left(1+\exp(\mathbf{w}^{\top}\mathbf{x})\right)^{2}}$
$=$ $\frac{\left\|\mathbf{w}\right\|_{2}}{1+\exp(-\mathbf{w}^{\top}\mathbf{x})}\leq\left\|\mathbf{w}\right\|_{2}.$
Since $\left\|\nabla^{2}f(\mathbf{x})\right\|_{2}\leq\left\|\mathbf{w}\right\|_{2}\cdot\left\|\nabla f(\mathbf{x})\right\|_{2},\forall\mathbf{x}\in\mathbb{R}^{n}$, we conclude that $f\in\mathcal{F}_{\widehat{\ell}}(\left\|\cdot\right\|_{2})$ with $\widehat{\ell}(\alpha)=\left\|\mathbf{w}\right\|_{2}\alpha$ (Chen et al., 2023b). Clearly, when $\left\|\mathbf{w}\right\|_{2}\gg 1$, generalized smooth is more appropriate because it improves upon the standard smooth by a factor of $\left\|\mathbf{w}\right\|_{2}$.
$\ell*$-smoothness
We follow a similar procedure as above:
$\left\|\nabla f(\mathbf{x})\right\|_{\infty}=\frac{\left|C\right|\left\|\mathbf{w}\right\|_{\infty}}{1+\exp(\mathbf{w}^{\top}\mathbf{x})},$
and compute
$\sup_{\mathbf{h}\in\mathbb{R}^{n}}\frac{\left\|\nabla^{2}f(\mathbf{x})\mathbf{h}\right\|_{\infty}}{\left\|\mathbf{h}\right\|_{1}}=$ $\frac{\left|C\right|}{\exp(-\mathbf{w}^{\top}\mathbf{x})\left(1+\exp(\mathbf{w}^{\top}\mathbf{x})\right)^{2}}\sup_{\mathbf{h}\in\mathbb{R}^{n}}\frac{\left\|\mathbf{w}\mathbf{w}^{\top}\mathbf{h}\right\|_{\infty}}{\left\|\mathbf{h}\right\|_{1}}$
$=$ $\frac{\left|C\right|\left\|\mathbf{w}\right\|_{\infty}}{\exp(-\mathbf{w}^{\top}\mathbf{x})\left(1+\exp(\mathbf{w}^{\top}\mathbf{x})\right)^{2}}\sup_{\mathbf{h}\in\mathbb{R}^{n}}\frac{\left|\mathbf{w}^{\top}\mathbf{h}\right|}{\left\|\mathbf{h}\right\|_{1}}$
$=$ $\frac{\left\|\nabla f(\mathbf{x})\right\|_{\infty}}{1+\exp(-\mathbf{w}^{\top}\mathbf{x})}\sup_{\mathbf{h}\in\Delta_{d}}\left|\mathbf{w}^{\top}\mathbf{h}\right|$
$\leq$ $\left\|\nabla f(\mathbf{x})\right\|_{\infty}\left\|\mathbf{w}\right\|_{\infty}.$
Therefore, we conclude that $f\in\mathcal{F}_{\widetilde{\ell}}(\left\|\cdot\right\|_{1})$ with $\widetilde{\ell}(\alpha)=\left\|\mathbf{w}\right\|_{\infty}\alpha$.
Comparison
Similarly, we can define the ratio function $\phi(\mathbf{w})$ as in Section C.1.1, and the same conclusions can be drawn: our definition surpasses $\ell$-smoothness by a dimension-dependent factor of $1/\sqrt{n}$.
Logistic regression on a high-dimensional sphere
Similar to Section C.1.1, we now consider logistic regression on a high-dimensional sphere, with radius $R\gg 1$. In this case, $\left\|\mathbf{w}\right\|_{2}=R\gg 1$, and $\mathbf{w}$ is uniformly distributed on the $R$-radius sphere. Note that $\mathbf{w}/R$ can be seen as a random variable uniformly distributed on the unit sphere, then we we also have $\phi_{\mathbb{E}}(\mathbf{w}):=\frac{\mathbb{E}[\left\|\mathbf{w}\right\|_{\infty}]}{\mathbb{E}[\left\|\mathbf{w}\right\|_{2}]}=\mathbb{E}[\left\|\mathbf{w}\right\|_{\infty}]\preceq\sqrt{\frac{\log n}{n}}$. Again, we verify the effectiveness and benefits of our $\ell*$-smoothness in this practical scenario.
C.1.3 The cases for bounded domains
The examples and justification in Sections C.1.1 and C.1.2 targets unbounded domains like $\mathbb{R}^{n}$. Under Assumption J.2, the previous arguments are still valid and we emphasize that the advantage of $\ell*$-smoothness does not stem from unboundedness of the domain.
Unnormalized Softmax Logits
For $f(\mathbf{x})=C\exp(\mathbf{w}^{\top}\mathbf{x}+b),\mathbf{w},\mathbf{x}\in Y:=\{\mathbf{y}\in\mathbb{R}^{n}|\|\mathbf{y}\|_{2}\leq R\}$. Then the standard smoothness constant is $L=\max_{\mathbf{x}}\|\nabla^{2}f(\mathbf{x})\|_{2}\leq\left|C\right|\exp(R^{2}+b)R^{2}$. Unfortunately, it exhibits an exponential dependence on $R$, making traditional smoothness impractical. For instance, when $\left|C\right|=b=R=5$, we get $L>10^{15}$, which is clearly meaningless and necessitates a better formulation. One can verify that we still have $f\in\mathcal{F}_{\widehat{\ell}}(\left\|\cdot\right\|_{2})$ with $\widehat{\ell}(\alpha)=\|\mathbf{w}\|_{2}\alpha$ and $f\in\mathcal{F}_{\widetilde{\ell}}(\left\|\cdot\right\|_{1})$ with $\widetilde{\ell}(\alpha)=\|\mathbf{w}\|_{\infty}\alpha$.
Logistic Regression
For $f(\mathbf{x})=4C\log(1+\exp(-\mathbf{w}^{\top}\mathbf{x})),\mathbf{w},\mathbf{x}\in Y$, we have $L=\max_{\mathbf{x}}\allowbreak\|\nabla^{2}f(\mathbf{x})\|_{2}\leq\left|C\right|\|\mathbf{w}\|_{2}^{2}$ and $f\in\mathcal{F}_{\widehat{\ell}}(\left\|\cdot\right\|_{2}),\widehat{\ell}(\alpha)=\|\mathbf{w}\|_{2}\alpha;f\in\mathcal{F}_{\widetilde{\ell}}(\left\|\cdot\right\|_{1}),\widetilde{\ell}(\alpha)=\|\mathbf{w}\|_{\infty}\alpha$. The example remains meaningful when the domain and $\left|C\right|$ is relatively large.
Main takeaways
In general, it makes sense to consider $\ell*$-smoothness even if standard or global smoothness holds, and for some functions like unnormalized softmax logits or logistic regression, $\ell*$-smoothness provably improves upon $\ell$-smoothness with dimension-dependent gains, regardless of the domain being bounded or not.
C.2 Theoretical Justifications: Constant Link Functions
Consider the function $h(\mathbf{x})=\mathbf{x}^{\top}(\mathbf{1}_{n}-\mathbf{e}_{1})(\mathbf{1}_{n}-\mathbf{e}_{1})^{\top}\mathbf{x}/2,\mathbf{x}\in\Delta_{n}$, where $\Delta_{n}=\{\mathbf{x}\geq\mathbf{0},\mathbf{1}_{n}^{\top}\mathbf{x}=1|\mathbf{x}\in\mathbb{R}^{n}\}$ is the (n1)(n-1)-dimensional simplex. We have the following proposition.
Proposition C.1.
(i) $h\in\mathcal{F}_{\widehat{\ell}}(\left\|\cdot\right\|_{2})$ with $\widehat{\ell}(\cdot)\equiv n-1$; (ii) $h\in\mathcal{F}_{\widetilde{\ell}}(\left\|\cdot\right\|_{1})$ with $\widetilde{\ell}(\cdot)\equiv 1$.
Proof.
Denote $(\mathbf{1}_{n}-\mathbf{e}_{1})(\mathbf{1}_{n}-\mathbf{e}_{1})^{\top}$ by $A$. For the quadratic function $h(\mathbf{x})=\frac{1}{2}\mathbf{x}^{\top}A\mathbf{x}$ where $A=A^{\top}$, we have
$\nabla h(\mathbf{x})=A\mathbf{x},\quad\nabla^{2}h(\mathbf{x})=A.$ (21)
Since $A=(\mathbf{1}_{n}-\mathbf{e}_{1})(\mathbf{1}_{n}-\mathbf{e}_{1})^{\top}$ is a rank-1 matrix, we have
$\left\|\nabla^{2}h(\mathbf{x})\right\|_{2}=\left\|\mathbf{1}_{n}-\mathbf{e}_{1}\right\|_{2}\left\|\mathbf{1}_{n}-\mathbf{e}_{1}\right\|_{2}=\sqrt{n-1}\cdot\sqrt{n-1}=n-1.$ (22)
[…]
$\forall\mathbf{u}\in\Delta_{n},\left\|\nabla^{2}h(\mathbf{x})\mathbf{u}\right\|_{\infty}\leq\widetilde{\ell}(\left\|\nabla h(\mathbf{x})\right\|_{\infty})\left\|\mathbf{u}\right\|_{1}$ (23)
$\Longleftrightarrow$ $\sup_{\mathbf{u}\in\Delta_{n}}\left\|\nabla^{2}h(\mathbf{x})\mathbf{u}\right\|_{\infty}\leq\widetilde{\ell}(\left\|\nabla h(\mathbf{x})\right\|_{\infty}).$
We proceed by computing
$\sup_{\mathbf{u}\in\Delta_{n}}\left\|A\mathbf{u}\right\|_{\infty}=\sup_{\mathbf{u}\in\Delta_{n}}\left\|(\mathbf{1}_{n}-\mathbf{e}_{1})[(\mathbf{1}_{n}-\mathbf{e}_{1})^{\top}\mathbf{u}]\right\|_{\infty}$ (24)
$=$ $\sup_{\mathbf{u}\in\Delta_{n}}(\mathbf{1}_{n}-\mathbf{e}_{1})^{\top}\mathbf{u}=\sup_{\mathbf{u}\in\Delta_{n}}\sum_{i=2}^{n}u_{i}=1,$
where we make use of $\left\|\mathbf{1}_{n}-\mathbf{e}_{1}\right\|_{\infty}=1$. Therefore, we conclude that $\widetilde{\ell}(\cdot)\equiv 1$. ∎
Under $\ell$-smoothness, the function $h$ inevitably has a smoothness parameter that scales linearly to the dimension $n$. Whereas, under $\ell*$-smoothness, the dependence of the dimension no longer exists. The additional factor n1n-1 will inevitably incur a dimension-dependent convergence rate, which will be stated explicitly in Section C.4.
C.3 Empirical Justifications
(b) Comparison of $\mathcal{F}_{\widehat{\ell}}(\left\|\cdot\right\|_{2})$ and $h\in\mathcal{F}_{\widetilde{\ell}}(\left\|\cdot\right\|_{1})$ generalized smooth functions class for different values of $n$.
We focus on the following example function:
$h(\mathbf{x})=\frac{\left(\sum_{i=1}^{n}x_{i}^{2}\right)^{2}}{4}+\prod_{i=1}^{n}x_{i}^{2}+\sum_{i=1}^{n}e^{5x_{i}},\quad\forall\mathbf{x}\in\Delta_{n}.$ (25)
For the Euclidean setup $h\in\mathcal{F}_{\widehat{\ell}}(\left\|\cdot\right\|_{2})$ as well as the non-Euclidean setup $h\in\mathcal{F}_{\widetilde{\ell}}(\left\|\cdot\right\|_{1})$, the smoothness can be modeled by
$\left\|\nabla^{2}h(\mathbf{x})\right\|_{2}\leq\widehat{\ell}(\left\|\nabla h(\mathbf{x})\right\|_{2}),\text{ and }\sup_{\mathbf{u}\in\Delta_{n}}\left\|\nabla^{2}h(\mathbf{x})\mathbf{u}\right\|_{\infty}\leq\widetilde{\ell}(\left\|\nabla h(\mathbf{x})\right\|_{\infty}),$ (26)
[…]
$\widehat{\ell}(\alpha)=\widehat{L}_{0}+\widehat{L}_{1}\alpha,\quad\widetilde{\ell}(\alpha)=\widetilde{L}_{0}+\widetilde{L}_{1}\alpha.$ (27)
We find this characterization reasonable based on observations from Figure 5(b): when $n>20$, the scattered points align into a straight line. To this end, we fit the values of $\widehat{L}_{0}$, $\widehat{L}_{1}$, $\widetilde{L}_{0}$, and $\widetilde{L}_{1}$ using $500$ randomly sampled points from the simplex $\Delta_{n}$.
Based on the modeling in (27), we can see that the slopes $\widehat{L}_{1}$ and $\widetilde{L}_{1}$ determine the magnitude of smoothness along the optimization trajectory (e.g., for some constant $0<G<\infty$, $\widehat{\ell}(G)=\widehat{L}_{0}+\widehat{L}_{1}G$), to a great extent. To clarify, note that the definition of $G$ in (5) has a closed-form solution:
[…]
$=$ $F\widetilde{L}_{1}+\sqrt{2F\widetilde{L}_{0}+F^{2}\widetilde{L}_{1}^{2}}\leq 2F\widetilde{L}_{1}+\sqrt{2F\widetilde{L}_{0}},$
where $f(\mathbf{x}_{0})-f^{*}$ is denoted by $F$. The above relation implies that $G$ grows linearly w.r.t. $\widetilde{L}_{1}$ and $\sqrt{\widetilde{L}_{0}}$. As a result, by (5), we compute the following effective smoothness parameter:
$L=\widetilde{\ell}(2G)=\widehat{L}_{0}+2\widehat{L}_{1}G\leq\widehat{L}_{0}+4F\widehat{L}_{1}^{2}+2F\widehat{L}_{1}\sqrt{2F\widetilde{L}_{0}}.$ (29)
Therefore, we conclude that $\widetilde{L}_{1}$ is the dominant term affecting the smoothness parameter $L$.
Due to the reasoning above, we concentrate on the slopes $\widetilde{L}_{1},\widehat{L}_{1}$ to analyze and compare the characteristics of the link functions $\widehat{\ell}$ and $\widetilde{\ell}$. To further explore this, we examine the ratio $\widetilde{L}_{1}/\widehat{L}_{1}$ for various values of $n$. Specifically, we vary $n$ from $6$ to $198$ in increments of $3$ and fit a function of the form $g(n)=a\cdot n^{-b},a,b>0$. The results, shown in Figure 5, reveal that the gap between $\widetilde{\ell}$ and $\widehat{\ell}$ is approximately at the order of $O(n^{-0.4})$.
Figure 5: The slope ratio $\widetilde{L}_{1}/\widehat{L}_{1}$ w.r.t. dimension $n$ and the fitted curve.
C.4 Comparison on the Convergence Rates
[…]
$\frac{\widehat{\ell}(\left\|\nabla f(\mathbf{x}_{0})\right\|_{2})\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|_{2}^{2}}{2T}\leq\frac{\widehat{\ell}(\left\|\nabla f(\mathbf{x}_{0})\right\|_{2})}{2T}\triangleq R_{1}(T),$ (30)
where we use $\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|_{2}^{2}\leq\sup_{\mathbf{x},\mathbf{y}\in\Delta_{n}}\left\|\mathbf{x}-\mathbf{y}\right\|_{2}^{2}=1$. For mirror descent, as established in Theorem 3.5, the convergence rate is given by
$\frac{\widetilde{\ell}(G)B(\mathbf{x}_{*},\mathbf{x}_{0})}{T}\leq\frac{\widetilde{\ell}(G)\ln n}{T}\triangleq R_{2}(T),$ (31)
[…]
$\frac{R_{2}(T)}{R_{1}(T)}=\frac{\widetilde{\ell}(G)\ln n}{\widehat{\ell}(\left\|\nabla f(\mathbf{x}_{0})\right\|_{2})}=\frac{\ln n}{n-1},$ (32)
suggesting that mirror descent defined in (6) provably improves the convergence rate by a factor of $n/\ln n$. The effect will become prominent when the dimension is large.
[…]
$\frac{R_{2}(T)}{R_{1}(T)}=\frac{\widetilde{\ell}(G)\ln n}{\widehat{\ell}(\left\|\nabla f(\mathbf{x}_{0})\right\|_{2})}\simeq\frac{\widetilde{\ell}(\left\|\nabla f(\mathbf{x}_{0})\right\|_{\infty})\ln n}{\widehat{\ell}(\left\|\nabla f(\mathbf{x}_{0})\right\|_{2})}\simeq\frac{\widetilde{L}_{1}\ln n}{\widehat{L}_{1}}\simeq\frac{\ln n}{n^{0.4}},$ (33)
underlining the advantage of mirror descent over gradient descent by reducing the dependency on the dimensionality factor. Obviously, the justifications for non-constant link functions in Section C.1 also follow from the above discussions, which we omit for simplicity.
Appendix D Properties of Generalized Smooth Function Classes
In this section, we demonstrate some basic properties of the function class $\mathcal{F}_{\ell}(\left\|\cdot\right\|)$ as well as $\mathcal{F}_{\ell,r}(\left\|\cdot\right\|)$. Equation 34 is a generalized version of Lemma A.4 of Li et al. (2023a), which only supports $f\in\mathcal{F}_{\ell}(\left\|\cdot\right\|_{2})$. Since the original derivations did not utilize the inherent property of the Euclidean norm, i.e., $\langle\cdot,\cdot\rangle=\left\|\cdot\right\|_{2}^{2}$, we can adapt its proof to accommodate an arbitrary norm $\left\|\cdot\right\|$.
Lemma D.1.
Let $f\in\mathcal{F}_{\ell}(\left\|\cdot\right\|)$. For any $\widetilde{\mathbf{x}},\mathbf{x}\in\mathcal{X}$, denote $\mathbf{x}(s):=s\widetilde{\mathbf{x}}+(1-s)\mathbf{x}$. If $\mathbf{x}(s)\in\mathcal{X}$ holds for all $0\leq s\leq 1$ and $\left\|\widetilde{\mathbf{x}}-\mathbf{x}\right\|\leq\frac{G}{\ell(\left\|\nabla f(\mathbf{x})\right\|_{*}+G)}$ for any $G\in\mathbb{R}_{++}$, then we have
$\max_{0\leq s\leq 1}\left\|\nabla f(\mathbf{x}(s))\right\|_{*}\leq\left\|\nabla f(\mathbf{x})\right\|_{*}+G.$ (34)
Proof.
First, define $g(s):=\left\|\nabla f(\mathbf{x}(s))\right\|_{*}$ for $0\leq s\leq 1$. Since $\mathbf{x}(s)\in\mathcal{X}$, $\mathbf{g}(s)$ is differentiable almost everywhere by Definition 2.1. Hence, the following relation holds almost everywhere over $(0,1)$:
$g^{\prime}(s)=$ $\lim_{t\to s}\frac{g(t)-g(s)}{t-s}\leq\lim_{t\to s}\frac{\left\|\nabla f(\mathbf{x}(t))-\nabla f(\mathbf{x}(s))\right\|_{*}}{t-s}$ (35)
$=$ $\left\|\lim_{t\to s}\frac{\nabla f(\mathbf{x}(t))-\nabla f(\mathbf{x}(s))}{t-s}\right\|_{*}=\left\|\nabla^{2}f(\mathbf{x}(s))(\widetilde{\mathbf{x}}-\mathbf{x})\right\|_{*}\leq\ell(g(s))\left\|\widetilde{\mathbf{x}}-\mathbf{x}\right\|,$
where the first inequality uses the triangle inequality and the last inequality uses Definition 2.1. Next, we define $h(u):=\ell(u)\left\|\widetilde{\mathbf{x}}-\mathbf{x}\right\|$ and $k(u):=\int_{0}^{u}1/h(v)\mathrm{d}v$. We immediately conclude that $g^{\prime}(s)\leq h(g(s))$ holds almost everywhere over $(0,1)$, which satisfies the condition of a generalized version of Grönwall’s inequality (Li et al., 2023a, Lemma A.3). Hence, we have
$k(\left\|\nabla f(\widetilde{\mathbf{x}})\right\|_{*})=k(g(1))\leq k(g(0))+1=k(\left\|\nabla f(\mathbf{x}\right\|_{*}))+1.$ (36)
Based on this result, we can derive
$k(\left\|\nabla f(\widetilde{\mathbf{x}})\right\|_{*})\left\|\widetilde{\mathbf{x}}-\mathbf{x}\right\|\leq$ $k(g(0))\left\|\widetilde{\mathbf{x}}-\mathbf{x}\right\|+\left\|\widetilde{\mathbf{x}}-\mathbf{x}\right\|$ (37)
$\leq$ $\int_{0}^{g(0)}\frac{\left\|\widetilde{\mathbf{x}}-\mathbf{x}\right\|}{\ell(u)\left\|\widetilde{\mathbf{x}}-\mathbf{x}\right\|}\mathrm{d}u+\frac{G}{\ell(\left\|\nabla f(\mathbf{x})\right\|_{*}+G)}$
$\leq$ $\int_{0}^{\left\|\nabla f(\mathbf{x})\right\|_{*}}\frac{1}{\ell(u)}\mathrm{d}u+\int_{\left\|\nabla f(\mathbf{x})\right\|_{*}}^{\left\|\nabla f(\mathbf{x})\right\|_{*}+G}\frac{1}{\ell(u)}\mathrm{d}u$
$=$ $k(\left\|\nabla f(\mathbf{x})\right\|_{*}+G)\left\|\widetilde{\mathbf{x}}-\mathbf{x}\right\|,$
where the second inequality uses the bound for $\left\|\widetilde{\mathbf{x}}-\mathbf{x}\right\|$ and the last inequality use the non-decreasing property of $\ell$. By the definition of $k$, it is obvious to see that $k$ is increasing. Then we can conclude that
$\left\|\nabla f(\widetilde{\mathbf{x}})\right\|_{*}\leq\left\|\nabla f(\mathbf{x})\right\|_{*}+G.$ (38)
Note that RHS of (38) involves the anchor point $\mathbf{x}$. Under the given conditions, the closed line segment between $\widetilde{\mathbf{x}}$ and $\mathbf{x}$ is contained in $\mathcal{X}$. Then, for any $0\leq\tilde{s}\leq 1$, we also have $\left\{\mathbf{x}(s)|0\leq s\leq\tilde{s}\right\}\subseteq\mathcal{X}$ and $\left\|\mathbf{x}(\tilde{s})-\mathbf{x}\right\|\leq\frac{G}{\ell(\left\|\nabla f(\mathbf{x})\right\|_{*}+G)}$. Therefore, we can invoke (38) to show that
$\left\|\nabla f(\mathbf{x}(\tilde{s}))\right\|_{*}\leq\left\|\nabla f(\mathbf{x})\right\|_{*}+G,\quad\forall\tilde{s}\in[0,1].$ (39)
Maximizing over $0\leq\tilde{s}\leq 1$ yields the desired result. ∎
In the following, we provide the omitted proof of Proposition 2.6.
Proof of Proposition 2.6.
The proof is very similar to that of Li et al. (2023a, Proposition 3.2), except that the derivations based on $\left\|\cdot\right\|_{2}$ should be replaced by an arbitrary norm $\left\|\cdot\right\|$.
Part 1. $\mathcal{F}_{\ell,r}(\left\|\cdot\right\|)\subseteq\mathcal{F}_{\ell}(\left\|\cdot\right\|)$.
Given an $\mathbf{x}\in\mathcal{X}$ where $\nabla^{2}f(\mathbf{x})$ exists, according to Definition 2.4, the following relation holds
$\left\|\nabla f(\mathbf{x}+s\mathbf{h})-\nabla f(\mathbf{x})\right\|_{*}\leq\ell(\left\|\nabla f(\mathbf{x})\right\|_{*})\cdot s\left\|\mathbf{h}\right\|$ (40)
for any $\mathbf{h}\in\mathcal{X}$ and any $s\in\mathbb{R}_{++}$ satisfying $s\left\|\mathbf{h}\right\|\leq r(\left\|\nabla f(\mathbf{x})\right\|_{*})$. Then, based on the continuity of the dual norm, we deduce that
$\left\|\nabla^{2}f(\mathbf{x})\mathbf{h}\right\|_{*}=$ $\left\|\lim_{s\to 0}\frac{\nabla f(\mathbf{x}+s\mathbf{h})-\nabla f(\mathbf{x})}{s}\right\|_{*}$ (41)
$=$ $\lim_{s\to 0}\left\|\frac{\nabla f(\mathbf{x}+s\mathbf{h})-\nabla f(\mathbf{x})}{s}\right\|_{*}\leq\ell(\left\|\nabla f(\mathbf{x})\right\|_{*})\cdot\left\|\mathbf{h}\right\|.$
Next, we prove that $\nabla^{2}f(\mathbf{x})$ exists almost everywhere. We briefly list the following two key steps, the detailed derivations can be found in Li et al. (2023a, Proposition 3.2).
  1. 1.
    Invoke Rademacher’s Theorem (Evans, 2018) to show that $f$ is twice differentiable almost everywhere within the ball $\mathcal{B}(\mathbf{x},r(\left\|\nabla f(\mathbf{x})\right\|_{*}))$.
  2. 2.
    Leverage the discretization as well as the covering technique to show that $f$ is twice differentiable almost everywhere within the domain $\mathcal{X}$.
[…]
First, we show that $\mathcal{B}(\mathbf{x},\widetilde{r}(\left\|\nabla f(\mathbf{x})\right\|_{*}))\subseteq\mathcal{X}$ for any $\mathbf{x}\in\mathcal{X}$ by contradiction. For any $\widetilde{\mathbf{x}}\in\mathcal{B}(\mathbf{x},\widetilde{r}(\left\|\nabla f(\mathbf{x})\right\|_{*}))$, suppose $\widetilde{\mathbf{x}}\notin\mathcal{X}$. Define
$\mathbf{x}(s):=s\widetilde{\mathbf{x}}+(1-s)\mathbf{x},\ 0\leq s\leq 1\text{ and }s_{b}:=\inf\left\{s|\mathbf{x}(s)\notin\mathcal{X}\right\}.$ (42)
By the definition of $s_{b}$ and Assumption 2.5, we have $\lim_{s\to s_{b}}f(\mathbf{x}(s))=\infty$ and $\mathbf{x}(s)\in\mathcal{X}$ for $0\leq s<s_{b}$. Then we have
$\forall s\in[0,s_{b}):f(\mathbf{x}(s))=$ $f(\mathbf{x})+\int_{0}^{s}\langle\nabla f(\mathbf{x}(u)),\widetilde{\mathbf{x}}-\mathbf{x}\rangle{\mathrm{d}}u$ (43)
$\leq$ $f(\mathbf{x})+s\cdot\max_{0\leq u\leq s}\left\|\nabla f(\mathbf{x}(u))\right\|_{*}\cdot\left\|\widetilde{\mathbf{x}}-\mathbf{x}\right\|$
$\leq$ $f(\mathbf{x})+\left(\left\|\nabla f(\mathbf{x})\right\|_{*}+G\right)\left\|\widetilde{\mathbf{x}}-\mathbf{x}\right\|<\infty,$
where the last inequality uses Equation 34 because the following condition is satisfied
$\left\|\widetilde{\mathbf{x}}-\mathbf{x}\right\|\leq\widetilde{r}(\left\|\nabla f(\mathbf{x})\right\|_{*})=G/\ell(\left\|\nabla f(\mathbf{x})\right\|_{*}+G)$ (44)
by definition. This contradicts $\lim_{s\to s_{b}}f(\mathbf{x}(s))=\infty$. Thus we have $\widetilde{\mathbf{x}}\in\mathcal{X}$ and consequently, $\mathcal{B}(\mathbf{x},\widetilde{r}(\left\|\nabla f(\mathbf{x})\right\|_{*}))\subseteq\mathcal{X}$. Next, we prove (2) in Definition 2.4. For any $\mathbf{x}_{1},\mathbf{x}_{2}\in\mathcal{B}(\mathbf{x},\widetilde{r}(\left\|\nabla f(\mathbf{x})\right\|_{*}))$, define $\mathbf{x}_{1,2}(s):=s\mathbf{x}_{1}+(1-s)\mathbf{x}_{2}$ for $0\leq s\leq 1$. Since $\mathbf{x}_{1,2}(s)\in\mathcal{B}(\mathbf{x},\widetilde{r}(\left\|\nabla f(\mathbf{x})\right\|_{*}))$, we obtain
$\left\|\nabla f(\mathbf{x}_{1})-\nabla f(\mathbf{x}_{2})\right\|_{*}=$ $\left\|\int_{0}^{1}\nabla^{2}f\left(\mathbf{x}_{1,2}(s)\right)(\mathbf{x}_{1}-\mathbf{x}_{2}){\mathrm{d}}s\right\|_{*}$ (45)
$\leq$ $\int_{0}^{1}\ell\left(\left\|\nabla f\left(\mathbf{x}_{1,2}(s)\right)\right\|_{*}\right)\cdot\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|{\mathrm{d}}s$
$\leq$ $\max_{0\leq s\leq 1}\ell\left(\left\|\nabla f\left(\mathbf{x}_{1,2}(s)\right)\right\|_{*}\right)\cdot\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|$
$\leq$ $\ell\left(\left\|\nabla f(\mathbf{x})\right\|_{*}+G\right)\cdot\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|,$
where the first inequality uses Definition 2.1 and the last inequality uses Equation 34 and the non-decreasing property of $\ell$. ∎
We present the following effective or local smoothness property derived under generalized smoothness. Lemma 2.7 is extremely useful as it closes the gap between generalized smoothness and classical smoothness.
Proof of Lemma 2.7.
To prove this lemma, we adapt the existing result for $f\in\mathcal{F}_{\ell,r}(\left\|\cdot\right\|_{2})$ (Li et al., 2023a, Lemma 3.3). Specifically, we establish three analogous statements for $f\in\mathcal{F}_{\ell,r}(\left\|\cdot\right\|)$ and then leverage the equivalence between generalized smooth function classes to conclude that these statements also hold for $f\in\mathcal{F}_{\ell}(\left\|\cdot\right\|)$. By Definition 2.4, we have $\ell(\left\|\nabla f(\mathbf{x})\right\|_{*})\leq\ell(G)$ and $r(G)\leq r(\left\|\nabla f(\mathbf{x})\right\|_{*})$. Hence, it follows that (i) $\mathcal{B}(\mathbf{x},r(G))\subseteq\mathcal{B}(\mathbf{x},r(\left\|\nabla f(\mathbf{x})\right\|_{*}))\subseteq\mathcal{X}$; and (ii) $\forall\mathbf{x}_{1},\mathbf{x}_{2}\in\mathcal{B}(\mathbf{x},r(G))$:
$\left\|\nabla f(\mathbf{x}_{1})-\nabla f(\mathbf{x}_{2})\right\|_{*}\leq\ell(\left\|\nabla f(\mathbf{x})\right\|_{*})\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|\leq\ell(G)\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|.$ (46)
Statements (i) and (ii) are similar to statements 1 and 2 in the lemma. Now we proceed to the quadratic upper bound. Define $\mathbf{x}(s)=s\mathbf{x}_{1}+(1-s)\mathbf{x}_{2}$ for $0\leq s\leq 1$. It’s clear that $\mathbf{x}(s),\mathbf{x}_{2}\in\mathcal{B}(\mathbf{x},r(G))$, so according to (46), we have
$\left\|\nabla f(\mathbf{x}(s))-\nabla f(\mathbf{x}_{2})\right\|_{*}\leq\ell(G)\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|.$ (47)
Then we can derive a local quadratic upper bound (denoted by (iii)) as follows
$f(\mathbf{x}_{1})-f(\mathbf{x}_{2})=$ $\int_{0}^{1}\langle\nabla f(\mathbf{x}(s)),\mathbf{x}_{1}-\mathbf{x}_{2}\rangle\mathrm{d}s$ (48)
$=$ $\langle\nabla f(\mathbf{x}_{2}),\mathbf{x}_{1}-\mathbf{x}_{2}\rangle+\int_{0}^{1}\langle\nabla f(\mathbf{x}(s))-\nabla f(\mathbf{x}_{2}),\mathbf{x}_{1}-\mathbf{x}_{2}\rangle\mathrm{d}s$
$\leq$ $\langle\nabla f(\mathbf{x}_{2}),\mathbf{x}_{1}-\mathbf{x}_{2}\rangle+\int_{0}^{1}\left\|\nabla f(\mathbf{x}(s))-\nabla f(\mathbf{x}_{2})\right\|_{*}\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|\mathrm{d}s$
$\leq$ $\langle\nabla f(\mathbf{x}_{2}),\mathbf{x}_{1}-\mathbf{x}_{2}\rangle+\frac{\ell(G)}{2}\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|^{2}.$
At last, we apply the equivalence of generalized smooth function classes to show that the same derivations also hold for $f\in\mathcal{F}_{\ell}(\left\|\cdot\right\|)$. In this case, it suffices to set $L=\ell(2G)$ and $r(G)=G/L$ according to Proposition 2.6. Then the statements (i), (ii), and (iii) can be easily transformed into the ones in Lemma 2.7, which concludes the proof. ∎
The following lemma is introduced out of convenience, for the sake of frequent usages in the next few sections.
Lemma D.2.
Under Assumption 3.3, let $\left\|\nabla f(\mathbf{x})\right\|_{*}\leq G\in\mathbb{R}_{+}$ for any given $\mathbf{x}\in\mathcal{X}$. If $0<\eta\leq\gamma/\ell(2G)$ for some $\gamma\in(0,1)$, then for any $\mathbf{x}_{1},\mathbf{x}_{2}\in\mathcal{B}(\mathbf{x},G/\ell(2G))$ and any $\mathbf{x}_{3},\mathbf{x}_{4}\in\mathcal{X}$,
$\eta\langle\nabla f(\mathbf{x}_{1})-\nabla f(\mathbf{x}_{2}),\mathbf{x}_{3}-\mathbf{x}_{4}\rangle\leq\frac{\gamma^{2}}{2\beta}\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|^{2}+\frac{\beta}{2}\left\|\mathbf{x}_{3}-\mathbf{x}_{4}\right\|^{2},$ (49)
holds for any $\beta\in\mathbb{R}_{++}$.
Proof.
Since $\left\|\nabla f(\mathbf{x})\right\|_{*}\leq G$ and $\mathbf{x}_{1},\mathbf{x}_{2}\in\mathcal{B}(\mathbf{x},G/\ell(2G))$, by Lemma 2.7, we have
$\left\|\nabla f(\mathbf{x}_{1})-\nabla f(\mathbf{x}_{2})\right\|_{*}\leq\ell(2G)\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|.$ (50)
According to Hölder’s Inequality, we have
$\forall\beta\in\mathbb{R}_{++}:\eta\langle\nabla f(\mathbf{x}_{1})-\nabla f(\mathbf{x}_{2}),\mathbf{x}_{3}-\mathbf{x}_{4}\rangle\leq\frac{\eta^{2}}{2\beta}\left\|\nabla f(\mathbf{x}_{1})-\nabla f(\mathbf{x}_{2})\right\|_{*}^{2}+\frac{\beta}{2}\left\|\mathbf{x}_{3}-\mathbf{x}_{4}\right\|^{2}.$ (51)
Combining these two inequalities with the condition that $0<\eta\leq\gamma/\ell(2G)$ completes the proof. ∎
Lastly, we provide the omitted proof of Lemma 3.4 in Section 3.2Lemma 3.4 is also a generalized version of the reversed Polyak-Łojasiewicz Inequality in Li et al. (2023a, Lemma 3.5). Since their derivations rely on the inherent properties of the Euclidean norm, we need to take a different approach to prove this lemma.
Proof of Lemma 3.4.
For any $\mathbf{x}\in\mathcal{X}$ and any $\mathbf{h}$ satisfying $\left\|\mathbf{h}\right\|\leq\left\|\nabla f(\mathbf{x})\right\|_{*}/\ell(2\left\|\nabla f(\mathbf{x})\right\|_{*})$, by Lemma 2.7, the following relation holds:
$\langle\nabla f(\mathbf{x}),\mathbf{h}\rangle-\frac{\ell(2\left\|\nabla f(\mathbf{x})\right\|_{*})}{2}\left\|\mathbf{h}\right\|^{2}\leq f(\mathbf{x})-f(\mathbf{x}-\mathbf{h})\leq f(\mathbf{x})-f^{*}.$ (52)
In the unconstrained setting, we can leverage the conjugate of the square norm, i.e., $\sup_{\mathbf{h}}\langle\mathbf{x},\mathbf{h}\rangle-\left\|\mathbf{h}\right\|^{2}/2=\left\|\mathbf{x}\right\|_{*}^{2}/2$ (Orabona, 2019, Example 5.11) to cope with LHS of (52). Now that we are dealing with the vector lying within a small ball centered at $\mathbf{x}$, we should apply the clipping technique to lower bound the inner product. Let $\mathbf{u}$ be any unit vector with $\langle\nabla f(\mathbf{x}),\mathbf{u}\rangle=\left\|\nabla f(\mathbf{x})\right\|_{*}$. Then we scale $\mathbf{u}$ to fit the radius of the ball $\mathcal{B}(\mathbf{x},\left\|\nabla f(\mathbf{x})\right\|_{*}/\ell(2\left\|\nabla f(\mathbf{x})\right\|_{*}))$, i.e., set $\mathbf{v}=\mathbf{u}\cdot\left\|\nabla f(\mathbf{x})\right\|_{*}/\ell(2\left\|\nabla f(\mathbf{x})\right\|_{*})$ so that $\left\|\mathbf{v}\right\|=\left\|\nabla f(\mathbf{x})\right\|_{*}/\ell(2\left\|\nabla f(\mathbf{x})\right\|_{*})$. Since (52) holds for any $\mathbf{h}$ satisfying $\left\|\mathbf{h}\right\|\leq\left\|\nabla f(\mathbf{x})\right\|_{*}/\ell(2\left\|\nabla f(\mathbf{x})\right\|_{*})$, we take $\mathbf{h}=\mathbf{v}$ and obtain
$\frac{\left\|\nabla f(\mathbf{x})\right\|_{*}}{\ell(2\left\|\nabla f(\mathbf{x})\right\|_{*})}\langle\nabla f(\mathbf{x}),\mathbf{u}\rangle-\frac{\ell(2\left\|\nabla f(\mathbf{x})\right\|_{*})}{2}\left[\frac{\left\|\nabla f(\mathbf{x})\right\|_{*}}{\ell(2\left\|\nabla f(\mathbf{x})\right\|_{*})}\right]^{2}$ (53)
$=$ $\frac{\left\|\nabla f(\mathbf{x})\right\|_{*}^{2}}{2\ell(2\left\|\nabla f(\mathbf{x})\right\|_{*})}\leq f(\mathbf{x})-f^{*}.$
The following lemma goes one step further to give a quantitative characterization between the gradient’s dual norm and the suboptimality gap.
Lemma D.3.
Under Assumption 3.3, for a given $\mathbf{x}\in\mathcal{X}$ satisfying $f(\mathbf{x})-f^{*}\leq F$, we have $\left\|\nabla f(\mathbf{x})\right\|_{*}\leq G<\infty$ and $G^{2}=2\ell(2G)F$, where $G:=\sup\left\{\alpha\in\mathbb{R}_{+}|\alpha^{2}\leq 2F\ell(2\alpha)\right\}$.
Proof.
Based on Lemma 3.4, the proof is analogous to that of Li et al. (2023a, Corollary 3.6), which involves merely an $f\in\mathcal{F}_{\ell}(\left\|\cdot\right\|_{2})$. Since we have a sub-quadratic $\ell$, we clearly have $\lim_{\alpha\to\infty}\frac{\alpha^{2}}{2\ell\left(2\alpha\right)}=\infty$. So for $F\in\mathbb{R}_{+},\exists\alpha_{F}\in\mathbb{R}_{++}$, such that $\forall\alpha>\alpha_{F},\frac{\alpha^{2}}{2\ell\left(2\alpha\right)}>F$. Therefore, if $\frac{\alpha^{2}}{2\ell\left(2\alpha\right)}\leq F$ holds for some $\alpha$, then it must be that $\alpha\leq\alpha_{F}$. Hence, our construction of $G$ implies that $G\leq\alpha_{F}<\infty$. We conclude that $\left\|\nabla f(\mathbf{x})\right\|_{*}\leq G$ by Lemma 3.4. Lastly, $G^{2}=2\ell(2G)F$ is directly implied by the compactness of the set $\left\{\alpha\in\mathbb{R}_{+}|\alpha^{2}\leq 2F\ell(2\alpha)\right\}$. ∎
Appendix E Analysis for Mirror Descent
We present the following standard lemma for mirror descent, which can be found in Bubeck and others (2015) and Lan (2020).
Lemma E.1.
For any $\mathbf{g}\in\mathcal{E}^{*},\mathbf{x}_{0}\in\mathcal{X}$, define $\mathbf{x}^{+}=\mathcal{P}_{\mathbf{x}_{0}}(\mathbf{g})$, then we have $\left\|\mathbf{x}^{+}-\mathbf{x}_{0}\right\|\leq\left\|\mathbf{g}\right\|_{*}$.
Proof.
Using the first-order optimality condition for mirror descent as well as the three-point equation for Bregman functions on the update (Nemirovski, 2004), we have
$\langle\mathbf{g},\mathbf{x}^{+}-\mathbf{x}\rangle\leq B(\mathbf{x},\mathbf{x}_{0})-B(\mathbf{x},\mathbf{x}^{+})-B(\mathbf{x}^{+},\mathbf{x}_{0}),\quad\forall\mathbf{x}\in\mathcal{X}.$ (54)
Take $\mathbf{x}=\mathbf{x}_{0}$ and rearrange it,
$\langle\mathbf{g},\mathbf{x}_{0}-\mathbf{x}^{+}\rangle\geq B(\mathbf{x}_{0},\mathbf{x}^{+})+B(\mathbf{x}^{+},\mathbf{x}_{0}).$ (55)
Using Cauchy-Schwarz Inequality and the property of Bregman divergence, we obtain
$\left\|\mathbf{g}\right\|_{*}\left\|\mathbf{x}_{0}-\mathbf{x}^{+}\right\|\geq$ $\langle\mathbf{g},\mathbf{x}_{0}-\mathbf{x}^{+}\rangle$ (56)
$\geq$ $B(\mathbf{x}_{0},\mathbf{x}^{+})+B(\mathbf{x}^{+},\mathbf{x}_{0})\geq\frac{1}{2}\left\|\mathbf{x}_{0}-\mathbf{x}^{+}\right\|^{2}+\frac{1}{2}\left\|\mathbf{x}^{+}-\mathbf{x}_{0}\right\|^{2}$
If $\left\|\mathbf{x}_{0}-\mathbf{x}^{+}\right\|=0$Lemma E.1 is naturally true. Otherwise, dividing both sides of the inequality by $\left\|\mathbf{x}_{0}-\mathbf{x}^{+}\right\|$ yields the result. ∎
Lemma E.2.
Under Assumptions 2.5 and 3.3, for any $\mathbf{x}_{0}\in\mathcal{X}$ satisfying $f(\mathbf{x}_{0})-f^{*}\leq F$, define $G:=\sup\left\{\alpha\in\mathbb{R}_{+}|\alpha^{2}\leq 2F\ell(2\alpha)\right\}$ and $\mathbf{x}^{+}=\mathcal{P}_{\mathbf{x}_{0}}(\eta\nabla f(\mathbf{x}_{0}))$. If $0<\eta\leq 1/\ell(2G)$, we have $f(\mathbf{x}^{+})\leq f(\mathbf{x}_{0})$ and $\left\|\nabla f(\mathbf{x}^{+})\right\|_{*}\leq G<\infty$.
Proof.
Choose $\mathbf{g}=\eta\nabla f(\mathbf{x}_{0})$ and $\mathbf{x}=\mathbf{x}_{0}$ in (54), we obtain
$\langle\eta\nabla f(\mathbf{x}_{0}),\mathbf{x}^{+}-\mathbf{x}_{0}\rangle\leq-B(\mathbf{x}_{0},\mathbf{x}^{+})-B(\mathbf{x}^{+},\mathbf{x}_{0}).$ (57)
Under Assumption 2.5, we know that $f^{*}>$-$\infty$, which gives $f(\mathbf{x}_{0})-f^{*}<\infty$. By Lemma D.3, we immediately deduce that $\left\|\nabla f(\mathbf{x}_{0})\right\|_{*}\leq G<\infty$. Then we invoke Lemma E.1 to control the distance between two iterates:
$\left\|\mathbf{x}^{+}-\mathbf{x}_{0}\right\|\leq\eta\left\|\nabla f(\mathbf{x}_{0})\right\|_{*}\leq\frac{\left\|\nabla f(\mathbf{x}_{0})\right\|_{*}}{\ell(2G)}\leq\frac{G}{\ell(2G)},$ (58)
where we use $\eta\leq 1/\ell(2G)$. The conditions of Lemma 2.7 are now satisfied, so we use the effective smoothness property to obtain
$f(\mathbf{x}^{+})\leq f(\mathbf{x}_{0})+\langle\nabla f(\mathbf{x}_{0}),\mathbf{x}^{+}-\mathbf{x}_{0}\rangle+\frac{\ell(2G)}{2}\left\|\mathbf{x}^{+}-\mathbf{x}_{0}\right\|^{2}.$ (59)
Multiplying by $\eta$ and adding this to (57), we have
$\eta f(\mathbf{x}^{+})\leq$ $\eta f(\mathbf{x}_{0})+\frac{\eta\ell(2G)}{2}\left\|\mathbf{x}^{+}-\mathbf{x}_{0}\right\|^{2}-B(\mathbf{x}_{0},\mathbf{x}^{+})-B(\mathbf{x}^{+},\mathbf{x}_{0})$ (60)
$\leq$ $\eta f(\mathbf{x}_{0})+\frac{\eta\ell(2G)-1}{2}\left\|\mathbf{x}^{+}-\mathbf{x}_{0}\right\|^{2}-B(\mathbf{x}_{0},\mathbf{x}^{+})$
$\leq$ $\eta f(\mathbf{x}_{0}),$
where the second inequality uses $B(\mathbf{x}^{+},\mathbf{x}_{0})\geq\frac{1}{2}\left\|\mathbf{x}^{+}-\mathbf{x}_{0}\right\|^{2}$ and the last inequality uses $\eta\leq 1/\ell(2G)$ and the non-negativity of Bregman divergence. Finally, we obtain the following descent property:
$f(\mathbf{x}^{+})-f^{*}\leq f(\mathbf{x}_{0})-f^{*}.$ (61)
Invoking Lemma D.3 again finishes the proof. ∎
Proof of Theorem 3.5.
According to the descent property in (61), we have $f(\mathbf{x}_{t+1})-f^{*}\leq f(\mathbf{x}_{t})-f^{*}\leq\cdots\leq f(\mathbf{x}_{0})-f^{*}$. By Lemma E.2, we deduce that for $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G<\infty$ holds for all $t\in\mathbb{N}$, where $G:=\sup\left\{\alpha\in\mathbb{R}_{+}|\alpha^{2}\leq 2(f(\mathbf{x}_{0})-f^{*})\ell(2\alpha)\right\}$. After establishing such an upper bound for the dual norm of the gradients along the trajectory, we can cope with generalized smooth functions in a similar way as the global smooth functions. Similar to (54), we consider the update of mirror descent in (6):
$\forall\mathbf{x}\in\mathcal{X},\ \langle\eta\nabla f(\mathbf{x}_{t}),\mathbf{x}_{t+1}-\mathbf{x}\rangle\leq B(\mathbf{x},\mathbf{x}_{t})-B(\mathbf{x},\mathbf{x}_{t+1})-B(\mathbf{x}_{t+1},\mathbf{x}_{t}).$ (62)
Note that the local Lipschitz constant $\ell(2\left\|\nabla f(\mathbf{x}_{t})\right\|_{*})$ is bounded by $L=\ell(2G)$. By Lemma E.1 and $\eta\leq\frac{1}{L}$, we have
$\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|\leq\eta\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq\frac{G}{L}.$ (63)
Therefore, the distance between $\mathbf{x}_{t+1}$ and $\mathbf{x}_{t}$ is small enough to satisfy the conditions in Lemma 2.7. Using the effective smoothness characterization, we obtain
$f(\mathbf{x}_{t+1})\leq f(\mathbf{x}_{t})+\langle\nabla f(\mathbf{x}_{t}),\mathbf{x}_{t+1}-\mathbf{x}_{t}\rangle+\frac{L}{2}\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|^{2}.$ (64)
Next, we divide the inner product term $\langle\eta\nabla f(\mathbf{x}_{t}),\mathbf{x}_{t+1}-\mathbf{x}\rangle$ into two parts, and bound them separately:
$\langle\nabla f(\mathbf{x}_{t}),\mathbf{x}_{t+1}-\mathbf{x}\rangle=$ $\langle\nabla f(\mathbf{x}_{t}),\mathbf{x}_{t}-\mathbf{x}\rangle+\langle\nabla f(\mathbf{x}_{t}),\mathbf{x}_{t+1}-\mathbf{x}_{t}\rangle$ (65)
$\geq$ $f(\mathbf{x}_{t})-f(\mathbf{x})+f(\mathbf{x}_{t+1})-f(\mathbf{x}_{t})-\frac{L}{2}\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|^{2}$
$=$ $f(\mathbf{x}_{t+1})-f(\mathbf{x})-\frac{L}{2}\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|^{2},$
where the inequality uses the convexity of $f$ and (64). Combining (62) with (65), we derive
$\eta\left[f(\mathbf{x}_{t+1})-f(\mathbf{x})\right]-\frac{\eta L}{2}\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|^{2}\leq B(\mathbf{x},\mathbf{x}_{t})-B(\mathbf{x},\mathbf{x}_{t+1})-B(\mathbf{x}_{t+1},\mathbf{x}_{t})$ (66)
After rearrangement and due to the fact that $B(\mathbf{x}_{t+1},\mathbf{x}_{t})\geq\frac{1}{2}\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|^{2}$, we have
$\eta\left[f(\mathbf{x}_{t+1})-f(\mathbf{x})\right]\leq$ $B(\mathbf{x},\mathbf{x}_{t})-B(\mathbf{x},\mathbf{x}_{t+1})+\frac{\eta L-1}{2}\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|^{2}$ (67)
$\leq$ $B(\mathbf{x},\mathbf{x}_{t})-B(\mathbf{x},\mathbf{x}_{t+1})$
Telescoping it and using Jensen’s Inequality,
$f(\overline{\mathbf{x}}_{T})-f(\mathbf{x})\leq$ $\frac{\sum_{t=0}^{T-1}\eta\left[f(\mathbf{x}_{t+1})-f(\mathbf{x})\right]}{\eta T}$ (68)
$\leq$ $\frac{\sum_{t=0}^{T-1}B(\mathbf{x},\mathbf{x}_{t})-B(\mathbf{x},\mathbf{x}_{t+1})}{\eta T}\leq\frac{B(\mathbf{x},\mathbf{x}_{0})}{\eta T},$
Taking $\mathbf{x}=\mathbf{x}_{*}$ gives the desired convergence for the average-iterate. As for the last iterate, recall the descent property in (61), we have $\sum_{t=0}^{T-1}f(\mathbf{x}_{t+1})-f(\mathbf{x})\geq T[f(\mathbf{x}_{T})-f(\mathbf{x})]$. Thus, the convergence rate is exactly the same as that of $\overline{\mathbf{x}}_{T}$. ∎
Appendix F Analysis for Accelerated Mirror Descent
The lemmas for this section are stated under the conditions in Theorem 3.7. For simplicity, we do not specify this point below.
Lemma F.1.
Define $e_{t}:=\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|$ for any $t\in\mathbb{N}_{+}$, then we have $e_{1}=e_{2}=0$. If
$e_{t-1}\leq G/L,\quad\left\|\nabla f(\mathbf{x}_{t-2})\right\|_{*}\leq G$ (69)
hold for any $t\geq 2$, we further have
$e_{t}=\frac{2}{t+1}\left\|\mathbf{x}_{t-1}-\mathbf{z}_{t-1}\right\|\leq\frac{(t-2)(\eta(t-1)+t)}{t(t+1)}e_{t-1}+\frac{(t-2)(t-1)\eta G}{t(t+1)L},\quad\forall t\geq 2.$ (70)
Proof.
By (8), the following holds for any $t\geq 3$:
$\mathbf{y}_{t}-\mathbf{x}_{t-1}=\alpha_{t}(\mathbf{z}_{t-1}-\mathbf{x}_{t-1})=\alpha_{t}(1-\alpha_{t-1})(\mathbf{z}_{t-1}-\mathbf{x}_{t-2}).$ (71)
We consider the last term above:
$\left\|\mathbf{z}_{t-1}-\mathbf{x}_{t-2}\right\|\leq\left\|\mathbf{z}_{t-1}-\mathbf{z}_{t-2}\right\|+\left\|\mathbf{z}_{t-2}-\mathbf{x}_{t-2}\right\|\leq\eta_{t-1}\left\|\nabla f(\mathbf{y}_{t-1})\right\|_{*}+\frac{e_{t-1}}{\alpha_{t-1}},$ (72)
where the last inequality is due to (8) and Lemma E.1. Then we use a similar idea as in Equation 34 to bound $\left\|\nabla f(\mathbf{y}_{t-1})\right\|_{*}$:
$\left\|\nabla f(\mathbf{y}_{t-1})\right\|_{*}\leq\left\|\nabla f(\mathbf{y}_{t-1})-\nabla f(\mathbf{x}_{t-2})\right\|_{*}+\left\|\nabla f(\mathbf{x}_{t-2})\right\|_{*}\leq Le_{t-1}+G,$ (73)
where we use (69) and Lemma 2.7. Combine the previous results, we obtain
$e_{t}\leq$ $\alpha_{t}(1-\alpha_{t-1})\left[\eta_{t-1}(Le_{t-1}+G)+\frac{e_{t-1}}{\alpha_{t-1}}\right]$ (74)
$\leq$ $\frac{2}{t+1}\cdot\frac{t-2}{t}\left[\frac{\eta(t-1)G}{2L}+\frac{\eta(t-1)e_{t-1}}{2}+\frac{te_{t-1}}{2}\right]$
$\leq$ $\frac{(t-2)(t-1)}{t(t+1)}\cdot\frac{\eta G}{L}+\frac{(t-2)(\eta(t-1)+t)}{t(t+1)}\cdot e_{t-1}.$
For $e_{2}$, we also have $e_{2}=2/(2+1)\left\|\mathbf{x}_{1}-\mathbf{z}_{1}\right\|=0$ based on (8). $e_{1}=0$ is a direct consequence of our initialization. ∎
Lemma F.2.
If $\mathbf{x}_{t},\mathbf{y}_{t}$ satisfies $\left\|\nabla f(\mathbf{y}_{t})\right\|_{*}\leq 2G$ and $\left\|\mathbf{x}_{t}-\mathbf{y}_{t}\right\|\leq 2G/L$, then for any $\mathbf{x}\in\mathcal{X}$:
$\frac{\eta t(t+1)}{4L}\left[f(\mathbf{x}_{t})-f(\mathbf{x})\right]\leq\frac{\eta t(t-1)}{4L}\left[f(\mathbf{x}_{t-1})-f(\mathbf{x})\right]+B(\mathbf{x},\mathbf{z}_{t-1})-B(\mathbf{x},\mathbf{z}_{t}),\forall t\in\mathbb{N}_{+}.$ (75)
Proof.
In view of Lemma 2.7 as well as the given condition, we have
$f(\mathbf{x}_{t})\leq$ $f(\mathbf{y}_{t})+\langle\nabla f(\mathbf{y}_{t}),\mathbf{x}_{t}-\mathbf{y}_{t}\rangle+\frac{L}{2}\left\|\mathbf{x}_{t}-\mathbf{y}_{t}\right\|^{2}$ (76)
$=$ $(1-\alpha_{t})\left[f(\mathbf{y}_{t})+\langle\nabla f(\mathbf{y}_{t}),\mathbf{x}_{t-1}-\mathbf{y}_{t}\rangle\right]$
$+\alpha_{t}\left[f(\mathbf{y}_{t})+\langle\nabla f(\mathbf{y}_{t}),\mathbf{z}_{t}-\mathbf{y}_{t}\rangle\right]+\frac{\alpha_{t}^{2}L}{2}\left\|\mathbf{z}_{t}-\mathbf{z}_{t-1}\right\|^{2}$
$\leq$ $(1-\alpha_{t})f(\mathbf{x}_{t-1})+\alpha_{t}\left[f(\mathbf{y}_{t})+\langle\nabla f(\mathbf{y}_{t}),\mathbf{z}_{t}-\mathbf{y}_{t}\rangle+\frac{1}{\eta_{t}}B(\mathbf{z}_{t},\mathbf{z}_{t-1})\right],$
where the first equality is due to $\mathbf{x}_{t}=(1-\alpha_{t})\mathbf{x}_{t-1}+\alpha_{t}\mathbf{z}_{t}$ and $\mathbf{x}_{t}-\mathbf{y}_{t}=\alpha_{t}(\mathbf{z}_{t}-\mathbf{z}_{t-1})$; the first inequality uses the convexity of $f$, the property of Bregman divergence and the fact $\alpha_{t}L\leq\eta_{t}$ (since $\eta\leq 1$). Similar to (54), the following holds for any $\mathbf{x}\in\mathcal{X}$:
$\langle\nabla f(\mathbf{y}_{t}),\mathbf{z}_{t}-\mathbf{x}\rangle\leq\frac{B(\mathbf{x},\mathbf{z}_{t-1})-B(\mathbf{x},\mathbf{z}_{t})-B(\mathbf{z}_{t},\mathbf{z}_{t-1})}{\eta_{t}}.$ (77)
Plugging the above relation into the previous derivations, we obtain
$f(\mathbf{x}_{t})\leq$ $(1-\alpha_{t})f(\mathbf{x}_{t-1})+\alpha_{t}\left[f(\mathbf{y}_{t})+\langle\nabla f(\mathbf{y}_{t}),\mathbf{x}-\mathbf{y}_{t}\rangle+\frac{B(\mathbf{x},\mathbf{z}_{t-1})-B(\mathbf{x},\mathbf{z}_{t})}{\eta_{t}}\right]$ (78)
$\leq$ $(1-\alpha_{t})f(\mathbf{x}_{t-1})+\alpha_{t}f(\mathbf{x})+\frac{\alpha_{t}}{\eta_{t}}\left[B(\mathbf{x},\mathbf{z}_{t-1})-B(\mathbf{x},\mathbf{z}_{t})\right],$
where the second step uses the convexity of $f$. A simple rearrangement yields
$\frac{\eta_{t}}{\alpha_{t}}\left[f(\mathbf{x}_{t})-f(\mathbf{x})\right]\leq\frac{(1-\alpha_{t})\eta_{t}}{\alpha_{t}}\left[f(\mathbf{x}_{t-1})-f(\mathbf{x})\right]+B(\mathbf{x},\mathbf{z}_{t-1})-B(\mathbf{x},\mathbf{z}_{t}).$ (79)
Specifying the value of $\alpha_{t}$ and $\eta_{t}$ completes the proof. ∎
Lemma F.3.
The trajectory generated by (8) satisfies the following relations for all $t\in\mathbb{N}_{+}$:
  1. 1.
    (bounded suboptimality gap) $\quad f(\mathbf{x}_{t})-f^{*}\leq f(\mathbf{x}_{0})-f^{*}$.
  2. 2.
    (bounded gradients) $\qquad\qquad\ \left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G<\infty,\left\|\nabla f(\mathbf{y}_{t})\right\|_{*}\leq 2G<\infty$.
  3. 3.
    (bounded sequence) $\qquad\qquad\ \ e_{t}\leq G/L$.
  4. 4.
    (bounded trajectory) $\qquad\qquad\ \ \mathbf{x}_{t},\mathbf{z}_{t}\in\mathcal{B}\left(\mathbf{x}_{*},\sqrt{2B(\mathbf{x}_{*},\mathbf{x}_{0})}\right)$.
Proof.
First, we prove the bounded trajectory property. Define the Lyapunov function $E_{t}$ as
$E_{t}:=\frac{\eta t(t+1)}{4L}\left[f(\mathbf{x}_{t})-f^{*}\right]+B(\mathbf{x}_{*},\mathbf{z}_{t})\geq B(\mathbf{x}_{*},\mathbf{z}_{t})\geq\frac{1}{2}\left\|\mathbf{x}_{*}-\mathbf{z}_{t}\right\|^{2}.$ (80)
Equation 75 states that $E_{t}\leq E_{t-1}\leq\cdots E_{1}\leq E_{0}=B(\mathbf{x},\mathbf{x}_{0})$. Combining with (80), we deduce that $\left\|\mathbf{z}_{t}-\mathbf{x}_{*}\right\|\leq\sqrt{2B(\mathbf{x}_{*},\mathbf{x}_{0})}$, implying $\mathbf{z}_{t}\in\mathcal{B}\left(\mathbf{x}_{*},\sqrt{2B(\mathbf{x}_{*},\mathbf{x}_{0})}\right)$. For $\mathbf{x}_{t}$, recall the update rule of $\mathbf{x}_{t}=(1-\alpha_{t})\mathbf{x}_{t-1}+\alpha_{t}\mathbf{z}_{t}$ and that $\mathbf{x}_{0}=\mathbf{z}_{0}$. We deduce that $\mathbf{x}_{t}$ is a convex combination of $\left\{\mathbf{z}_{s}\right\}_{s=0}^{t}$. By convexity of $\mathcal{B}\left(\mathbf{x}_{*},\sqrt{2B(\mathbf{x}_{*},\mathbf{x}_{0})}\right)$, it holds that $\mathbf{x}_{t}\in\mathcal{B}\left(\mathbf{x}_{*},\sqrt{2B(\mathbf{x}_{*},\mathbf{x}_{0})}\right)$. For conditions 1 to 3, we proceed by induction, as specified below.
Part 1. Base Case
For $t=1$, we clearly have $\mathbf{y}_{1}=\mathbf{z}_{0}=\mathbf{x}_{0}$ and thus $\left\|\nabla f(\mathbf{y}_{1})\right\|_{*}\leq G$. Also, we have $\left\|\mathbf{x}_{1}-\mathbf{y}_{1}\right\|=\left\|\mathbf{z}_{1}-\mathbf{z}_{0}\right\|\leq\eta_{1}\left\|\nabla f(\mathbf{z}_{0})\right\|_{*}\leq G/(2L)$. In this scenario, the conditions of Equation 75 are satisfied, we immediately obtain
$\frac{\eta}{2L}\left[f(\mathbf{x}_{1})-f(\mathbf{x}_{0})\right]\leq-B(\mathbf{z}_{0},\mathbf{z}_{1})\leq 0,$ (81)
by choosing $\mathbf{x}=\mathbf{x}_{0}=\mathbf{z}_{0}$. Then we conclude that $f(\mathbf{x}_{1})-f^{*}\leq f(\mathbf{x}_{0})-f^{*}$. By Lemma D.3, we have $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G<\infty$. By definition, we have $e_{1}=0$, which completes the proof for this part.
Part 2. Induction Step
Suppose the statements hold for all $t\leq s-1$ where $s\geq 2$. Our goal is to show that they also hold for $t=s$. The main idea is to show the conditions of Equation 75 hold via Equation 70. Then we establish the descent property under Equation 75, which is the key to the bounded suboptimality gap as well as the bounded gradients. First, we consider the third statement by splitting the timeline.
Case (i). $s\geq\tau\geq 3$
Recall the definition of $\tau$ in Theorem 3.7:
$\tau=\left\lceil\frac{4\sqrt{2B(\mathbf{x}_{*},\mathbf{x}_{0})}L}{G}\right\rceil-1=\left\lceil\frac{8\sqrt{2B(\mathbf{x}_{*},\mathbf{x}_{0})}}{\frac{2G}{L}}\right\rceil-1\geq\left\lceil\frac{8\sqrt{2B(\mathbf{x}_{*},\mathbf{x}_{0})}}{2\sqrt{2B(\mathbf{x}_{*},\mathbf{x}_{0})}}\right\rceil-1\geq 3.$ (82)
where we use $2G/L\leq 2\sqrt{2B(\mathbf{x}_{*},\mathbf{x}_{0})}$ (Yu et al., 2024, Proposition A.6). Then we have
$e_{s}=$ $\frac{2}{s+1}\left\|\mathbf{x}_{t-1}-\mathbf{z}_{t-1}\right\|\leq\frac{2}{\tau+1}\left(\left\|\mathbf{x}_{t-1}-\mathbf{x}_{*}\right\|+\left\|\mathbf{z}_{t-1}-\mathbf{x}_{*}\right\|\right)$ (83)
$\leq$ $\frac{2}{\tau+1}\cdot 2\sqrt{2B(\mathbf{x}_{*},\mathbf{x}_{0})}\leq\frac{4\sqrt{2B(\mathbf{x}_{*},\mathbf{x}_{0})}}{\frac{4\sqrt{2B(\mathbf{x}_{*},\mathbf{x}_{0})}L}{G}}\leq\frac{G}{L}.$
Case (ii). $2\leq s\leq\tau-1$
From the induction basis, we see that the conditions of Equation 70 are satisfied. Then we have
$e_{s}\leq$ $\frac{(s-2)(\eta(s-1)+s)}{s(s+1)}e_{s-1}+\frac{(s-2)(s-1)\eta G}{s(s+1)L}$ (84)
$\leq$ $\underbrace{\frac{(\tau-3)(\eta(\tau-2)+\tau-1)}{\tau(\tau-1)}}_{q_{\tau}}\cdot e_{s-1}+\underbrace{\frac{(\tau-3)(\tau-2)\eta}{\tau(\tau-1)}}_{a_{\tau}}\cdot\frac{G}{L}.$
We aim to bound $e_{s}$ as a contraction mapping. For the ratio parameter, we have
$q_{\tau}=\frac{\tau-3}{\tau}+\frac{(\tau-3)(\tau-2)\eta}{\tau(\tau-1)}\leq 1-\frac{3}{\tau}+\frac{3}{2\tau}<1,$ (85)
where we use the definition of $\eta$. Then we obtain
$e_{s}-\frac{a_{\tau}}{1-q_{\tau}}\leq q_{\tau}(e_{s-1}-\frac{a_{\tau}}{1-q_{\tau}})\leq\cdots\leq q_{\tau}^{s-2}(e_{2}-\frac{a_{\tau}}{1-q_{\tau}})\overset{e_{2}=0}{\leq}0.$ (86)
Thus,
$e_{s}\leq\frac{a_{\tau}}{1-q_{\tau}}\leq\frac{\frac{3}{2\tau}}{\frac{3}{\tau}-\frac{3}{2\tau}}\cdot\frac{G}{L}\leq\frac{G}{L}.$ (87)
Therefore, we have proven the third statement.
Next, we show that the conditions of Equation 75 are satisfied.
$\left\|\nabla f(\mathbf{y}_{s})\right\|_{*}\leq\left\|\nabla f(\mathbf{y}_{s})-\nabla f(\mathbf{x}_{s-1})\right\|_{*}+\left\|\nabla f(\mathbf{x}_{s-1})\right\|_{*}\leq Le_{s-1}+G\leq 2G,$ (88)
where we use the induction basis $\left\|\nabla f(\mathbf{x}_{s-1})\right\|_{*}\leq G,e_{s}\leq G/L$ and Lemma 2.7. We also have
$\left\|\mathbf{x}_{t}-\mathbf{y}_{t}\right\|=\alpha_{t}\left\|\mathbf{z}_{t}-\mathbf{z}_{t-1}\right\|\leq\alpha_{t}\eta_{t}\left\|\nabla f(\mathbf{y}_{t})\right\|_{*}\leq\frac{t\eta G}{(t+1)L}\leq\frac{G}{L}.$ (89)
Then we invoke Equation 75 to get the following
$\frac{\eta s(s+1)}{4L}\left[f(\mathbf{x}_{s})-f(\mathbf{x}_{0})\right]\leq\frac{\eta s(s-1)}{4L}\left[f(\mathbf{x}_{s-1})-f(\mathbf{x}_{0})\right]+B(\mathbf{x}_{0},\mathbf{z}_{s-1})-B(\mathbf{x}_{0},\mathbf{z}_{s}),$ (90)
which also holds for all $1\leq t\leq s-1$ by induction. Summing from $t=1$ to $t=s$, we obtain
$\frac{\eta s(s+1)}{4L}\left[f(\mathbf{x}_{s})-f(\mathbf{x}_{0})\right]\leq-B(\mathbf{x}_{0},\mathbf{z}_{s})\leq 0,$ (91)
which yields $f(\mathbf{x}_{t})-f^{*}\leq f(\mathbf{x}_{0})-f^{*}$. $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G<\infty$ is immediately derived using Lemma D.3. Finally, the three statements are all proven, which completes the proof. ∎
Proof of Theorem 3.7.
By Lemma F.3, we immediately conclude that
$\max\{\left\|\nabla f(\mathbf{y}_{t})\right\|_{*},2\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\}\leq 2G<\infty,\quad\forall t\in\mathbb{N}_{+}.$ (92)
Similar to (89), we have $\left\|\mathbf{x}_{t}-\mathbf{y}_{t}\right\|\leq 2G/L$. Then, we invoke Equation 75 to derive the following:
$\frac{\eta t(t+1)}{4L}\left[f(\mathbf{x}_{t})-f(\mathbf{x})\right]\leq\frac{\eta t(t-1)}{4L}\left[f(\mathbf{x}_{t-1})-f(\mathbf{x})\right]+B(\mathbf{x},\mathbf{z}_{t-1})-B(\mathbf{x},\mathbf{z}_{t}),$ (93)
which holds for all $t\in\mathbb{N}_{+}$ and $\mathbf{x}\in\mathcal{X}$. Summing from $1$ to $T$ and choosing $\mathbf{x}=\mathbf{x}_{*}$, we obtain
$\frac{\eta T(T+1)}{4L}\left[f(\mathbf{x}_{T})-f^{*}\right]\leq B(\mathbf{x}_{*},\mathbf{z}_{0})-B(\mathbf{x}_{*},\mathbf{z}_{T})\leq B(\mathbf{x}_{*},\mathbf{x}_{0}).$ (94)
A simple rearrangement yields the desired convergence result. ∎
Appendix G Analysis for Optimistic Mirror Descent
Lemma G.1.
Consider the $t$-th iteration of the optimistic mirror descent algorithm defined in (10). The following relations hold:
$\eta(f(\mathbf{y}_{t})-f(\mathbf{x}_{t-1}))\leq$ $\eta\langle\nabla f(\mathbf{y}_{t})-\nabla f(\mathbf{y}_{t-1}),\mathbf{y}_{t}-\mathbf{x}_{t-1}\rangle$ (95)
$-\frac{1}{2}\left\|\mathbf{x}_{t}-\mathbf{y}_{t}\right\|^{2}-\frac{1}{2}\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|^{2}-\frac{1}{2}\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|^{2},$
$\eta(f(\mathbf{x}_{t})-f(\mathbf{x}_{t-1}))\leq$ $\eta\langle\nabla f(\mathbf{x}_{t})-\nabla f(\mathbf{y}_{t-1}),\mathbf{y}_{t}-\mathbf{x}_{t-1}\rangle$ (96)
$+\eta\langle\nabla f(\mathbf{x}_{t})-\nabla f(\mathbf{y}_{t}),\mathbf{x}_{t}-\mathbf{y}_{t}\rangle$
$-\frac{1}{2}\left\|\mathbf{x}_{t}-\mathbf{y}_{t}\right\|^{2}-\frac{1}{2}\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|^{2}-\frac{1}{2}\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|^{2}.$
Proof.
By Equation 139, we have
$\forall\mathbf{x}\in\mathcal{X},\langle\eta\nabla f(\mathbf{y}_{t-1}),\mathbf{y}_{t}-\mathbf{x}\rangle\leq B(\mathbf{x},\mathbf{x}_{t-1})-B(\mathbf{x},\mathbf{y}_{t})-B(\mathbf{y}_{t},\mathbf{x}_{t-1});$ (97)
$\forall\mathbf{x}\in\mathcal{X},\langle\eta\nabla f(\mathbf{y}_{t}),\mathbf{x}_{t}-\mathbf{x}\rangle\leq B(\mathbf{x},\mathbf{x}_{t-1})-B(\mathbf{x},\mathbf{x}_{t})-B(\mathbf{x}_{t},\mathbf{x}_{t-1}).$ (98)
Choose $\mathbf{x}=\mathbf{x}_{t}$ in (97), $\mathbf{x}=\mathbf{x}_{t-1}$ in (98) and then add them together:
$\langle\eta\nabla f(\mathbf{y}_{t-1}),\mathbf{y}_{t}-\mathbf{x}_{t}\rangle+\langle\eta\nabla f(\mathbf{y}_{t}),\mathbf{x}_{t}-\mathbf{x}_{t-1}\rangle$ (99)
$\leq$ $-B(\mathbf{x}_{t},\mathbf{y}_{t})-B(\mathbf{y}_{t},\mathbf{x}_{t-1})-B(\mathbf{x}_{t-1},\mathbf{x}_{t}).$
By convexity of $f$ and the property of Bregman function, we obtain
$\eta(f(\mathbf{y}_{t})-f(\mathbf{x}_{t-1}))$ (100)
$\leq$ $\langle\eta\nabla f(\mathbf{y}_{t}),\mathbf{y}_{t}-\mathbf{x}_{t-1}\rangle$
$=$ $\eta\langle\nabla f(\mathbf{y}_{t})-\nabla f(\mathbf{y}_{t-1}),\mathbf{y}_{t}-\mathbf{x}_{t}\rangle+\langle\eta f(\mathbf{y}_{t-1}),\mathbf{y}_{t}-\mathbf{x}_{t}\rangle+\langle\eta f(\mathbf{y}_{t}),\mathbf{x}_{t}-\mathbf{x}_{t-1}\rangle$
$\leq$ $\eta\langle\nabla f(\mathbf{y}_{t})-\nabla f(\mathbf{y}_{t-1}),\mathbf{y}_{t}-\mathbf{x}_{t}\rangle-\frac{1}{2}\left\|\mathbf{x}_{t}-\mathbf{y}_{t}\right\|^{2}-\frac{1}{2}\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|^{2}-\frac{1}{2}\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|^{2}.$
To prove the second bound, we first notice that
$\langle\nabla f(\mathbf{x}_{t}),\mathbf{x}_{t}-\mathbf{x}_{t-1}\rangle=$ $\underbrace{\langle\nabla f(\mathbf{y}_{t-1}),\mathbf{y}_{t}-\mathbf{x}_{t-1}\rangle}_{A_{t}}+\underbrace{\langle\nabla f(\mathbf{x}_{t})-\nabla f(\mathbf{y}_{t-1}),\mathbf{y}_{t}-\mathbf{x}_{t-1}\rangle}_{B_{t}}$ (101)
$+\underbrace{\langle\nabla f(\mathbf{y}_{t}),\mathbf{x}_{t}-\mathbf{y}_{t-1}\rangle}_{C_{t}}+\underbrace{\langle\nabla f(\mathbf{x}_{t})-\nabla f(\mathbf{y}_{t}),\mathbf{x}_{t}-\mathbf{y}_{t}\rangle}_{D_{t}}.$
Similar to the proof of Lemma H.3, we bound the four terms separately. To bound $A_{t}$ and $C_{t}$, we choose $\mathbf{x}=\mathbf{x}_{t-1}$ in (97), $\mathbf{x}=\mathbf{y}_{t}$ in (98). Adding them up, we have
$\eta(A_{t}+C_{t})=$ $\langle\eta\nabla f(\mathbf{y}_{t-1}),\mathbf{y}_{t}-\mathbf{x}_{t-1}\rangle+\langle\eta\nabla f(\mathbf{y}_{t}),\mathbf{x}_{t}-\mathbf{y}_{t}\rangle$ (102)
$\leq$ $-B(\mathbf{x}_{t-1},\mathbf{y}_{t})-B(\mathbf{y}_{t},\mathbf{x}_{t})-B(\mathbf{x}_{t},\mathbf{x}_{t-1}).$
By convexity of $f$ and the property of Bregman function, we obtain
$\eta(f(\mathbf{x}_{t})-f(\mathbf{x}_{t-1}))\leq$ $\eta\langle\nabla f(\mathbf{x}_{t}),\mathbf{x}_{t}-\mathbf{x}_{t-1}\rangle$ (103)
$\leq$ $\eta(B_{t}+D_{t})-\frac{1}{2}\left\|\mathbf{x}_{t-1}-\mathbf{y}_{t}\right\|^{2}-\frac{1}{2}\left\|\mathbf{y}_{t}-\mathbf{x}_{t}\right\|^{2}-\frac{1}{2}\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|^{2},$
which completes the proof by plugging in the definitions of $B_{t}$ and $D_{t}$. ∎
Lemma G.2.
Under Assumptions 2.5 and 3.3, consider the updates in (10) and define $G:=\sup\{\alpha\in\mathbb{R}_{+}|\alpha^{2}\leq 2(f(\mathbf{x}_{0})-f^{*})\ell(2\alpha)\}$. If $0<\eta\leq\gamma/\ell(2G)$ for some $\gamma\in(0,1/3]$, the following statements holds for all $t\in\mathbb{N}_{+}$:
  1. 1.
    (bounded suboptimality gap) $\quad f(\mathbf{y}_{t})-f^{*}\leq f(\mathbf{x}_{0})-f^{*},f(\mathbf{x}_{t})-f^{*}\leq f(\mathbf{x}_{0})-f^{*}$;
  2. 2.
    (bounded gradients) $\quad\quad\quad\quad\ \left\|\nabla f(\mathbf{y}_{t})\right\|_{*}\leq G<\infty,\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G<\infty$.
  3. 3.
    (tracking $\{\mathbf{y}_{t}\}_{t\in\mathbb{N}}$ sequence) $\quad\left\|\mathbf{y}_{t}-\mathbf{y}_{t-1}\right\|\leq\frac{G}{L}\sum_{s=1}^{t}\gamma^{s}\leq\frac{\gamma G}{(1-\gamma)L}<\frac{G}{L}$.
Proof.
The proof resembles that of Lemma E.2, which involves induction. For convenience, we denote $L:=\ell(2G)$.
Part 1. Base Case
Consider the first round of (10). Similar to the proof of Lemma E.2, we argue that $f(\mathbf{y}_{0})-f^{*}=f(\mathbf{x}_{0})-f^{*}<\infty$. By Lemma 3.4, we have $\left\|\nabla f(\mathbf{y}_{0})\right\|_{*}=\left\|\nabla f(\mathbf{y}_{0})\right\|_{*}\leq G<\infty$. According to Equation 96, we have
$\eta(f(\mathbf{y}_{1})-f(\mathbf{x}_{0}))\leq$ $\eta\langle\nabla f(\mathbf{y}_{1})-\nabla f(\mathbf{y}_{0}),\mathbf{y}_{1}-\mathbf{x}_{0}\rangle$ (104)
$-\frac{1}{2}\left\|\mathbf{x}_{1}-\mathbf{y}_{1}\right\|^{2}-\frac{1}{2}\left\|\mathbf{y}_{1}-\mathbf{x}_{0}\right\|^{2}-\frac{1}{2}\left\|\mathbf{x}_{1}-\mathbf{x}_{0}\right\|^{2}.$
It suffices to control the inner product term. By Lemma E.1 as well as our initialization $\mathbf{x}_{0}=\mathbf{y}_{0}$, we have
$\left\|\mathbf{y}_{1}-\mathbf{y}_{0}\right\|=\left\|\mathbf{y}_{1}-\mathbf{x}_{0}\right\|\leq\eta\left\|\nabla f(\mathbf{x}_{0})\right\|_{*}\leq\frac{\gamma G}{L}\leq\frac{G}{L}.$ (105)
Then we use Lemma D.2 to establish the following upper bound:
$\eta\langle\nabla f(\mathbf{y}_{1})-\nabla f(\mathbf{y}_{0}),\mathbf{y}_{1}-\mathbf{x}_{1}\rangle\leq\frac{\gamma^{2}}{2}\left\|\mathbf{y}_{1}-\mathbf{y}_{0}\right\|^{2}+\frac{1}{2}\left\|\mathbf{y}_{1}-\mathbf{x}_{1}\right\|^{2}.$ (106)
Noticing $\mathbf{x}_{0}=\mathbf{y}_{0}$ and $\gamma^{2}<1$, we obtain
$\eta(f(\mathbf{y}_{1})-f(\mathbf{x}_{0}))\leq-\frac{1}{2}\left\|\mathbf{x}_{1}-\mathbf{x}_{0}\right\|^{2}\leq 0.$ (107)
Now we clearly have $f(\mathbf{y}_{1})-f^{*}\leq f(\mathbf{x}_{0})-f^{*}$ as $\eta>0$. By Lemma D.3, we conclude that $\left\|\nabla f(\mathbf{y}_{1})\right\|_{*}\leq G$. Next, we prove the other half of the statements. By Equation 96, we have
$\eta(f(\mathbf{x}_{1})-f(\mathbf{x}_{0}))\leq$ $\eta\langle\nabla f(\mathbf{x}_{1})-\nabla f(\mathbf{y}_{0}),\mathbf{y}_{1}-\mathbf{x}_{0}\rangle+\eta\langle\nabla f(\mathbf{x}_{1})-\nabla f(\mathbf{y}_{1}),\mathbf{x}_{1}-\mathbf{y}_{1}\rangle$ (108)
$-\frac{1}{2}\left\|\mathbf{x}_{0}-\mathbf{y}_{1}\right\|^{2}-\frac{1}{2}\left\|\mathbf{y}_{1}-\mathbf{x}_{1}\right\|^{2}-\frac{1}{2}\left\|\mathbf{x}_{1}-\mathbf{x}_{0}\right\|^{2}.$
In the language of Equation 96, we shall establish an upper bound for the terms $\eta B_{1}$ and $\eta D_{1}$, i.e., the inner product terms in the above inequality. By Lemma E.1, we have
$\left\|\mathbf{x}_{1}-\mathbf{y}_{0}\right\|=\left\|\mathbf{x}_{1}-\mathbf{x}_{0}\right\|\leq\eta\left\|\nabla f(\mathbf{y}_{1})\right\|_{*}\leq\frac{\gamma G}{L}\leq\frac{G}{L},$ (109)
and by Equation 139,
$\left\|\mathbf{x}_{1}-\mathbf{y}_{1}\right\|\leq\eta\left\|\nabla f(\mathbf{y}_{1})-\nabla f(\mathbf{y}_{0})\right\|_{*}\leq\eta L\left\|\mathbf{y}_{1}-\mathbf{y}_{0}\right\|\leq\frac{\gamma^{2}G}{L}\leq\frac{G}{L},$ (110)
where the second inequality uses $\left\|\mathbf{y}_{1}-\mathbf{y}_{0}\right\|\leq\frac{G}{L}$ and Lemma 2.7. Then we invoke Lemma D.2 to derive the following bounds for $\eta B_{1}$ (set $\beta=1$) as well as $\eta D_{1}$ (set $\beta=\frac{1}{2}$):
$\eta B_{1}=\eta\langle\nabla f(\mathbf{x}_{1})-\nabla f(\mathbf{y}_{0}),\mathbf{y}_{1}-\mathbf{x}_{0}\rangle\leq\frac{\gamma^{2}}{2}\left\|\mathbf{x}_{1}-\mathbf{y}_{0}\right\|^{2}+\frac{1}{2}\left\|\mathbf{y}_{1}-\mathbf{x}_{0}\right\|^{2}.$ (111)
$\eta D_{1}=\eta\langle\nabla f(\mathbf{x}_{1})-\nabla f(\mathbf{y}_{1}),\mathbf{x}_{1}-\mathbf{y}_{1}\rangle\leq\frac{1}{2}\left\|\mathbf{x}_{1}-\mathbf{y}_{1}\right\|^{2}.$ (112)
Combine (108) with the above inequalities, we obtain
$\eta(f(\mathbf{x}_{1})-f(\mathbf{x}_{0}))\leq$ $\frac{\gamma^{2}}{2}\left\|\mathbf{x}_{1}-\mathbf{y}_{0}\right\|^{2}+\frac{1}{2}\left\|\mathbf{y}_{1}-\mathbf{x}_{0}\right\|^{2}+\frac{1}{2}\left\|\mathbf{x}_{1}-\mathbf{y}_{1}\right\|^{2}$ (113)
$-\frac{1}{2}\left\|\mathbf{x}_{0}-\mathbf{y}_{1}\right\|^{2}-\frac{1}{2}\left\|\mathbf{y}_{1}-\mathbf{x}_{1}\right\|^{2}-\frac{1}{2}\left\|\mathbf{x}_{1}-\mathbf{x}_{0}\right\|^{2}$
$\leq$ $-\frac{1-\gamma^{2}}{2}\left\|\mathbf{x}_{1}-\mathbf{x}_{0}\right\|^{2}\overset{\gamma^{2}<1}{\leq}0,$
where we use the initialization setup of $\mathbf{x}_{0}=\mathbf{y}_{0}$. Consequently, we conclude that $f(\mathbf{x}_{1})-f^{*}\leq f(\mathbf{x}_{0})-f^{*}$. $\left\|\nabla f(\mathbf{y}_{1})\right\|_{*}\leq G$ is straightforward after a standard application of Lemma D.3.
Part 2. Induction Step
Suppose the function values and the gradients are bounded before the $t$-th iteration. Thus we have $\left\|\nabla f(\mathbf{y}_{t-1})\right\|_{*}\leq G,\left\|\nabla f(\mathbf{x}_{t-1})\right\|_{*}\leq G$ and $\left\|\mathbf{y}_{t-1}-\mathbf{y}_{t-2}\right\|\leq\frac{G}{L}\sum_{s=1}^{t-1}\gamma^{s}\leq\frac{\gamma G}{(1-\gamma)L}$. We prove statement 3 first, which tracks the amount of change of the sequence $\{\mathbf{y}_{t}\}_{t\in\mathbb{N}}$.
$\left\|\mathbf{y}_{t}-\mathbf{y}_{t-1}\right\|\leq$ $\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|+\left\|\mathbf{x}_{t-1}-\mathbf{y}_{t-1}\right\|$ (114)
$\leq$ $\eta\left\|\nabla f(\mathbf{y}_{t-1})\right\|_{*}+\eta\left\|\nabla f(\mathbf{y}_{t-1})-\nabla f(\mathbf{y}_{t-2})\right\|_{*}$
$\leq$ $\eta G+\eta L\left\|\mathbf{y}_{t-1}-\mathbf{y}_{t-2}\right\|$
$\leq$ $\frac{\gamma G}{L}+\frac{\gamma G}{L}\sum_{s=1}^{t-1}\gamma^{s}=\frac{G}{L}\sum_{s=1}^{t}\gamma^{s}\leq\frac{\gamma G}{(1-\gamma)L},$
where the second inequality uses Equation 139 and the third one use Lemma 2.7. Now we examine the rest statements for the $t$-th round of the algorithm. By Equation 96, we have
$\eta(f(\mathbf{x}_{t})-f(\mathbf{x}_{t-1}))$ (115)
$\leq$ $\underbrace{\eta\langle\nabla f(\mathbf{x}_{t})-\nabla f(\mathbf{y}_{t-1}),\mathbf{y}_{t}-\mathbf{x}_{t-1}\rangle}_{B_{t}}+\underbrace{\eta\langle\nabla f(\mathbf{x}_{t})-\nabla f(\mathbf{y}_{t}),\mathbf{x}_{t}-\mathbf{y}_{t}\rangle}_{D_{t}}$
$-\frac{1}{2}\left\|\mathbf{x}_{t}-\mathbf{y}_{t}\right\|^{2}-\frac{1}{2}\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|^{2}-\frac{1}{2}\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|^{2}.$
To leverage the effective smoothness property, we need to carefully control the distance quantity, namely $\left\|\mathbf{x}_{t}-\mathbf{y}_{t}\right\|$ and $\left\|\mathbf{x}_{t}-\mathbf{y}_{t-1}\right\|$. Based on the bound of $\left\|\mathbf{y}_{t}-\mathbf{y}_{t-1}\right\|$, by Equations 139 and 2.7 we have
$\left\|\mathbf{x}_{t}-\mathbf{y}_{t}\right\|\leq$ $\eta\left\|\nabla f(\mathbf{y}_{t})-\nabla f(\mathbf{y}_{t-1})\right\|_{*}\leq\eta L\left\|\mathbf{y}_{t}-\mathbf{y}_{t-1}\right\|$ (116)
$\leq$ $\frac{G}{L}\sum_{s=2}^{t+1}\gamma^{s}\leq\frac{\gamma^{2}G}{(1-\gamma)L}<\frac{G}{L},$
and
$\left\|\mathbf{x}_{t}-\mathbf{y}_{t-1}\right\|\leq\left\|\mathbf{x}_{t}-\mathbf{y}_{t}\right\|+\left\|\mathbf{y}_{t}-\mathbf{y}_{t-1}\right\|\leq\frac{(\gamma^{2}+\gamma)G}{(1-\gamma)L}<\frac{G}{L}.$ (117)
With these prepared, we invoke Lemma D.2 and choose $\beta=2\gamma^{2}/(1-2\gamma)$ therein to derive
$B_{t}\leq$ $\frac{1-2\gamma}{4}\left\|\mathbf{x}_{t}-\mathbf{y}_{t-1}\right\|^{2}+\frac{\gamma^{2}}{1-2\gamma}\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|^{2}$ (118)
$\leq$ $\frac{1-2\gamma}{2}\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|^{2}+\frac{1-2\gamma}{2}\left\|\mathbf{x}_{t-1}-\mathbf{y}_{t-1}\right\|^{2}+\frac{1}{2}\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|^{2},$
where we use the value range requirement of $\gamma\in(0,1/3]$. Similarly, choosing $\beta=\gamma$ in Lemma D.2 gives
$D_{t}=\eta\langle\nabla f(\mathbf{x}_{t})-\nabla f(\mathbf{y}_{t}),\mathbf{x}_{t}-\mathbf{y}_{t}\rangle\leq\gamma\left\|\mathbf{x}_{t}-\mathbf{y}_{t}\right\|^{2}.$ (119)
Adding (118) and (119) to (115), we obtain
$\eta(f(\mathbf{x}_{t})-f(\mathbf{x}_{t-1}))$ (120)
$\leq$ $\frac{1-2\gamma}{2}\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|^{2}+\frac{1-2\gamma}{2}\left\|\mathbf{x}_{t-1}-\mathbf{y}_{t-1}\right\|^{2}+\frac{1}{2}\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|^{2}+\gamma\left\|\mathbf{x}_{t}-\mathbf{y}_{t}\right\|^{2}$
$-\frac{1}{2}\left\|\mathbf{x}_{t}-\mathbf{y}_{t}\right\|^{2}-\frac{1}{2}\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|^{2}-\frac{1}{2}\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|^{2}$
$\leq$ $\frac{1-2\gamma}{2}\left\|\mathbf{x}_{t-1}-\mathbf{y}_{t-1}\right\|^{2}-\frac{1-2\gamma}{2}\left\|\mathbf{x}_{t}-\mathbf{y}_{t}\right\|^{2}.$
By induction, it’s easy to verify that the above relation holds for all iterations before the current one. Summing them up, we have
$\eta(f(\mathbf{x}_{t-1})-f(\mathbf{x}_{0}))$ $=\sum_{s=1}^{t-1}\eta(f(\mathbf{x}_{s})-f(\mathbf{x}_{s-1}))$ (121)
$\leq\sum_{s=1}^{t-1}\frac{1-2\gamma}{2}\left\|\mathbf{x}_{s-1}-\mathbf{y}_{s-1}\right\|^{2}-\frac{1-2\gamma}{2}\left\|\mathbf{x}_{s}-\mathbf{y}_{s}\right\|^{2}$
$\overset{\mathbf{x}_{0}=\mathbf{y}_{0}}{=}-\frac{1-2\gamma}{2}\left\|\mathbf{x}_{t-1}-\mathbf{y}_{t-1}\right\|^{2}.$
Combining (120) with (121), we obtain the following bound for the suboptimality gap of $\mathbf{x}_{t}$:
$\eta(f(\mathbf{x}_{t})-f(\mathbf{x}_{0}))\leq-\frac{1-2\gamma}{2}\left\|\mathbf{x}_{t}-\mathbf{y}_{t}\right\|^{2}\leq 0.$ (122)
Obviously, we have $f(\mathbf{x}_{t})-f^{*}\leq f(\mathbf{x}_{0})-f^{*}$ as $\eta>0$. Applying Lemma D.3 yields $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G$. It remains to show the suboptimality as well as the gradient bound for $\mathbf{y}_{t}$. By Equation 96, we have
$\eta(f(\mathbf{y}_{t})-f(\mathbf{x}_{t-1}))$ (123)
$\leq$ $\underbrace{\eta\langle\nabla f(\mathbf{y}_{t})-\nabla f(\mathbf{x}_{t-1}),\mathbf{y}_{t}-\mathbf{x}_{t-1}\rangle}_{E_{t}}+\underbrace{\eta\langle\nabla f(\mathbf{x}_{t-1})-\nabla f(\mathbf{y}_{t-1}),\mathbf{y}_{t}-\mathbf{x}_{t-1}\rangle}_{F_{t}}$
$-\frac{1}{2}\left\|\mathbf{x}_{t}-\mathbf{y}_{t}\right\|^{2}-\frac{1}{2}\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|^{2}-\frac{1}{2}\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|^{2}.$
Before using Lemma D.2 to bound the terms $E_{t}$ and $F_{t}$, we need to verify the applicability of Lemma 2.7. By Lemma E.1, we have
$\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|\leq\eta\left\|\nabla f(\mathbf{y}_{t-1})\right\|_{*}\leq\frac{\gamma G}{L}\leq\frac{G}{L}.$ (124)
Choosing $\beta=\gamma$ in Lemma D.2, $E_{t}$ can be bounded as follows:
$E_{t}=\eta\langle\nabla f(\mathbf{y}_{t})-\nabla f(\mathbf{x}_{t-1}),\mathbf{y}_{t}-\mathbf{x}_{t-1}\rangle\leq\gamma\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|^{2}.$ (125)
Similar routine should be done for $F_{t}$. By Equations 139 and 2.7, we have
$\left\|\mathbf{x}_{t-1}-\mathbf{y}_{t-1}\right\|\leq$ $\eta\left\|\nabla f(\mathbf{y}_{t-1})-\nabla f(\mathbf{y}_{t-2})\right\|_{*}\leq\eta L\left\|\mathbf{y}_{t-1}-\mathbf{y}_{t-2}\right\|$ (126)
$\leq$ $\frac{G}{L}\sum_{s=1}^{t-1}\gamma^{s+1}\leq\frac{\gamma^{2}G}{(1-\gamma)L}<\frac{G}{L}.$
Choosing $\beta=\gamma^{2}/(1-2\gamma)$ in Lemma D.2, $F_{t}$ can be bounded as follows:
$F_{t}\leq\frac{1-2\gamma}{2}\left\|\mathbf{x}_{t-1}-\mathbf{y}_{t-1}\right\|^{2}+\frac{\gamma^{2}}{2(1-2\gamma)}\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|^{2}.$ (127)
Adding (125) and (127) to (123), we obtain
$\eta(f(\mathbf{y}_{t})-f(\mathbf{x}_{t-1}))$ (128)
$\leq$ $\gamma\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|^{2}+\frac{1-2\gamma}{2}\left\|\mathbf{x}_{t-1}-\mathbf{y}_{t-1}\right\|^{2}+\frac{\gamma^{2}}{2(1-2\gamma)}\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|^{2}$
$-\frac{1}{2}\left\|\mathbf{x}_{t}-\mathbf{y}_{t}\right\|^{2}-\frac{1}{2}\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|^{2}-\frac{1}{2}\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|^{2}$
$\leq$ $\frac{1-2\gamma}{2}\left\|\mathbf{x}_{t-1}-\mathbf{y}_{t-1}\right\|^{2}-\frac{1-2\gamma}{2}\left\|\mathbf{x}_{t}-\mathbf{y}_{t}\right\|^{2},$
where the last inequality is due to $\gamma^{2}/(1-2\gamma)+2\gamma\leq 1$ for any $\gamma\in(0,1/3]$. Adding (121) to the above inequality, we have
$\eta(f(\mathbf{y}_{t})-f(\mathbf{x}_{0}))\leq-\frac{1-2\gamma}{2}\left\|\mathbf{x}_{t}-\mathbf{y}_{t}\right\|^{2}\leq 0.$ (129)
We immediately see that $f(\mathbf{y}_{t})-f^{*}\leq f(\mathbf{x}_{0})-f^{*}$. Then $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G$ is a direct result based on Lemma D.3. Combining Part 1 with Part 2 finishes the proof. ∎
Proof of Equation 11.
Applying Equation 139 to the updates defined in (10), we have
$\forall\mathbf{x}\in\mathcal{X},\eta\langle\nabla f(\mathbf{y}_{t}),\mathbf{y}_{t}-\mathbf{x}\rangle\leq$ $B(\mathbf{x},\mathbf{x}_{t-1})-B(\mathbf{x},\mathbf{x}_{t})-B(\mathbf{x}_{t},\mathbf{y}_{t})-B(\mathbf{y}_{t},\mathbf{x}_{t-1})$ (130)
$+\eta\langle\nabla f(\mathbf{y}_{t})-\nabla f(\mathbf{y}_{t-1}),\mathbf{y}_{t}-\mathbf{x}_{t}\rangle.$
By Lemma G.2, we have $\left\|\mathbf{y}_{t}-\mathbf{y}_{t-1}\right\|\leq\frac{G}{L}\sum_{s=1}^{t}1/3^{s}\leq\frac{G}{2L}<\frac{G}{L}$ as we specify $\gamma=1/3$, which satisfies the condition in Lemma D.2. Hence, the following result holds:
$\eta\langle\nabla f(\mathbf{y}_{t})-\nabla f(\mathbf{y}_{t-1}),\mathbf{y}_{t}-\mathbf{x}_{t}\rangle\leq\frac{1}{12}\left\|\mathbf{y}_{t}-\mathbf{y}_{t-1}\right\|^{2}+\frac{1}{3}\left\|\mathbf{y}_{t}-\mathbf{x}_{t}\right\|^{2},$ (131)
where we set $\beta=2/3$ in Lemma D.2. Plugging this into (130) and recalling the property of Bregman divergence, we have
$\eta\langle\nabla f(\mathbf{y}_{t}),\mathbf{y}_{t}-\mathbf{x}\rangle\leq$ $B(\mathbf{x},\mathbf{x}_{t-1})-B(\mathbf{x},\mathbf{x}_{t})-\frac{1}{2}\left\|\mathbf{x}_{t}-\mathbf{y}_{t}\right\|^{2}-\frac{1}{2}\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|^{2}$ (132)
$+\frac{1}{6}\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|^{2}+\frac{1}{6}\left\|\mathbf{x}_{t-1}-\mathbf{y}_{t-1}\right\|^{2}+\frac{1}{3}\left\|\mathbf{y}_{t}-\mathbf{x}_{t}\right\|^{2}$
$\leq$ $B(\mathbf{x},\mathbf{x}_{t-1})-B(\mathbf{x},\mathbf{x}_{t})+\frac{1}{6}\left\|\mathbf{x}_{t-1}-\mathbf{y}_{t-1}\right\|^{2}-\frac{1}{6}\left\|\mathbf{x}_{t}-\mathbf{y}_{t}\right\|^{2},$
which holds for any $\mathbf{x}\in\mathcal{X}$. Telescoping sum, we obtain
$\sum_{t=1}^{T}\langle\eta\nabla f(\mathbf{y}_{t}),\mathbf{y}_{t}-\mathbf{x}\rangle\leq B(\mathbf{x},\mathbf{x}_{0})-B(\mathbf{x},\mathbf{x}_{T})-\frac{1}{6}\left\|\mathbf{x}_{T}-\mathbf{y}_{T}\right\|^{2}\leq B(\mathbf{x},\mathbf{x}_{0}).$ (133)
By Jensen’s Inequality, we derive the following convergence rate:
$f(\overline{\mathbf{y}}_{T})-f^{*}\leq\frac{1}{\eta T}\sum_{t=1}^{T}\langle\eta\nabla f(\mathbf{y}_{t}),\mathbf{y}_{t}-\mathbf{x}_{*}\rangle\leq\frac{B(\mathbf{x}_{*},\mathbf{x}_{0})}{\eta T},$ (134)
where $\overline{\mathbf{y}}_{T}=\sum_{t=1}^{T}\mathbf{y}_{t}/T$. ∎
Appendix H Analysis for Mirror Prox
The following lemma describes some basics of the main building blocks of the algorithms that require two prox-mapping per iteration (Korpelevich, 1976; Popov, 1980). The techniques primarily stem from the seminal work of Nemirovski (2004). For completeness, we provide the proof here.
Lemma H.1.
For any $\mathbf{g}_{1},\mathbf{g}_{2}\in\mathcal{E}^{*}$ and any $\mathbf{x}_{0}\in\mathcal{X}$ satisfying the following algorithmic dynamics:
$\mathbf{x}_{1}=\mathcal{P}_{\mathbf{x}_{0}}(\mathbf{g}_{1}),\quad\mathbf{x}_{2}=\mathcal{P}_{\mathbf{x}_{0}}(\mathbf{g}_{2}),$ (135)
we have
$\quad\forall\mathbf{x}\in\mathcal{X},\langle\mathbf{g}_{1},\mathbf{x}_{1}-\mathbf{x}\rangle\leq B(\mathbf{x},\mathbf{x}_{0})-B(\mathbf{x},\mathbf{x}_{1})-B(\mathbf{x}_{1},\mathbf{x}_{0}),$ (136)
$\quad\forall\mathbf{x}\in\mathcal{X},\langle\mathbf{g}_{2},\mathbf{x}_{2}-\mathbf{x}\rangle\leq B(\mathbf{x},\mathbf{x}_{0})-B(\mathbf{x},\mathbf{x}_{2})-B(\mathbf{x}_{2},\mathbf{x}_{0}),$ (137)
$\forall\mathbf{x}\in\mathcal{X},\langle\nabla f(\mathbf{x}_{1}),\mathbf{x}_{1}-\mathbf{x}\rangle\leq$ $B(\mathbf{x},\mathbf{x}_{0})-B(\mathbf{x},\mathbf{x}_{2})-B(\mathbf{x}_{2},\mathbf{x}_{1})-B(\mathbf{x}_{1},\mathbf{x}_{0})$ (138)
$-\langle\mathbf{g}_{1}-\mathbf{g}_{2},\mathbf{x}_{1}-\mathbf{x}_{2}\rangle+\langle\nabla f(\mathbf{x}_{1})-\mathbf{g}_{2},\mathbf{x}_{1}-\mathbf{x}\rangle,$
and the continuity of the prox-mapping is characterized by
$\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|=\left\|\mathcal{P}_{\mathbf{x}_{0}}(\mathbf{g}_{1})-\mathcal{P}_{\mathbf{x}_{0}}(\mathbf{g}_{2})\right\|\leq\left\|\mathbf{g}_{1}-\mathbf{g}_{2}\right\|_{*}.$ (139)
Proof.
Since $\psi(\cdot)$ is 1-strongly convex, we can leverage an existing tool (Nemirovski, 2004, Lemma 2.1), which exploits the continuity of the operator $\mathcal{P}_{\mathbf{x}}(\cdot)$, to show (139). (136) and (137) is immediately proven after using the first-order optimality condition for projection under Bregman divergence (Nemirovski, 2004, Lemma 3.1) on both updates. Taking $\mathbf{x}=\mathbf{x}_{2}$ in (136) and adding (137), we obtain
$\langle\mathbf{g}_{2},\mathbf{x}_{2}-\mathbf{x}\rangle+\langle\mathbf{g}_{1},\mathbf{x}_{1}-\mathbf{x}_{2}\rangle\leq B(\mathbf{x},\mathbf{x}_{0})-B(\mathbf{x},\mathbf{x}_{2})-B(\mathbf{x}_{2},\mathbf{x}_{1})-B(\mathbf{x}_{1},\mathbf{x}_{0}),\quad\forall\mathbf{x}\in\mathcal{X}.$ (140)
Rearrange the terms as below:
$\forall\mathbf{x}\in\mathcal{X},\langle\mathbf{g}_{2},\mathbf{x}_{1}-\mathbf{x}\rangle\leq$ $B(\mathbf{x},\mathbf{x}_{0})-B(\mathbf{x},\mathbf{x}_{2})-B(\mathbf{x}_{2},\mathbf{x}_{1})-B(\mathbf{x}_{1},\mathbf{x}_{0})$ (141)
$-\langle\mathbf{g}_{1}-\mathbf{g}_{2},\mathbf{x}_{1}-\mathbf{x}_{2}\rangle.$
Adding $\langle\nabla f(\mathbf{x}_{1})-\mathbf{g}_{2},\mathbf{x}_{1}-\mathbf{x}\rangle$ to both sides of this inequality completes the proof. ∎
Lemma H.2.
Under Assumptions 2.5 and 3.3, consider the updates in (12). If $0<\eta\leq 1/2\ell(2G)$ where $G:=\sup\left\{\alpha\in\mathbb{R}_{+}|\alpha^{2}\leq 2(f(\mathbf{x}_{0})-f^{*})\ell(2\alpha)\right\}$, the following statements holds for all $t\in\mathbb{N}_{+}$:
  1. 1.
    (descent property) $\quad\quad f(\mathbf{x}_{t})\leq f(\mathbf{x}_{t-1}),f(\mathbf{y}_{t})\leq f(\mathbf{x}_{t-1})$;
  2. 2.
    (bounded gradients) $\quad\left\|\nabla f(\mathbf{y}_{t})\right\|_{*}\leq G<\infty,\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G<\infty$.
Proof.
We prove this lemma by induction.
Part 1. Base Case
We consider the case where $t=1$. Similar to the proof of Lemma E.2, we argue that $f(\mathbf{y}_{0})-f^{*}=f(\mathbf{x}_{0})-f^{*}<\infty$. By Lemma 3.4, we have $\left\|\nabla f(\mathbf{y}_{0})\right\|_{*}=\left\|\nabla f(\mathbf{y}_{0})\right\|_{*}\leq G<\infty$. Now we look at the sequence $\left\{\mathbf{y}_{t}\right\}_{t\in\mathbb{N}}$ that mirror prox maintains, whose dynamics are exactly like mirror descent. For convenience, we denote $\ell(2G)$ by $L$. As $\eta\leq 1/(2L)\leq 1/L$, we can apply Lemma E.2 to conclude that $f(\mathbf{y}_{1})\leq f(\mathbf{x}_{0})=f(\mathbf{y}_{0})$ and $\left\|\nabla f(\mathbf{y}_{1})\right\|_{*}\leq G$. Next, we apply Lemma H.3 to the first round of the mirror prox algorithm. It’s straightforward to verify that the conditions are satisfied for the case where $s=1$. Hence, we have $f(\mathbf{x}_{1})\leq f(\mathbf{x}_{0})$. By Lemma D.3, we conclude that $\left\|\nabla f(\mathbf{x}_{1})\right\|_{*}\leq G$.
Part 2. Induction Step
Suppose the statements hold for all $t\leq s-1$ where $s>2$. So we have $\left\|\nabla f(\mathbf{y}_{s-1})\right\|_{*}\leq G$ and $\left\|\nabla f(\mathbf{x}_{s-1})\right\|_{*}\leq G$. We also have $f(\mathbf{x}_{s-1})\leq f(\mathbf{x}_{s-2})\leq f(\mathbf{x}_{0})$. Now we consider the case where $t=s$. Similar to the base case, we invoke Lemma E.2 to immediately conclude that $f(\mathbf{y}_{s})\leq f(\mathbf{x}_{s-1})\leq f(\mathbf{x}_{0})$ and $\left\|\nabla f(\mathbf{y}_{s})\right\|_{*}\leq G$. Next, we prove the other half of the statements. Note that the conditions in Lemma H.3 are all satisfied by the previous derivations. So we invoke Lemma H.3 to deduce that $f(\mathbf{x}_{s})\leq f(\mathbf{x}_{s-1})$, which immediately implies $f(\mathbf{x}_{s})\leq f(\mathbf{x}_{0})$. Then $\left\|\nabla f(\mathbf{x}_{s})\right\|_{*}\leq G$ is a direct result based on Lemma D.3. Combining Part 1 with Part 2 finishes the proof. ∎
Lemma H.3.
Under Assumptions 2.5 and 3.3, consider the following updates for any $s\in\mathbb{N}_{+}$:
$\mathbf{y}_{s}=\mathcal{P}_{\mathbf{x}_{s-1}}(\eta\nabla f(\mathbf{x}_{s-1})),\quad\mathbf{x}_{s}=\mathcal{P}_{\mathbf{x}_{s-1}}(\eta\nabla f(\mathbf{y}_{s})),$ (142)
which satisfies $0<\eta\leq 1/2\ell(2G)$ for some $G\in\mathbb{R}_{+}$. If $\left\|\nabla f(\mathbf{x}_{s-1})\right\|_{*}\leq G$ and $\left\|\nabla f(\mathbf{y}_{s})\right\|_{*}\leq G$ hold, then we have $f(\mathbf{x}_{s})\leq f(\mathbf{x}_{s-1})$.
Proof.
Notice that
$\langle\nabla f(\mathbf{x}_{s}),\mathbf{x}_{s}-\mathbf{x}_{s-1}\rangle=$ $\underbrace{\langle\nabla f(\mathbf{y}_{s}),\mathbf{x}_{s}-\mathbf{y}_{s}\rangle}_{A_{s}}+\underbrace{\langle\nabla f(\mathbf{x}_{s})-\nabla f(\mathbf{y}_{s}),\mathbf{x}_{s}-\mathbf{y}_{s}\rangle}_{B_{s}}$ (143)
$+\underbrace{\langle\nabla f(\mathbf{x}_{s-1}),\mathbf{y}_{s}-\mathbf{x}_{s-1}\rangle}_{C_{s}}$
$+\underbrace{\langle\nabla f(\mathbf{x}_{s})-\nabla f(\mathbf{x}_{s-1}),\mathbf{y}_{s}-\mathbf{x}_{s-1}\rangle}_{D_{s}}.$
We bound the four terms separately. $A_{s}$ and $C_{s}$ can be bounded by the property of prox-mapping (Nemirovski, 2004, Lemma 3.1). Choosing $\mathbf{x}_{1}=\mathbf{y}_{s},\mathbf{x}_{0}=\mathbf{x}_{s-1},\mathbf{g}_{1}=\eta\nabla f(\mathbf{x}_{s-1}),\mathbf{x}=\mathbf{x}_{s-1}$ in (136), we obtain
$\eta C_{s}=\langle\eta\nabla f(\mathbf{x}_{s-1}),\mathbf{y}_{s}-\mathbf{x}_{s-1}\rangle\leq-B(\mathbf{x}_{s-1},\mathbf{y}_{s})-B(\mathbf{y}_{s},\mathbf{x}_{s-1}).$ (144)
Choosing $\mathbf{x}_{2}=\mathbf{x}_{s},\mathbf{x}_{0}=\mathbf{x}_{s-1},\mathbf{g}_{2}=\eta\nabla f(\mathbf{y}_{s}),\mathbf{x}=\mathbf{y}_{s}$ in (137), we obtain
$\eta A_{s}=\langle\eta\nabla f(\mathbf{y}_{s}),\mathbf{x}_{s}-\mathbf{y}_{s}\rangle\leq B(\mathbf{y}_{s},\mathbf{x}_{s-1})-B(\mathbf{y}_{s},\mathbf{x}_{s})-B(\mathbf{x}_{s},\mathbf{x}_{s-1}).$ (145)
Adding them together, we have
$\eta A_{s}+\eta C_{s}\leq$ $-B(\mathbf{x}_{s-1},\mathbf{y}_{s})-B(\mathbf{y}_{s},\mathbf{x}_{s})-B(\mathbf{x}_{s},\mathbf{x}_{s-1})$ (146)
$\leq$ $-\frac{1}{2}\left\|\mathbf{x}_{s-1}-\mathbf{y}_{s}\right\|^{2}-\frac{1}{2}\left\|\mathbf{y}_{s}-\mathbf{x}_{s}\right\|^{2}-\frac{1}{2}\left\|\mathbf{x}_{s}-\mathbf{x}_{s-1}\right\|^{2}.$
For $B_{s}$ and $D_{s}$, we shall use the effective smoothness property. For convenience, denote by $L:=\ell(2G)$ as we did in Lemma 2.7. Observe that
$\left\|\mathbf{x}_{s}-\mathbf{x}_{s-1}\right\|\leq\eta\left\|\nabla f(\mathbf{y}_{s})\right\|_{*}\leq\frac{G}{2L}\leq\frac{G}{L}$ (147)
by Lemma E.1 and the given condition. Then we call Lemma D.2 to bound $\eta D_{s}$ as follows:
$\eta D_{s}=\eta\langle\nabla f(\mathbf{x}_{s})-\nabla f(\mathbf{x}_{s-1}),\mathbf{y}_{s}-\mathbf{x}_{s-1}\rangle\leq\frac{1}{8}\left\|\mathbf{x}_{s}-\mathbf{x}_{s-1}\right\|^{2}+\frac{1}{2}\left\|\mathbf{y}_{s}-\mathbf{x}_{s-1}\right\|^{2}$ (148)
As for $B_{s}$, we still need to verify that the distance between $\mathbf{x}_{s}$ and $\mathbf{y}_{s}$ is small enough in order to apply Lemma 2.7. We utilize (139) to show that
$\left\|\mathbf{x}_{s}-\mathbf{y}_{s}\right\|=\left\|\mathcal{P}_{\mathbf{x}_{s-1}}(\eta\nabla f(\mathbf{y}_{s}))-\mathcal{P}_{\mathbf{x}_{s-1}}(\eta\nabla f(\mathbf{x}_{s-1}))\right\|\leq\eta\left\|\nabla f(\mathbf{y}_{s})-\nabla f(\mathbf{x}_{s-1})\right\|_{*}.$ (149)
Then by Lemma E.1 and $\left\|\nabla f(\mathbf{x}_{s-1})\right\|_{*}\leq G$, it follows
$\left\|\mathbf{y}_{s}-\mathbf{x}_{s-1}\right\|\leq\eta\left\|\nabla f(\mathbf{x}_{s-1})\right\|_{*}\leq\frac{G}{2L}\leq\frac{G}{L}.$ (150)
Using Lemma 2.7, we deduce that
$\left\|\mathbf{x}_{s}-\mathbf{y}_{s}\right\|\leq\eta\left\|\nabla f(\mathbf{y}_{s})-\nabla f(\mathbf{x}_{s-1})\right\|_{*}\leq\eta L\left\|\mathbf{y}_{s}-\mathbf{x}_{s-1}\right\|\leq\frac{G}{4L}\leq\frac{G}{L}.$ (151)
Utilizing Lemma 2.7 again along with Cauchy-Schwarz Inequality, we have
$\eta B_{s}=$ $\eta\langle\nabla f(\mathbf{x}_{s})-\nabla f(\mathbf{y}_{s}),\mathbf{x}_{s}-\mathbf{y}_{s}\rangle$ (152)
$\leq$ $\eta\left\|\nabla f(\mathbf{x}_{s})-\nabla f(\mathbf{y}_{s})\right\|_{*}\left\|\mathbf{x}_{s}-\mathbf{y}_{s}\right\|\leq\eta L\left\|\mathbf{x}_{s}-\mathbf{y}_{s}\right\|^{2}\leq\frac{1}{2}\left\|\mathbf{x}_{s}-\mathbf{y}_{s}\right\|^{2}.$
Adding up (146), (148) and (152), we obtain
$\eta(A_{s}+B_{s}+C_{s}+D_{s})\leq$ $-\frac{1}{2}\left\|\mathbf{x}_{s-1}-\mathbf{y}_{s}\right\|^{2}-\frac{1}{2}\left\|\mathbf{y}_{s}-\mathbf{x}_{s}\right\|^{2}-\frac{1}{2}\left\|\mathbf{x}_{s}-\mathbf{x}_{s-1}\right\|^{2}$ (153)
$+\frac{1}{8}\left\|\mathbf{x}_{s}-\mathbf{x}_{s-1}\right\|^{2}+\frac{1}{2}\left\|\mathbf{y}_{s}-\mathbf{x}_{s-1}\right\|^{2}+\frac{1}{2}\left\|\mathbf{x}_{s}-\mathbf{y}_{s}\right\|^{2}$
$=$ $-\frac{3}{8}\left\|\mathbf{x}_{s}-\mathbf{x}_{s-1}\right\|^{2}\leq 0.$
Since $\eta>0$, it directly yields
$f(\mathbf{x}_{s})-f(\mathbf{x}_{s-1})\leq\langle\nabla f(\mathbf{x}_{s}),\mathbf{x}_{s}-\mathbf{x}_{s-1}\rangle=A_{s}+B_{s}+C_{s}+D_{s}\leq 0.$ (154)
Proof of Equation 13.
By Lemma H.2, we have $\left\|\nabla f(\mathbf{y}_{t})\right\|_{*}\leq G<\infty,\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G<\infty$ for all $t\in\mathbb{N}_{+}$. Since the bound for $\nabla f(\mathbf{x}_{0})=\nabla f(\mathbf{y}_{0})$ naturally holds, so we obtain
$\max\left\{\left\|\nabla f(\mathbf{y}_{t})\right\|_{*},\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\right\}\leq G<\infty,\quad\forall t\in\mathbb{N}.$ (155)
Now it suffices to show the convergence rate is at the order of $O(1/T)$. Choosing $\mathbf{x}_{0}=\mathbf{x}_{t-1},\mathbf{x}_{1}=y_{t},\mathbf{g}_{1}=\eta\nabla f(\mathbf{x}_{t-1}),\mathbf{x}_{2}=\mathbf{x}_{t},\mathbf{g}_{2}=\eta\nabla f(\mathbf{y}_{t})$ in Equation 139, we have
$\forall\mathbf{x}\in\mathcal{X},\langle\eta\nabla f(\mathbf{y}_{t}),\mathbf{y}_{t}-\mathbf{x}\rangle\leq$ $B(\mathbf{x},\mathbf{x}_{t-1})-B(\mathbf{x},\mathbf{x}_{t})-B(\mathbf{x}_{t},\mathbf{y}_{t})-B(\mathbf{y}_{t},\mathbf{x}_{t-1})$ (156)
$-\eta\langle\nabla f(\mathbf{x}_{t-1})-\nabla f(\mathbf{y}_{t}),\mathbf{y}_{t}-\mathbf{x}_{t}\rangle.$
By Lemma E.1 and (155), we have
$\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|\leq\eta\left\|\nabla f(\mathbf{x}_{t-1})\right\|_{*}\leq\frac{G}{2L}\leq\frac{G}{L}.$ (157)
As $\eta\leq 1/(2L)$, we set $\beta=1$ in Lemma D.2 to derive
$\eta\langle\nabla f(\mathbf{y}_{t})-\nabla f(\mathbf{x}_{t-1}),\mathbf{y}_{t}-\mathbf{x}_{t}\rangle\leq\frac{1}{8}\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|^{2}+\frac{1}{2}\left\|\mathbf{y}_{t}-\mathbf{x}_{t}\right\|^{2}.$ (158)
Combine (156) with (158):
$\forall\mathbf{x}\in\mathcal{X},\langle\eta\nabla f(\mathbf{y}_{t}),\mathbf{y}_{t}-\mathbf{x}\rangle\leq$ $B(\mathbf{x},\mathbf{x}_{t-1})-B(\mathbf{x},\mathbf{x}_{t})-\frac{1}{2}\left\|\mathbf{x}_{t}-\mathbf{y}_{t}\right\|^{2}-\frac{1}{2}\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|^{2}$ (159)
$+\frac{1}{8}\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|^{2}+\frac{1}{2}\left\|\mathbf{y}_{t}-\mathbf{x}_{t}\right\|^{2}$
$\leq$ $B(\mathbf{x},\mathbf{x}_{t-1})-B(\mathbf{x},\mathbf{x}_{t})-\frac{3}{8}\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|^{2},$
where we also use the property of Bregman divergence. By convexity of $f$, we have
$\forall\mathbf{x}\in\mathcal{X},f(\mathbf{y}_{t})-f(\mathbf{x})\leq\langle\nabla f(\mathbf{y}_{t}),\mathbf{y}_{t}-\mathbf{x}\rangle\leq\frac{B(\mathbf{x},\mathbf{x}_{t-1})-B(\mathbf{x},\mathbf{x}_{t})}{\eta}.$ (160)
Telescoping and taking $\mathbf{x}=\mathbf{x}_{*}$ on both sides of the inequality, we obtain
$\sum_{t=1}^{T}[f(\mathbf{y}_{t})-f^{*}]\leq\frac{B(\mathbf{x}_{*},\mathbf{x}_{0})-B(\mathbf{x}_{*},\mathbf{x}_{T})}{\eta}\leq\frac{B(\mathbf{x}_{*},\mathbf{x}_{0})}{\eta}.$ (161)
Since the output is the average of iterates, i.e., $\overline{\mathbf{y}}_{T}=\sum_{t=1}^{T}\mathbf{y}_{t}/T$, the proof is finished by Jensen’s inequality and the convexity of $f$. ∎
Appendix I Analysis of Stochastic Mirror Descent
The lemmas in this section hold under the conditions of Theorem 4.2. For brevity, we do not restate this point.
Lemma I.1.
Suppose $\eta_{s}\leq\frac{1}{2L}$ and that $\left\|\nabla f(\mathbf{x}_{s})\right\|_{*}\leq G,\left\|\mathbf{g}_{s}\right\|_{*}\leq\frac{G}{\eta_{s+1}L}$ for all $0\leq s\leq t-1$. For any $0<\delta<1$, with probability at least $1-\delta$, for each $t\in[T]$, it holds that
$f(\mathbf{x}_{t})-f(\mathbf{x}_{0})\leq 16\sigma^{2}\log\left(\frac{2}{\delta}\right)\sum_{s=1}^{t}\frac{\eta_{s}^{2}}{\sum_{k=s}^{t}\eta_{k}}.$ (162)
Proof.
This lemma is built upon Liu and Zhou (2025b, Lemma 4.3). Under the conditions of $\left\|\nabla f(\mathbf{x}_{s})\right\|_{*}\leq G,\left\|\mathbf{g}_{s}\right\|_{*}\leq\frac{G}{\eta L}$, in each step $t$ where we utilize $\ell*$-smoothness, it degenerates to the standard $L$-smoothness in Liu and Zhou (2025b). Thus, the proof of Lemma 4.3 in Liu and Zhou (2025b) remains valid. Substituting arbitrary $\mathbf{x}=\mathbf{x}_{0}$, noting that $M=0$ and $1\leq 2\log\left(\frac{2}{\delta}\right)$ finishes the proof. ∎
Lemma I.2 (Bounded suboptimality gaps).
With probability at least $1-\frac{\delta}{2}$, for all $t\in[T]$, we have
$f(\mathbf{x}_{t})-f^{*}\leq f(\mathbf{x}_{0})-f^{*}+\frac{64\sigma}{\sqrt{\delta}}.$ (163)
In the meantime, it holds that
$\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G,\ \left\|\mathbf{g}_{t}\right\|_{*}\leq\frac{G}{\eta_{t+1}L},\ \left\|\boldsymbol{\epsilon}_{t}\right\|_{*}\leq\frac{G}{2\eta_{s+1}L}\quad\forall t-1\in[T].$ (164)
Proof.
The essence underlying this lemma is to enforce the happening of events $\left\{A_{t}\right\}_{t=0}^{T-1}$ and $\left\{B_{t}\right\}_{t=0}^{T-1}$. For the latter, which governs the gradient estimates, we invoke Chebyshev’s inequality under Assumption 4.1:
$\mathbb{P}\left(\bigcup_{t=0}^{T-1}B_{t}\right)\leq\sum_{t=0}^{T-1}\mathbb{P}\left(\left\|\boldsymbol{\epsilon}_{t}\right\|_{*}>\frac{G}{2\eta_{s+1}L}\right)\leq\frac{4\eta^{2}L^{2}\sigma^{2}}{G^{2}}\leq\frac{4\eta^{2}G^{4}\delta}{128^{2}G^{2}}\leq\frac{\delta}{4},$ (165)
where the last inequality is due to $\eta\leq 32/G$ together with the following relation induced by Lemma D.3:
$G^{2}=2LF=2L\left(f(\mathbf{x}_{0})-f^{*}+\frac{64\sigma}{\sqrt{\delta}}\right)\geq\frac{128L\sigma}{\sqrt{\delta}}.$ (166)
On the other hand, for the former $\left\{A_{t}\right\}_{t=0}^{T-1}$, which governs the suboptimality gaps, we seek help from Lemma I.1. We proceed by induction to bound the failure probability of events $\left\{A_{t}\right\}_{t=0}^{T-1}$. For $t=0$, $A_{0}$ trivially holds. By Lemma D.3, we have $\left\|\nabla f(\mathbf{x}_{0})\right\|_{*}\leq G$. Since $B_{0}$ holds with probability at least $1-\frac{\delta}{4T}$ according to (165), we have
$\eta_{1}\left\|\mathbf{g}_{0}\right\|_{*}\leq\eta_{1}\left\|\nabla f(\mathbf{x}_{0})\right\|_{*}+\eta_{1}\left\|\boldsymbol{\epsilon}_{0}\right\|_{*}\leq\frac{G}{2L}+\frac{\eta_{1}G}{2\eta_{1}L}=\frac{G}{L}.$
Then, we invoke Lemma I.1 to derive
$f(\mathbf{x}_{1})-f(\mathbf{x}_{0})\leq\frac{16\eta\sigma^{2}}{\sqrt{T}}\log\left(\frac{8T}{\delta}\right)=\frac{16\sigma}{\sqrt{T}}\sqrt{\log\left(\frac{8T}{\delta}\right)}\quad\text{w.p.}\geq 1-\frac{\delta}{4T},$
[…]
$f(\mathbf{x}_{t})-f(\mathbf{x}_{0})\leq\frac{16\sigma}{\sqrt{T}}\sqrt{\log\left(\frac{8T}{\delta}\right)}\sum_{s=1}^{t}\frac{\eta}{\sqrt{T}(t-s+1)}\leq 16\sigma\left(1+\log(t)\right)\sqrt{\frac{\log\left(\frac{8T}{\delta}\right)}{T}}$
$\leq$ $32\sigma\log(T)\sqrt{\frac{\log\left(\frac{8T}{\delta}\right)}{T}}\leq\frac{64\sigma}{e}\sqrt{\log\left(\frac{8}{\delta}\right)+3}\leq 37\sigma\sqrt{\log\left(\frac{8}{\delta}\right)}\leq\frac{64\sigma}{\sqrt{\delta}},$
[…]
With the above preparations, we are ready to prove the main theorem.
Proof of Theorem 4.2.
By (164) in Lemma I.2, we deduce that throughout the trajectory of (14), we can view the behavior of $\ell*$-smooth functions as the standard smooth functions with effective smoothness parameter $L$. In this way, the conditions for Lemma I.1 are satisfied. According to Eq.(30) in Liu and Zhou (2025b), with probability at least $1-\frac{\delta}{2}$, we have
$f(\mathbf{x}_{T})-f^{*}\leq\frac{4B(\mathbf{x}_{*},\mathbf{x}_{0})}{\sum_{t=1}^{T}\eta_{t}}+4\sigma^{2}\left(1+2\log\left(\frac{4}{\delta}\right)\right)\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\sum_{s=t}^{T}\eta_{s}}.$
Note that the above relation must build upon the premise of Lemma I.2, i.e., the occurrence of events $\cup_{t=0}^{T-1}A_{t}$ and $\cup_{t=0}^{T-1}B_{t}$. We take a union bound among these events to obtain
$f(\mathbf{x}_{T})-f^{*}\leq\frac{4B(\mathbf{x}_{*},\mathbf{x}_{0})}{\sum_{t=1}^{T}\eta_{t}}+16\sigma^{2}\log\left(\frac{4}{\delta}\right)\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\sum_{s=t}^{T}\eta_{s}},\quad\text{w.p.}\geq 1-\delta.$
Noting that $\eta_{t}=\min\left\{\frac{1}{2L},\frac{64}{G},\sqrt{\frac{B(\mathbf{x}_{*},\mathbf{x}_{0})}{\sigma^{2}\log\left(\frac{1}{\delta}\right)\log(T)}}\right\}$, we finally arrive at
$f(\mathbf{x}_{T})-f^{*}\leq$ $O\left(\frac{LB(\mathbf{x}_{*},\mathbf{x}_{0})}{T}+\sigma\sqrt{\frac{B(\mathbf{x}_{*},\mathbf{x}_{0})\log\left(\frac{1}{\delta}\right)\log(T)}{T}}+\frac{GB(\mathbf{x}_{*},\mathbf{x}_{0})}{\sqrt{T}}\right)$
[…]
which holds sufficiently large $T$ and small $\delta$. ∎
Appendix J Extension for Stochastic Convex Optimization Under New Noise Model
In this section, we provide new convergence results for SCO under $\ell*$-smoothness and a newly proposed noise model (Assumption J.1). Section J.1 introduces the new noise condition. Section J.2 gives the main result, providing a high-probability anytime convergence guarantee. Section J.3 compares it with other popular noise assumptions, where one can see why this new formulation can be meaningful and more expressive. Section J.4 provides the omitted proof.
J.1 Generalized Bounded Noise Model
We formally introduce the generalized bounded noise assumption as below.
Assumption J.1.
For all $t\in\mathbb{N}$, the stochastic gradients are unbiased:
$\mathbb{E}_{t-1}[\boldsymbol{\epsilon}_{t}]=0,$ (167)
where the expectation $\mathbb{E}_{t-1}$ is conditioned on the past stochasticity $\left\{\boldsymbol{\epsilon}_{s}\right\}_{s=0}^{t-1}$. The noise satisfies the generalized bounded condition w.r.t. a non-decreasing polynomial function $\sigma:\mathbb{R}_{+}\mapsto\mathbb{R}_{++}$:
$\left\|\boldsymbol{\epsilon}_{t}\right\|_{*}\leq\sigma(\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}),\quad\text{a.s.}$ (168)
Moreover, the degree of the polynomial $\sigma$ is finite: $\operatorname{deg}(\sigma)<+\infty$.
This generalized bounded noise [^7] The noise is bounded when the gradient is bounded but can become unbounded if the gradient explodes. assumption is inspired by the affine variance condition initially proposed by Bottou et al. (2018), where the noise level depends on the norm of the true gradient and is captured via an affine function. Assumption J.1 encompasses the uniformly bounded noise (Zhang et al., 2020b) and affine noise (Liu et al., 2023b; Hong and Lin, 2024) condition with $\sigma(\alpha)\equiv\sigma$ and $\sigma(\alpha)\equiv\sigma_{0}+\sigma_{1}\alpha$, respectively. Moreover, we underline that Assumption J.1 is never stronger than the commonly assumed finite moment or sub-Gaussian noise condition, with thorough discussions deferred to Section J.3.
For theoretical analysis, we also need to following bounded domain assumption, which is also widely adopted in the context of mirror descent (Zhang et al., 2023, 2024b; Yu et al., 2024; Bai et al., 2025).
Assumption J.2.
The domain $\mathcal{X}$ is bounded in the sense that $\max_{\mathbf{x}\in\mathcal{X}}\psi(\mathbf{x})-\min_{\mathbf{x}\in\mathcal{X}}\psi(\mathbf{x})\leq D^{2}$.
Note that Assumption J.2 does not imply $\sup_{\mathbf{x},\mathbf{y}}B(\mathbf{x},\mathbf{y})\leq D^{2}$, which is used in Eldowa and Paudice (2024, Lemma 9) to simplify the analysis. This stronger assumption is rarely used and can not be satisfied by the simplest instance of Bregman divergence induced by negative entropy functions.
J.2 Theoretical Guarantees
Based on our new noise model, establishing a theoretical guarantee is highly challenging for the sake of the following two quantities that need to be controlled.
Suboptimality gap   Due to the inherent stochasticity, the sequence $\left\{f(\mathbf{x}_{t})-f^{*}\right\}_{t\in\mathbb{N}}$ may become unbounded. Recall from deterministic optimization, we directly link suboptimality gaps to the last-iterate convergence. However, stochastic gradient descent (SGD) (Robbins and Monro, 1951) primarily focuses on the average-iterate, and therefore, such a link does not hold. A fruitful line of work (Shamir and Zhang, 2013; Harvey et al., 2019; Orabona, 2020; Jain et al., 2021) explore the last-iterate convergence of SGD but often remain confined to Euclidean settings, compact domains, or uniformly bounded noise. Our solution is motivated by Orabona (2020), who elegantly converts average-iterate convergence of SGD to last-iterate in-expectation convergence for non-smooth objectives within Euclidean domains. Through a careful analysis of martingales and a refinement of their techniques, we extend the framework to $\ell*$-smooth objectives in non-Euclidean settings. Leveraging this conversion framework, we prove that suboptimality gaps are bounded for all iterations simultaneously with high probability, indicating that the gradients are also bounded for all $t\in\mathbb{N}$. Such “anytime” characteristic is vital for our analysis, since Lemma 2.7 needs to be applied to every round of SMD.
Dual norm of the stochastic gradient   In the context of SCO, we still need to exploit the local smoothness property (Lemma 2.7), which requires the distance between consecutive iterates to be bounded. In the deterministic case, we have $\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|\leq\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}$, and by Lemma 3.4, the boundedness of $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}$ is inferred from finite suboptimality gaps. In contrast, for the stochastic case, it holds that $\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|\leq\left\|\mathbf{g}_{t}\right\|_{*}$, where the challenge lies in bounding $\left\|\mathbf{g}_{t}\right\|_{*}$. Clearly, the self-bounding property in Lemma 3.4 fails to bridge the stochastic gradient and the suboptimality gap. To address this, we leverage the noise model in Assumption J.1 to connect the noise level with the dual norm of the true gradient. As the suboptimality gap is bounded with high probability, the boundedness of the true gradient follows directly, implying that the noise is almost surely bounded via $\sigma(\cdot)$. Finally, we apply a union bound over these events to establish bounds on the stochastic gradient.
Based on the above reasoning, our analysis framework significantly differs from that of Li et al. (2023a, Section 5.2), who utilize a stopping-time technique (Williams, 1991) to show that the two key quantities are bounded before the stopping time $T$ with high probability. However, their method has two limitations: (i) it requires prior knowledge of the total number of iterations $T$; (ii) it lacks a convergence guarantee for the outputs generated before the final round. Moreover, their approach is ill-suited for SMD, as it heavily relies on the Euclidean inner product such that $\langle\mathbf{x},\mathbf{x}\rangle=\left\|\mathbf{x}\right\|_{2}^{2}$. Our main result is presented in Theorem J.3, whose proof is deferred to Section J.4.
Theorem J.3.
Under Assumptions 2.5, J.2, 3.3 and J.1, set
$0<\eta_{t}\leq\min\left\{\frac{D}{\sigma(\left\|\nabla f(\mathbf{x}_{t-1})\right\|_{*})\sqrt{t}},\right.\left.\frac{\min\left\{1,\left\|\nabla f(\mathbf{x}_{t-1})\right\|_{*}/\sigma(\left\|\nabla f(\mathbf{x}_{t-1})\right\|_{*})\right\}}{2\ell(2\left\|\nabla f(\mathbf{x}_{t-1})\right\|_{*})}\right\}.$ (169)
With high probability, we have the following anytime convergence rate:
$f(\overline{\mathbf{x}}_{t})-f^{*}=\widetilde{O}\left(\frac{1}{\sqrt{t}}\right),\quad\forall t\in\mathbb{N}_{+},$ (170)
where $\overline{\mathbf{x}}_{t}=\sum_{s=1}^{t}\eta_{s}\mathbf{x}_{s}/\sum_{s=1}^{t}\eta_{s}$. Moreover, the gradients along the trajectory satisfy $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\allowbreak=\widetilde{O}(1),\forall t\in\mathbb{N}$.
Remark J.4.
Theorem J.3 achieves the classic $\widetilde{O}(1/\sqrt{t})$ anytime convergence rate (Shamir and Zhang, 2013; Zhang et al., 2024a; Bai et al., 2025) in the classic SCO. Note that existing results rely on a deterministic learning rate schedule proportional to $1/\sqrt{t}$. However, this strategy is too coarse to capture the local features of the objective under $\ell*$-smoothness. Consequently, we incorporate local curvature information into the learning rates, akin to the approach in Xie et al. (2024b), who consider online convex optimization under $\ell$-smoothness.
J.3 Comparison of Noise Assumptions
Below, we briefly compare Assumption J.1 with commonly used assumptions in stochastic optimization.
Bounded noise
When $\sigma(\cdot)$ degenerates into a positive constant, Assumption J.1 recovers the uniformly bounded noise condition in the seminal work on $(L_{0},L_{1})$-smoothness (Zhang et al., 2020b), and it is also commonly adopted in other subfields (Koloskova et al., 2023; Zhang et al., 2023; Yu et al., 2024).
Affine noise
If $\sigma(\alpha)=\sigma_{0}+\sigma_{1}\alpha$ or $\sigma(\alpha)=\sqrt{\sigma_{0}^{2}+\sigma_{1}^{2}\alpha^{2}}$Assumption J.1 recovers the almost sure $(\sigma_{0},\sigma_{1})$-affine noise condition in Liu et al. (2023b); Hong and Lin (2024) and Attia and Koren (2023), respectively. Its in-expectation version is also widely adopted in the existing literature (Shi et al., 2021; Jin et al., 2021; Wang et al., 2023a; Jiang et al., 2024).
Finite moment
Another popular setting is that the stochastic gradient has bounded $p$-th moment for some $p\in(1,2]$ (Sadiev et al., 2023; Liu et al., 2024), which includes the uniformly bounded variance condition (Ghadimi and Lan, 2012, 2013; Arjevani et al., 2023) with $p=2$. We underline that the generalized bounded noise is never stronger than this finite moment condition (cf. Proposition J.5).
Sub-Gaussian noise
This condition is strictly stronger than the finite variance condition and is extremely effective in deriving high-probability bounds (Juditsky et al., 2011; Liu and Zhou, 2024). As demonstrated by Proposition J.5Assumption J.1 can not imply sub-Gaussian noise.
Generalized heavy-tailed noise
Recently, Liu and Zhou (2025a) propose a relaxation of the traditional heavy-tailed noises assumption (e.g., finite moment condition) in the form $\mathbb{E}_{t-1}[\left\|\boldsymbol{\epsilon}_{t}\right\|_{2}^{p}]\leq\sigma_{0}^{p}+\sigma_{1}^{p}\left\|\nabla f(\mathbf{x}_{t})\right\|_{2}^{p}$ for some $p\in(1,2],\sigma_{0},\sigma_{1}\in\mathbb{R}_{+}$. The almost sure version of the generalized heavy-tailed noise corresponds to our formulation with $\sigma(\alpha)=\sigma_{0}^{p}+\sigma_{1}^{p}\alpha^{p}$. While Assumption J.1 and this condition do not imply each other, both can model heavy-tailed noise environments.
In real-world applications, it is common that the noise level is positively correlated with the true signal, as considered in most multiplicative noise models (Sancho et al., 1982; López-Martínez and Fabregas, 2003; Aubert and Aujol, 2008; Hodgkinson and Mahoney, 2021). To this end, we leverage an arbitrary non-decreasing polynomial $\sigma(\cdot)$ to capture this relation. We believe this formulation is expressive enough to accommodate diverse stochastic environments where the noise is potentially unbounded (Hwang, 1986; Loh and Wainwright, 2011; Diakonikolas et al., 2023) or has heavy tails (Zhang and Zhou, 2018; Gurbuzbalaban et al., 2021; Cutkosky and Mehta, 2021).
Lastly, we present a simple proof to demonstrate the relative strength of our generalized bounded noise assumption compared to the finite moment condition. A similar argument is presented in Liu and Zhou (2025a, Example A.1), who construct a counterexample that satisfies the generalized heavy-tailed noise condition but does not satisfy the finite moment condition.
Proposition J.5.
Assumption J.1 does not imply the finite moment condition given below:
$\forall t\in\mathbb{N},p\in(1,2],\quad\mathbb{E}_{t-1}[\left\|\boldsymbol{\epsilon}_{t}\right\|_{*}^{p}]<+\infty,$ (171)
where the expectation $\mathbb{E}_{t-1}$ is conditioned on the past stochasticity $\left\{\boldsymbol{\epsilon}_{s}\right\}_{s=0}^{t-1}$.
Proof.
Consider the special case where
[…]
In this case, the finite moment condition degenerates to the finite variance assumption (Ghadimi and Lan, 2012, 2013; Arjevani et al., 2023). It suffices to show that the variance could be unbounded under Assumption J.1 with equality condition in (iii). We proceed with
$\mathbb{E}_{t-1}\left[\left\|\boldsymbol{\epsilon}_{t}\right\|_{*}^{2}\right]=$ $\mathbb{E}_{t-1}\left[\sigma_{0}+\sigma_{1}\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\right]^{2}$ (172)
$=$ $\sigma_{0}^{2}+2\sigma_{0}\sigma_{1}\mathbb{E}_{t-1}\left[\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\right]+\sigma_{1}^{2}\mathbb{E}_{t-1}\left[\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}^{2}\right].$
Intuitively, $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}$ can not be directly controlled and thus could potentially diverge to infinity. For instance, consider a more extreme case where $f(\mathbf{x})=\mathbf{x}^{2}$ and the algorithm outputs $\mathbf{x}_{t}=t\mathbf{x}_{0}$. In this way, we have
$\mathbb{E}_{t-1}\left[\left\|\boldsymbol{\epsilon}_{t}\right\|_{*}^{2}\right]=$ $\sigma_{0}^{2}+4\sigma_{0}\sigma_{1}\mathbb{E}_{t-1}\left[\left\|\mathbf{x}_{t}\right\|_{*}\right]+4\sigma_{1}^{2}\mathbb{E}_{t-1}\left[\left\|\mathbf{x}_{t}\right\|_{*}^{2}\right]$ (173)
$=$ $\sigma_{0}^{2}+4t\sigma_{0}\sigma_{1}\mathbb{E}_{t-1}\left\|\mathbf{x}_{0}\right\|_{*}+4t^{2}\sigma_{1}^{2}\mathbb{E}_{t-1}\left\|\mathbf{x}_{0}\right\|_{*}^{2}\to\infty,\text{ if }t\to\infty,$
which clearly contradicts the finite variance condition. Hence, Assumption J.1 is not strictly stronger than the finite moment condition. ∎
J.4 Proof of Theorem J.3
The lemmas in this subsection hold under the conditions in Theorem J.3. We do not restate this point for simplicity. First, we present the following two technical lemmas. Lemma J.6 is the Hoeffding-Azuma inequality for martingales, which is often used as a standard tool to derive the high probability bound. Lemma J.7 is a powerful technique to transform from the convergence behavior from the average-iterate to the last-iterate.
Lemma J.6.
(Cesa-Bianchi and Lugosi, 2006) Let $V_{1},V_{2},\ldots$ be a martingale difference sequence with respect to some sequence $X_{1},X_{2},\ldots$ such that $V_{i}\in[A_{i},A_{i}+c_{i}]$ for some random variable $A_{i}$, measurable with respect to $X_{1},\ldots,X_{i-1}$ and a positive constant $c_{i}$. If $S_{n}=\sum_{i=1}^{n}V_{i}$, then for any $t>0$,
$\mathbb{P}[S_{n}>t]\leq\exp\left(-\frac{2t^{2}}{\sum_{i=1}^{n}c_{i}^{2}}\right).$ (174)
Lemma J.7.
(Orabona, 2020, Lemma 1) Let $\left\{\eta_{s}\right\}_{s=1}^{t}$ be a non-increasing sequence of positive numbers and $\left\{q_{s}\right\}_{s=1}^{t}$ be a sequence of positive numbers. Then
$\eta_{t}q_{t}\leq\frac{1}{t}\sum_{s=1}^{t}\eta_{s}q_{s}+\sum_{k=1}^{t-1}\frac{1}{k(k+1)}\sum_{s=t-k+1}^{t}\eta_{s}\left(q_{s}-q_{t-k}\right).$ (175)
Next, we move on to the dynamics of SMD defined in (14). For convenience, we define the following quantities, which are frequently used in our analysis.
$\forall t\in\mathbb{N},\quad G_{t}=\left\|\nabla f(\mathbf{x}_{t})\right\|_{*},\ L_{t}=\ell(2G_{t}),\ \sigma_{t}=\sigma(G_{t})$ (176)
Lemma J.8.
If $\eta_{t+1}\leq\min\left\{1/(2L_{t}),G_{t}/(2L_{t}\sigma_{t})\right\}$, then for any $\mathbf{x}\in\mathcal{X}$ we have
$\eta_{t+1}[f(\mathbf{x}_{t+1})-f(\mathbf{x})]\leq B(\mathbf{x},\mathbf{x}_{t})-B(\mathbf{x},\mathbf{x}_{t+1})+\eta_{t+1}\langle\boldsymbol{\epsilon}_{t},\mathbf{x}-\mathbf{x}_{t}\rangle+\eta_{t+1}^{2}\sigma_{t}^{2}.$ (177)
Proof.
By Equation 139, we can derive
$\forall\mathbf{x}\in\mathcal{X},\ \langle\eta_{t+1}\nabla f(\mathbf{x}_{t}),\mathbf{x}_{t+1}-\mathbf{x}\rangle\leq$ $B(\mathbf{x},\mathbf{x}_{t})-B(\mathbf{x},\mathbf{x}_{t+1})-B(\mathbf{x}_{t+1},\mathbf{x}_{t})$ (178)
$+\eta_{t+1}\langle\nabla f(\mathbf{x}_{t})-\mathbf{g}_{t},\mathbf{x}_{t+1}-\mathbf{x}\rangle.$
In view of the step size $\eta_{t+1}\leq\min\left\{1/(2L_{t}),G_{t}/(2L_{t}\sigma_{t})\right\}$, we exploit the generalized smoothness property as follows
$\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|\leq$ $\eta_{t+1}\left\|\mathbf{g}_{t}\right\|_{*}\leq\eta_{t+1}(\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}+\left\|\boldsymbol{\epsilon}_{t}\right\|_{*})$ (179)
$\leq$ $\frac{G_{t}}{2L_{t}}+\frac{G_{t}}{2L_{t}\sigma_{t}}\cdot\sigma(\left\|\nabla f(\mathbf{x}_{t})\right\|_{*})\leq\frac{G_{t}}{L_{t}},$
where the first step is due to Lemma E.1; the last step leverages the non-decreasing property of $\sigma(\cdot)$. Now that (179) holds, we can utilize the local smoothness property in Lemma 2.7 and we follow the same steps as in (64) and (65) to derive
$\langle\eta_{t+1}\nabla f(\mathbf{x}_{t}),\mathbf{x}_{t+1}-\mathbf{x}_{*}\rangle\geq\eta_{t+1}(f(\mathbf{x}_{t+1})-f^{*})-\frac{\eta_{t+1}L}{2}\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|^{2}.$ (180)
Combining the above inequality with (178), we obtain
$\eta_{t+1}[f(\mathbf{x}_{t+1})-f(\mathbf{x})]\leq$ $B(\mathbf{x},\mathbf{x}_{t})-B(\mathbf{x},\mathbf{x}_{t+1})-B(\mathbf{x}_{t+1},\mathbf{x}_{t})+\frac{\eta_{t+1}L}{2}\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|^{2}$ (181)
$+\eta_{t+1}\langle\boldsymbol{\epsilon}_{t},\mathbf{x}-\mathbf{x}_{t}\rangle+\eta_{t+1}\langle\boldsymbol{\epsilon}_{t},\mathbf{x}_{t}-\mathbf{x}_{t+1}\rangle$
$\leq$ $B(\mathbf{x},\mathbf{x}_{t})-B(\mathbf{x},\mathbf{x}_{t+1})+\eta_{t+1}\langle\boldsymbol{\epsilon}_{t},\mathbf{x}-\mathbf{x}_{t}\rangle-\frac{1}{2}\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|^{2}$
$+\frac{1}{4}\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|^{2}+\eta_{t+1}^{2}\left\|\boldsymbol{\epsilon}_{t}\right\|_{*}^{2}+\frac{1}{4}\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|^{2}$
$\leq$ $B(\mathbf{x},\mathbf{x}_{t})-B(\mathbf{x},\mathbf{x}_{t+1})+\eta_{t+1}\langle\boldsymbol{\epsilon}_{t},\mathbf{x}-\mathbf{x}_{t}\rangle+\eta_{t+1}^{2}\sigma_{t}^{2},$
where the second inequality uses Lemma D.2 and $\eta_{t+1}\leq 1/(2L_{t})$. ∎
Lemma J.9.
For any $\delta\in(0,1)$, define
[…]
$\widetilde{G}_{t}=\sup\left\{\alpha\in\mathbb{R}_{+}|\alpha^{2}\leq 2\widetilde{F}_{t}\ell(2\alpha)\right\},\quad\forall t\in\mathbb{N}_{+}.$
Under Assumptions 3.3 and J.1, with probability at least $1-\delta$, the following holds simultaneously for all $t\in\mathbb{N}_{+}$:
$f(\mathbf{x}_{t})-f^{*}\leq\widetilde{F}_{t}=\widetilde{O}(1),\ G_{t}\leq\widetilde{G}_{t}=\widetilde{O}(1),\ L_{t}=\widetilde{O}(1),\ \sigma_{t}=\widetilde{O}(1).$ (183)
Proof.
We apply Equation 177 to get
$\sum_{s=t_{1}}^{t_{2}}\eta_{s}\left[f(\mathbf{x}_{s})-f(\mathbf{x})\right]\leq B(\mathbf{x},\mathbf{x}_{t_{1}-1})-B(\mathbf{x},\mathbf{x}_{t_{2}})+\sum_{s=t_{1}-1}^{t_{2}-1}\left(\eta_{s+1}\langle\boldsymbol{\epsilon}_{s},\mathbf{x}-\mathbf{x}_{s}\rangle+\eta_{s+1}^{2}\sigma_{s}^{2}\right),$ (184)
which holds for any $\mathbf{x}\in\mathcal{X}$ and any indices $t_{1}\leq t_{2},t_{1},t_{2}\in\mathbb{N}$. For convenience, we define
$q_{s}:=f(\mathbf{x}_{s})-f^{*},\ V_{s}:=\eta_{s+1}\langle\boldsymbol{\epsilon}_{s},\mathbf{x}_{*}-\mathbf{x}_{s}\rangle,\ U_{s}^{k}:=\eta_{s+1}\langle\boldsymbol{\epsilon}_{s},\mathbf{x}_{t-k}-\mathbf{x}_{s}\rangle$ (185)
for all $0\leq s\leq t$, where $\mathbf{x}_{*}\in\operatorname*{argmin}_{\mathbf{x}\in\mathcal{X}}f(\mathbf{x})$ (the existence of $\mathbf{x}_{*}$ can be implied by Assumptions 2.5 and J.2). Obviously, we have $q_{s}\geq 0$ and $\left\{V_{s}\right\}_{s=0}^{t-1},\left\{U_{s}^{k}\right\}_{s=t-k+1}^{t-1}$ are martingale difference sequences. Then we have
[…]
The above bound usually relates to the average-iterate. To incorporate Lemma J.7, we consider the following relation:
$\sum_{s=t-k+1}^{t}\eta_{s}(q_{s}-q_{t-k})=\sum_{s=t-k+1}^{t}\eta_{s}[f(\mathbf{x}_{s})-f(\mathbf{x}_{t-k})]$ (187)
[…]
$\leq$ $\sum_{s=t-k+1}^{t}\eta_{s}^{2}\sigma_{s-1}^{2}+\sum_{s=t-k+1}^{t-1}U_{s}^{k}.$
With these preparations, we invoke Lemma J.7 to deduce
$\eta_{t}q_{t}\leq$ $\frac{1}{t}\sum_{s=1}^{t}\eta_{s}q_{s}+\sum_{k=1}^{t-1}\frac{1}{k(k+1)}\sum_{s=t-k+1}^{t}\eta_{s}\left(q_{s}-q_{t-k}\right)$ (188)
$\leq$ $\frac{D^{2}}{t}+\frac{1}{t}\sum_{s=1}^{t}\eta_{s}^{2}\sigma_{s-1}^{2}+\frac{1}{t}\sum_{s=0}^{t-1}V_{s}^{t}$
$+\sum_{k=1}^{t-1}\frac{1}{k(k+1)}\sum_{s=t-k+1}^{t}\eta_{s}^{2}\sigma_{s-1}^{2}+\sum_{k=1}^{t-1}\frac{1}{k(k+1)}\sum_{s=t-k+1}^{t-1}U_{s}^{k}.$
Next, we bound the terms in (188). Recalling the definition of $\eta_{s}\leq D/(\sigma_{s-1}\sqrt{s})$, we have
$\sum_{s=1}^{t}\eta_{s}^{2}\sigma_{s-1}^{2}$ $\leq D^{2}\sum_{s=1}^{t}\frac{1}{s}\leq D^{2}+D^{2}\int_{1}^{t}\frac{1}{s}\mathrm{d}s=D^{2}\left(1+\ln{t}\right),$ (189)
$\sum_{s=t-k+1}^{t}\eta_{s}^{2}\sigma_{s-1}^{2}$ $\leq D^{2}\sum_{s=t-k+1}^{t}\frac{1}{s}\leq D^{2}\int_{t-k}^{t}\frac{1}{s}\mathrm{d}s=D^{2}\ln\left(\frac{t}{t-k}\right)\leq\frac{kD^{2}}{t-k}.$
Then
$\sum_{k=1}^{t-1}\frac{1}{k(k+1)}\sum_{s=t-k+1}^{t}\eta_{s}^{2}\sigma_{s-1}^{2}\leq\sum_{k=1}^{t-1}\frac{D^{2}}{k(k+1)}\cdot\frac{k+1}{t-k}$ (190)
$=$ $\sum_{k=1}^{t-1}\frac{D^{2}}{kt}+\sum_{k=1}^{t-1}\frac{D^{2}}{(t-k)t}=\frac{2D^{2}}{t}\sum_{k=1}^{t-1}\frac{1}{t}\leq\frac{2D^{2}[1+\ln(t-1)]}{t}.$
Next, we utilize Lemma J.6 to bound the sum of the martingale difference sequence. In view of $\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|\leq 2\sqrt{2}D,\forall\mathbf{x}_{1},\mathbf{x}_{2}\in\mathcal{X}$ (cf. (Yu et al., 2024, Proposition A.6)), we have
$V_{s}\leq\eta_{s+1}\left\|\boldsymbol{\epsilon}_{s}\right\|_{*}\left\|\mathbf{x}_{*}-\mathbf{x}_{s}\right\|\leq\frac{D}{\sigma_{s}\sqrt{s+1}}\sigma_{s}\cdot 2\sqrt{2}D=\frac{2\sqrt{2}D^{2}}{\sqrt{s+1}},$ (191)
$U_{s}^{k}\leq\eta_{s+1}\left\|\boldsymbol{\epsilon}_{s}\right\|_{*}\left\|\mathbf{x}_{t-k}-\mathbf{x}_{s}\right\|\leq\frac{D}{\sigma_{s}\sqrt{s+1}}\sigma_{s}\cdot 2\sqrt{2}D=\frac{2\sqrt{2}D^{2}}{\sqrt{s+1}}.$
where we make use of $\eta_{s}\leq D/(\sigma_{s-1}\sqrt{s})$. Hence, we invoke Lemma J.6 to deduce that with probability at least $1-\delta/t$,
$\sum_{s=0}^{t-1}V_{s}^{t}\leq\sqrt{\frac{1}{2}\sum_{s=0}^{t-1}\left(\frac{2\sqrt{2}D^{2}}{\sqrt{s+1}}\right)^{2}\ln\frac{1}{\delta}}\leq 2D^{2}\sqrt{\left(1+\ln t\right)\ln\frac{t}{\delta}}.$ (192)
Similarly, for a given index $k$, with probability at least $1-\delta/t$,
$\sum_{s=t-k+1}^{t-1}U_{s}^{k}\leq\sqrt{\frac{1}{2}\sum_{s=t-k+1}^{t-1}\left(\frac{2\sqrt{2}D^{2}}{\sqrt{s+1}}\right)^{2}\ln\frac{1}{\delta}}\leq 2D^{2}\sqrt{\ln\left(\frac{t}{t-k+1}\right)\ln\frac{t}{\delta}}.$ (193)
Taking the union bound over $k\in[t-1]$ and $\left\{V_{s}^{t}\right\}_{s=0}^{t-1}$, we conclude that with probability at least $1-\delta$,
$\sum_{s=0}^{t-1}V_{s}^{t}\leq 2D^{2}\sqrt{\left(1+\ln t\right)\ln\frac{t}{\delta}},\ \sum_{s=t-k+1}^{t-1}U_{s}^{k}\leq 2D^{2}\sqrt{\ln\left(\frac{t}{t-k+1}\right)\ln\frac{t}{\delta}},$ (194)
hold simultaneously for all $k\in[t-1]$. We proceed to bound the second term.
$\sum_{k=1}^{t-1}\frac{1}{k(k+1)}\sum_{s=t-k+1}^{t-1}U_{s}^{k}$ (195)
$\leq$ $2D^{2}\sqrt{\ln\frac{t}{\delta}}\sum_{k=1}^{t-1}\frac{1}{k(k+1)}\cdot\ln\left(1+\frac{k}{t-k}\right)\cdot\frac{1}{\sqrt{\ln\left(\frac{t}{t-k}\right)}}$
$\leq$ $2D^{2}\sqrt{\ln\frac{t}{\delta}}\underbrace{\sum_{k=1}^{t-1}\frac{1}{k(t-k)}\cdot\frac{1}{\sqrt{\ln\left(\frac{t}{t-k}\right)}}}_{\texttt{term-A}}.$
Then,
$\texttt{term-A}=$ $\sum_{k=1}^{t-1}\frac{1}{kt}\cdot\frac{1}{\sqrt{\ln\left(\frac{t}{t-k}\right)}}+\sum_{k=1}^{t-1}\frac{1}{t(t-k)}\cdot\frac{1}{\sqrt{\ln\left(\frac{t}{t-k}\right)}}$ (196)
$=$ $\frac{1}{t}\sum_{k=1}^{t-1}\frac{1}{k}\left(\frac{1}{\sqrt{\ln\left(\frac{t}{t-k}\right)}}+\frac{1}{\sqrt{\ln\left(\frac{t}{k}\right)}}\right)$
[…]
where we use the relation $x/(1+x)\leq\ln(1+x),\forall x>-1$ and the following fact
$\frac{1}{\sqrt{\ln\left(\frac{t}{t-k}\right)}}+\frac{1}{\sqrt{\ln\left(\frac{t}{k}\right)}}\in\left(\frac{2}{\sqrt{\ln 2}},\frac{1}{\sqrt{\ln\left(1+\frac{1}{t-1}\right)}}+\frac{1}{\sqrt{\ln{t}}}\right].$ (197)
Now that each term in (188) has been calculated, we combine all these results to obtain
$\eta_{t}q_{t}\leq$ $\frac{D^{2}}{t}+\frac{D^{2}(1+\ln t)}{t}+\frac{2D^{2}[1+\ln(t-1)]}{t}$ (198)
$+\frac{2D^{2}\sqrt{\left(1+\ln t\right)\ln\frac{t}{\delta}}}{t}+2D^{2}\sqrt{\ln\frac{t}{\delta}}\left(\frac{1+\ln t}{\sqrt{t}}+\frac{1+\ln t}{t\sqrt{\ln t}}\right).$
which holds with probability at least $1-\delta$. Dividing by $\eta_{t}$ and taking union bound on $\left\{q_{s}\right\}_{s\in\mathbb{N}_{+}}$ (Zhang et al., 2024a), we obtain
$\forall t\in\mathbb{N}_{+},\quad q_{t}\leq$ $\frac{2L_{t-1}D^{2}(4+3\ln t)}{t}+\frac{4L_{t-1}D^{2}\sqrt{\left(1+\ln t\right)\ln\frac{2t^{3}}{\delta}}}{t}$ (199)
$+4L_{t-1}D^{2}\sqrt{\ln\frac{2t^{3}}{\delta}}\left(\frac{1+\ln t}{\sqrt{t}}+\frac{1+\ln t}{t\sqrt{\ln t}}\right)$
$+\frac{D\sigma_{t-1}(4+3\ln t)}{\sqrt{t}}+\frac{2D\sigma_{t-1}\sqrt{\left(1+\ln t\right)\ln\frac{2t^{3}}{\delta}}}{\sqrt{t}}$
$+2D\sigma_{t-1}\sqrt{\ln\frac{2t^{3}}{\delta}}(1+\ln t)+2D\sigma_{t-1}\sqrt{\ln\frac{2t^{3}}{\delta}}\frac{1+\ln t}{\sqrt{t\ln t}},$
where we use the well-known fact $\sum_{t=1}^{+\infty}1/t^{2}=\pi^{2}/6<2$. Denote by $F_{t}$ as the RHS of the above inequality. For all $2\leq t\leq T$, we have
[…]
We will show $\widetilde{F}_{t}=\widetilde{O}(1)$ by induction in the sequel. For the base case, we have $G_{0}=\left\|\nabla f(\mathbf{x}_{0})\right\|_{*}<\infty$. Hence $L_{0},\sigma_{0}$ are absolute constants. Since the high probability bound (199) holds, we clearly have $F_{1}\leq\widetilde{F}_{1}<\infty$ and $G_{1}\leq\widetilde{G}_{1}<\infty$ by Lemma D.3. For the induction step, we have $G_{t-1}=\widetilde{O}(1)$ from the induction basis. Then under Assumptions 3.3 and J.1,
$L_{t-1}=\ell(2G_{t-1})=\widetilde{O}(1),\ \sigma_{t-1}=\sigma(G_{t-1})=\widetilde{O}(1),$ (201)
hold with probability at least $1-\delta$, which implies $F_{t}=\widetilde{O}(1)$. Then we invoke Lemma D.3 again to deduce that $G_{t}=\widetilde{O}(1)$. Combining the base case and the induction step finishes the proof. ∎
Proof of Theorem J.3.
In the proof of Lemma J.9, we have shown that with probability at least $1-\delta$ (see (186), (189), (194) and (199)),
$\sum_{s=1}^{t}\eta_{s}q_{s}\leq D^{2}+D^{2}(1+\ln t)+2D^{2}\sqrt{\left(1+\ln t\right)\ln\frac{2t^{3}}{\delta}},\quad\forall t\in\mathbb{N}_{+}.$ (202)
For the average-iterate up to $t$-th round, by the convexity of $f$ and Jensen’s Inequality, we have
$f(\overline{\mathbf{x}}_{t})-f^{*}\leq\frac{\sum_{s=1}^{t}\eta_{s}q_{s}}{\sum_{s=1}^{t}\eta_{s}}\leq\frac{\overbrace{D^{2}(2+\ln t)+2D^{2}\sqrt{\left(1+\ln t\right)\ln\frac{2t^{3}}{\delta}}}^{D_{t}}}{\sum_{s=1}^{t}\eta_{s}}$ (203)
$\leq$ $\frac{D_{t}}{\sum_{s=1}^{t}\frac{\min\left\{1,G_{s-1}/\sigma_{s-1}\right\}}{2L_{s-1}}}+\frac{D_{t}}{\sum_{s=1}^{t}\frac{D}{\sigma_{s-1}\sqrt{s}}}\leq\frac{\widetilde{L}_{t-1}^{\max}D_{t}}{t}+\frac{\sigma_{t-1}^{\max}D_{t}}{D\sqrt{t}},$
where the second line is derived via the flooring technique (Liu et al., 2023c; Liu and Zhou, 2024); the last line uses $\sum_{s=1}^{t}1/\sqrt{s}\geq\sqrt{t}$ and the definitions below:
$\widetilde{L}_{t}^{\max}:=\max\left\{L_{0},\max_{1\leq s\leq t}\left[L_{s}\max\left(1,\frac{\sigma_{s-1}}{G_{s-1}}\right)\right]\right\},\quad\sigma_{t}^{\max}:=\max_{0\leq s\leq t}\sigma_{s}.$ (204)
By Lemma J.9, we have $G_{t}=\widetilde{O}(1),\ L_{t}=\widetilde{O}(1),\ \sigma_{t}=\widetilde{O}(1)$ that holds with high probability along the trajectory. Therefore, we have $\sigma_{t-1}^{\max}=\widetilde{O}(1),\widetilde{L}_{t-1}^{\max}=\widetilde{O}(1)$. In view of $D_{t}=\widetilde{O}(1)$, we conclude our proof by combining these quantities:
$f(\overline{\mathbf{x}}_{t})-f^{*}\leq$ $\frac{\widetilde{L}_{t-1}^{\max}\left(D^{2}(2+\ln t)+2D^{2}\sqrt{\left(1+\ln t\right)\ln\frac{2t^{3}}{\delta}}\right)}{t}$ (205)
$+\frac{\sigma_{t-1}^{\max}\left(D(2+\ln t)+2D\sqrt{\left(1+\ln t\right)\ln\frac{2t^{3}}{\delta}}\right)}{\sqrt{t}}=\widetilde{O}\left(\frac{1}{\sqrt{t}}\right).$
Appendix K Analysis for Non-convex Composite Mirror Descent
First, we introduce a standard yet useful lemma.
Lemma K.1 (Lemma 6.4 in Lan (2020)).
Let $\mathcal{G}_{t}$ be defined in (16), it holds that
$\langle\nabla f(\mathbf{x}_{t}),\mathcal{G}_{t}\rangle\geq\left\|\mathcal{G}_{t}\right\|^{2}+\frac{\phi(\mathbf{x}_{t+1})-\phi(\mathbf{x}_{t})}{\eta}.$
The proof of Theorem 5.1 stems from the textbook analysis of mirror descent in the non-convex case (Lan, 2020, Theorem 6.5), together with our previous generalized smooth analysis.
Proof of Theorem 5.1.
First, we provide a quick and intuitive explanation for our proof outline. Recall Lemma E.2, which bounds the suboptimality gaps $\left\{f(\mathbf{x}_{t})-f^{*}\right\}$ along the optimization trajectory without convexity. This is exactly the key to the derivations in the non-convex case. Since we have $f(\mathbf{x}_{t})-f^{*}\leq F$ and $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G$ for all $t\in\mathbb{N}$, we immediately leverage the effective smoothness constant $L$ and reduce the analysis for $\ell*$-smooth functions to the textbook analysis of standard $L$-smooth functions. Below, we present formal reasoning by induction.
When $t=0$, the induction basis is trivially satisfied. Suppose for all $s\leq t-1$ we have bounded suboptimality gaps and bounded gradients. Now that the conditions of Lemma 2.7 is satisfied, we apply Lemma 2.7 to the $t$-th iterate:
$f(\mathbf{x}_{t+1})\leq$ $f(\mathbf{x}_{t})+\langle\nabla f(\mathbf{x}_{t}),\mathbf{x}_{t+1}-\mathbf{x}_{t}\rangle+\frac{L}{2}\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|^{2}$
$\leq$ $f(\mathbf{x}_{t})-\eta\langle\nabla f(\mathbf{x}_{t}),\mathcal{G}_{t}\rangle+\frac{\eta^{2}L}{2}\left\|\mathcal{G}_{t}\right\|^{2}$
[…]
Since $\eta\leq 1/L$, after rearrangement we get
$f(\mathbf{x}_{t+1})\leq F(\mathbf{x}_{t+1})\leq F(\mathbf{x}_{t})-\frac{\eta}{2}\left\|\mathcal{G}_{t}\right\|^{2}\leq F(\mathbf{x}_{t})\leq\cdots\leq F(\mathbf{x}_{0}),$ (206)
which implies $f(\mathbf{x}_{t+1})-f^{*}\leq f(\mathbf{x}_{0})-f^{*}+\phi(\mathbf{x}_{0})$. So we deduce $\left\|\nabla f(\mathbf{x}_{t+1})\right\|_{*}\leq G$ by Lemma D.3. Hence, throughout the optimization trajectory, the bounded suboptimality gaps as well as gradients hold uniformly. Thus, we can start from (206) to perform a telescoping:
$\frac{1}{T}\sum_{t=0}^{t-1}\left\|\mathcal{G}_{t}\right\|^{2}\leq\frac{F(\mathbf{x}_{0})-F(\mathbf{x}_{T})}{\eta T}\leq\frac{F(\mathbf{x}_{0})-F^{*}}{\eta T},$
which completes the proof. ∎